A Complete Bibliography of Publications in North American Journal of Fisheries Management: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

09 June 2023 Version 1.01

Title word cross-reference

F_{1} [KH82]. μ [TH86a]. × [BR86, GA87, GM85, Gow86, JSC88, KH82, LJS82].

-ethanol [GB83].

1 [KH82]. 1-4 [GH83]. 1-Year-Old [KK84]. 12.0-Inch [Par84]. 1254 [BMM81]. 13 [Pit89]. 14.0-Inch [Par82]. 15-cm [Laa81]. 15.0-Inch [Nov87]. 17-Megawatt [BK85]. 174 [GB83].

2 [GB83]. 2- [GB83].

3-Year-Old [Nic84]. 305-mm [AO88].

500- [TH86a]. 58-Year [OC89].

7 [SB83].

Ability [GC84]. Abundance [Atk89, BMM87, DOFJ81, DPK84, FNN82, Gis85, HN82b, HR89, JS83, Kal87, KMR89, Kim88, MSB84, Mit81, NBH85, New89, RC89, RT85, Rob86, SM83, SB83]. AC [Hud85, SH86b]. Acceptance [AN84]. Accuracy [FGM84, MSG86, PB87, SM87]. Acidified [GSDK89]. Acoustics [SA84]. Acrylic [Dim85]. Activity [Gil82, PB82]. Adirondack [CK88, CRK88]. Adjustments [WMG85]. Adult [AGN81, BK86, BK85, BKJ89, HES86, Mun82, Mur85, SM85, SC88, STP83, Wal88]. after [Cla85, CCS89, EH83, Fra89, GH83, HS89, PGR88, SH89]. Against [ME82]. Age [AGN81, CB83, CW85, CM83, Eri83, FF82, HT89, PT84]. Ageing [QR82]. Ages [SB88a]. Aging [Box86a, CS87]. Alabama [MSB84, TG85, Eck83, JD88, KDP86]. Alaska [LKL87, Llo87, Mun82, Mur85, MK89]. Alaskan [HMK86, MS85a, PB88]. Aliceville [TG85]. Alkaline [GSDK89]. Allegheny [SA84]. Alleviate [SA84]. Allozymic [MG88]. along [PBP83, PW84]. Altered [PT84]. Alternative [PW84]. Altitude [Hub88]. American [BK86, BK85, Eck83, Ken88, LM87, MWTO89, PM87, SM83]. among [SH87]. Anadromous [BNC^{+}86, POB85]. Analyses [LAA83]. Analysis [BFB85a, BP86b, FF82, GSO89, HT89, HW83, Jen85, LD87, LUBO85, MKM89]. Analyzing [Kim88]. Anchor [Bul83, EC82, FM87, TC82]. Anchoring [Neg82]. Anesthetic [GM87, Gil89]. Angeles [PDA82]. Angler [BKSM89, CSH^{+}87, Com82, Dot82, Ebb87, Har83, IR89, JD88, MA84, Mat83, Mor85a, Mor85b, Par84, Sch86, SGO88, WGS84, Whi89]. Angler-Collected [Whi89]. Angler-Induced [Dot82]. Angler-Supplied [Ebb87]. Anglers [AN84, Car82b, CH88, DW81, Mat86, Sta81]. Angling [Cob88, Mar87, Mos83, MMC87, PB87, TMCS81]. anguillarum [BSS82]. Annual [BC87, Bea83, Enn83, OC89, RT85, SHR83]. Annulus [CCP89, QR82]. Apache [Rin85]. Apalachicola [WC83, WC85]. Apparatus [LRLM84]. Application [Cam85, GEB86, MKM89, SRH^{+}83, TP84]. Applied [BM85]. Applying [AK86]. Approach [Jac89, VJL82]. Approximating [Gus88]. Aquarium [SWH85]. Aquashade [BHN82]. Aquatic [CCS89, Gil82, Gis85, GA81, KMR89, LP82, WGWP84]. Area [Alm87, CA87, DOFJ81, GD86, MM82, MHK^{+}87]. Areas [Pof85, Qui88]. Arizona [MM82, Mef83, Rin85]. Arkansas [FH82, RH82]. Aroclor [BMM81]. Artificial [BMW86, CC89, Dom87, EGT^{+}84, IB84, LJ89, Mos83, PPP89]. Artificially [HR81]. Assess [Gab84, MGS87, Scr88]. Assessing [AS87, GEB86, MB86]. Assessment [BP86a, BMM87, Dew89, Don87, Enn85, HMK89, Jen86, KH89, MS82, Ret84, Wal88]. Assinica [Gow86]. Associated [Jew86, KMR89, MW86, SKB82]. Assumption [CRK88]. Atlantic [Alm87, BSS82, BJM82, BOS87, CR86, CW85, HG86, HRDM84, MM88, MK88, NG87, OB87, POB85, Pof85, Red86, SHS85, SMJR89, SMU84, VJL82].

Atmospheric [GSDK89]. Attack [SH89]. Attempt [PH85]. Attempted [Mef83]. Attempts [SA84]. Attraction [BK86]. Attractors [TS82, TG85]. Australia [MSK89]. Availability [BZM84, PH88]. Avoidance [BB82]. aztecus [KPR^{+}89].

Backfilled [NT87]. Bacterin [BSS82]. Bait [GH89, Nie82, NO88]. Bait-Fish [Nie82]. Baited [PCC81]. Baitfish [NH88]. Baits [PPP89, PW87b, RB87]. Bank [KD87, Sch86, BP86b, GSO89, Ove85]. Bank- [Sch86]. Banks [BMW86, PBP83, PN85, STP83, Wal88]. Bar [SGS88, VHJB83]. Barkley [BJ84]. Barrier [BWCQ88, PB88, STP83]. Based [AK86, FB82, Gar82, WMG85]. Basin [LT89, Ray88, Win89]. Basis [Rei87]. Bass [AO88, BR88b, Box82, Box86b, Box87, BML^{+}86, BKSM89, CC89, CCS89, CP82, CCP89, Dew88, DJ87, DRWM87, DRM^{+}89, DPK84, Ebb87, FNN82, FK83, FHB87, Gab87, GM85, Hal86, HGH82, Hub88, JDTK87, KP86, LMW87, Low81, MSC81, MR87, MSB84, Mir86, MM87, MS85b, Nov87, Par82, Par84, PW87a, Par89a, PH85, PT84, PB82, PGR88, POS87, Qui89, RR89, SM85, SHB85, SHP87, SWH85, SM82, SKB82, Szt85, TC82, VHJB83, WGW88, WGS84, Whi89, WC83, Yea85]. Bay [Eck83, GD86, HEJ87, Szt85, Cam85, Chi89, Fle87, GD86, Mat83, SHR83]. Beach [Lyo86, PPB89]. Beaked [Atk89]. Beaver [RH82]. Bed [AH86]. Bedload [AH83, HAD83]. before [HS89]. Behavior [HM85, MWTO89]. below [FMW88]. Benefits [And89]. Benzocaine [Gil89]. Bering [FNW82, MR88, WT84]. Better [Mor85a]. Between [SM85, CH85, CK88, DPK84, HD87, LD87, MG88, ME82, MS85b, Rei88, Ser83, Ser87, WGWP84, WS87, WS89]. Bias [BR88b]. Bighead [FH82]. Bioeconomic [SJT^{+}82]. Biological [Dew88, MB89, MR87, WT84]. Biologists [Tay81]. Biology [TB86, WC83]. Biomass [Atk89, GSO89, HM86, LMW87, PN89, PJ82, RH82]. Biotic [AK86, AS87]. Biphenyl [BMM81]. Bird [Mat89]. Black [BKSM89, HBS83, Low81, MG88, SSHP85, SKB82]. Bloaters [Say87]. Blue [And89, CRS87, KP86, KNR88, SBB88, SRH^{+}83, SHR83, TMCS81, TP84, Zal87]. Blueback [Bul83, MEB84, PB81]. Bluefin [HM86]. Bluegill [Ber82, BZM84, BR86, Cob88, GA87, KH82]. Bluegills [Gab87, LJ89, MS85b, RM82]. Boat [DW81, FB82, MSG86]. Boat-Based [FB82]. Boats [EH89, SC87]. Boldt [Cla85]. Bon [Eck83]. Bootstrap [KB85]. Boreal [MMC87]. Bottle [FGC86]. Bottom [Atk89]. Boulder [TW87]. Bow [KH89]. Box [BJFF87]. Branding [BOS87]. Brands [FP85]. Brief [BKJ89, Hol87, LGJR89, Whi89]. Briefs [BTPM89, BWCQ88, Gus88, Hub88, Jen89, MSK89]. British [Bre87, GP88, SM87]. Brook [AH86, CAG81, CK88, CRK88, DA89, EH83, Fli83, Fra89, Gow86, HLS81, Hud85, MRL83]. Brown [BM85, BNC^{+}86, CAG81, DS88, EH86, HP82, Har83, KPR^{+}89, MM85, WGH87, Wor81]. Brunswick [CR86]. Buffer [BTB85]. Bullheads [HBS83]. Buoyant [DHBZ89]. Bypass [GSZ^{+}88, NG87].

Caged [KNR88]. Cages [PW87b]. Calcified [SB88a]. Calculation [WMG85]. Calibrated [SGS88]. California [Har86, HK88, ME82, MVG83]. Callinectes [CRS87]. Canada [AFL83, Don87, Ret84]. Canals [NT87]. Canceling [SGO88]. Canoes [LRLM84]. Canopy [PN89]. Capacitance [PN83]. Capture [BKJ89, FGC86, Lyo86, PPB89, Ueb86]. Captured [DRM^{+}89, Fle87]. Carlin [Kal89]. Carolina [BM85, Bar85, HEJ87, LWL89, LUBO85, MM85, PB81, SMU84]. Carp [FH82, GB83, HM85, Hil86, MB81, SB88b]. Case [HWPS84, POS87]. Catch [AN84, Bar89, BKSM89, Chi89, Cla83, EH86, Hal86, HT89, HS83, IR89, Jen85, LUBO85, Mat83, Mor85b, Ove85, Par82, Par89a, RP83, SGG86, SRJ82, Ser82a, Ser83, TP84, THHS89]. Catch-and-Release [AN84, Bar89, SGG86]. Catchability [BK82]. Catchable [HFG85, Mor85a, Roh87, Vin87]. Catchable-Size [HFG85, Vin87]. Catches [BP86a, BKSM89, DHBZ89, GH89, HM86, IR89, MH88, MR88, Nov87, PCC81, Qui88, Red86, Wal88, WGS84]. Catching [SLA85]. Categorization [Gab84]. Catfish [AGN81, BHN82, ES87, GH89, Joh87, KNR88, Mos83, New89, PW87b, SWH85, SD87, SN88, TMCS81]. Catostomids [SH86a]. Caught [CC89, Dew88, MSK89, PPP89, PGR88, PDA82, SHB85, SHP87]. Caused [Bre87, Dot82]. Cauterization [Box82]. Cave [Axo81, BKSM89]. Census [FNN82, SM87]. Centimeter [CM83]. Central [HG86, VJL82, LWL89, MM82, Zal87]. Challenge [FL85]. Chambers [KDP86]. Changes [BR88a, BZM84, DS88, Enn85, Ken88, MW86, PBB88]. Channel [AGN81, BHN82, BMW86, ES87, GH89, KNR88, Mos83, New89, SWH85, SD87, SN88, TMCS81]. Channelized [CK85, EGT^{+}84, HN82b, New89]. Channels [BP86a]. Characteristics [Dew88, Fer86, Gar82, GD86, MT87, MR87, Pof85, RR89]. Charlotte [AFL83]. Charter [DLP89]. Cheek [FHB87]. Cheese [GH89]. Chemical [MBRG83, Mef83, Wor81]. Chemicals [BMM81, GM87]. Chesapeake [Chi89, SHR83]. China [HW83]. Chinook [BNA83, BMW86, Car82a, EE85, FL85, GSZ^{+}88, KT84, Mun82, MHK^{+}87, MR88, PB88, Ray88, Rei87, RW88, RC89, SM83, Wer88, Win89]. Chionoecetes [TMO89]. Chum [Bil86, MZ87, Mur85]. Circulation [Mos83]. City [Bea83]. Clam [HEJ87, Tow85]. Class [Hau83, RC88]. Classes [RW88, Ser87]. Classification [LP82]. Clay [Gil82]. Cleaning [McK87]. Climatic [Kal87]. Clipping [BOS87, Box82]. Closing [EH89]. cm [Laa81]. Coast [LWL89, Bry83, Low81]. Coastal [BP86a, HB85b, KPR^{+}89, NT87, PDA82]. Cod [FF82]. Coded [ES86, FHB87, HC88, KP86, MZ87]. Coho [Bam85, BB82, Car82a, CSP88, HK88, HES86, MWD86, PC84, Rei88, RHT88, SJ85]. Cohort [HW83, Ser87]. Cold [BOS87, SBB88]. Cold-Branding [BOS87]. Coldwater [SB88b]. Collapse [Eva81]. Collected [HBDZ89, Whi89]. Collection [BJFF87]. Colorado [AN84, LT89, MB89, PB82, SB87, SSB85, TBM84, Tyu88]. Columbia [BNC^{+}86, DD89, PH89, BR88b, Bre87, FMW88, GS87, GP88, Ray88, SM87, Win89]. Comments [CW85, RT86, SH86b]. Commercial [BM85, BJ83, BJ84, CH85, Fer86, FJ87, Hea85, HL87, LUBO85, Mac82, MR86, PW87b, RT85, RC86, SSHP85, SRH^{+}83, THHS89, Wer88, Wol81]. Commercially [Dew88, HT89]. Common [GB83, TGPS86]. Common-Property [TGPS86]. Communities [AK86, MS82, NT87]. Community [AS87, KDP86, MGS87]. Comparative [GM87, GD86, Joh87]. Compare [POS87]. Comparison [Bam85, BA88, Box86a, CS87, CK88, DHBZ89, EOS88, FM87, HLS81, HHHH83, HES86, NT87, Par89a, PW87b, PC84, RB87, ZF88]. Compatibility [Mac82]. Competition [ME82]. Compliance [Par84]. Composition [CM83, FF82, LUBO85, RH82]. Compositions [Chi89]. Computation [Mun82]. Computer [HM86]. Computerized [CM88]. Concentrated [GM85]. Concurrent [MRL83]. Condition [TMO89, WG83, WS89]. Conditions [DOFJ81]. Confidence [Gus88]. Conflict [CH85]. Connected [Hil89]. Connecticut [BK86, MK88]. Conservation [FL85]. Considerations [AK86, EC85, Rin88]. Constraints [DW81]. Consumption [HM85, PDA82]. Containment [PGR88]. Contemporary [Wri81]. Contents [Dim85, Mir86]. Contest [OC89]. Contribution [RHT88, WGF87]. Control [GB83, MBRG83, PW84, SJ83]. Controlled [Hud85]. Cooling [Bar84]. Cooperative [VJL82, WG83]. Coping [SM86]. Copper [MR86]. Coral [FK87]. Cord [BA88, MS86]. Correlations [Kal87]. Cost [CSP88]. Count [PB87]. Count-Interview [PB87]. Counting [WOH89]. County [KDP86, PDA82]. Court [Cla85]. Cove [HGH82]. Cover [BZM84, CP87, WGF87]. Covers [Har83]. Coves [RH82, WJ86]. CPUE [Jen85]. Crab [And89, Mil83, SRH^{+}83, SHR83, TP84]. Crabs [Bre87, CRS87, TMO89]. Crappie [Bea83, Box87, HBS83, Mit81, SSHP85]. Crappies [Box86a, MG88]. Crawfish [RP83]. Crayfish [HD87, NO88, RB87]. Created [HR81]. Creation [Yea85]. Creek [CRS87, Har83, HB86, KB82, Vin87, WL84, Wor81]. Creel [FNN82, Heg87, PB87, Roh86]. Crest [KPR^{+}89]. Croaker [HW83]. Crop [Bar84, Har83, SB87]. Crops [GEB86, MRL83, WJ86]. Cultch [Eck83]. Culverts [BG89]. Current [BG88]. Curves [Cra87]. Cutthroat [KKM85, SGG86]. Cyanide [Wil84]. Cyclic [KH82]. Cyclical [TP84].

Daily [GS87]. Dakota [HM85, MBRG83, SM82]. Dam [BK86, BK85, BWCQ88, DD89, PH89, BJFF87, FMW88, GS87, JD88]. Dams [GSZ^{+}88]. Data [Ebb87, FNN82, FGM84, SRJ82, Sch84, SS88, TBM84, Wil87b]. Day [BJFF87, DD89, Par89a]. Dealing [BP86b]. Debris [Bry83, MK89]. Decade [Bar89]. Decades [Don87]. Decision [Bai87, Cla85]. Deep [HR81]. Deep-Spawning [HR81]. Deepwater [DRM84]. Definition [Alm87]. Definitions [PB87]. Delphi [Cra87, Zub81]. Delta [SM83]. Demand [HB85a, LD88]. Densities [Hau83]. Density [CJH84, EH86, Hal86, RP83, Ser82a, Ser83, WMG85, WS87, WS89].

Depletion [MK89]. Deployed [LGJR89]. Deposition [GSDK89, HMK89, Ser82b]. Depressurization [FK83]. Deprivation [EH83]. Depth [BJM82, CCW87, HRDM84, SJ85, Ueb86]. Derivation [FF82]. Derived [CA87]. Describing [CM83]. Design [CW85, LRLM84, MWTO89, Rin88, WOH89]. Designing [BNHF88]. Designs [SRH^{+}83]. Detect [CRK88, GC84]. Detecting [PBB88]. Determinant [Hub88]. Determination [BTPM89, CB83, Eri83, MMS^{+}83]. Determine [DJ87, LO88, RC86]. Determined [Par89b, PT84, Tho85, WGS84]. Detonating [BA88, MS86]. Developed [Cra87]. Developing [Eva81]. Development [AFL83, Axo81, BNA83, Ray88]. Device [Dim85, JLZM87]. Devices [Scr88]. Dewatering [BNA83]. Dewaterings [Hol87]. Die-Offs [Mur85]. Diel [Atk89, BP86a, Gis85, RP83, Wal88]. Diet [HS83, SMJR89, SH87]. Differences [BKSM89, LD87, MG88, RC88]. Different [GAD86, HES86, Red86, WE86]. Differential [Mat89]. Differentiating [AGN81]. Differentiation [ST85]. Differing [WT84]. Digeranylamino [GB83]. Dillon [SSB85]. Dimensions [BTB85]. Direct [BG88]. Discards [HL87]. Discrimination [WGW88]. Dispersal [RC89]. Distance [Rei88]. Distribution [Bar84, BMM87, CCW87, CW85, CCS89, CP82, FH82, Hub88, KMR89, LD88, Mos83, Rob86, SB83]. Distributions [FMW88]. Diver [JLZM87]. Diver-Operated [JLZM87]. Diversion [MBRG83]. Diversions [McK87]. Docks [IB84]. Dogfish [MB86]. Domestic [Gow86]. Dorsal [Eri83]. Dover [Ueb86]. Downstream [CM86, Cus85, HS89]. Dredge [GA81, Tho85]. Dredging [Har86]. Drops [Joh87]. Drum [FGM84, Fer86, Mat86]. Drums [FJ87]. Dugout [LRLM84]. Dungeness [Bre87]. Durations [MSG86]. during [MM88, RH82, SHP87, SGO88, WG83]. Dynamics [HW83, ST85].

Each [Ser87]. Early [See87b]. Eastern [KK84, FNW82, MR88, WT84]. Ecological [Cus85]. Economic [And89, Eas82, GP88, Tow85, WT84]. Editorial [Ken89]. Eelgrass [FTCF84]. Eels [PM87]. Effect [Bea83, GH89, Hol87, MZ87, SSHP85, Wal88]. Effectively [Nol87]. Effectiveness [BWCQ88, CRS87, Laa81, NG87, Qui88, SMS88, TG85]. Effects [AN84, BG88, BNA83, BMM81, Cla83, Cob88, Cus85, DD89, EGT^{+}84, FMW88, FK83, FK87, GLC86, GA87, Gil82, GEB86, GA81, Har83, Har86, HMK86, HN82a, HBDZ89, HB86, HEJ87, HP88, JDTK87, KD87, KPR^{+}89, LKL87, MH88, MSB84, MS85b, MWTO89, Mos83, PW87a, PB87, Ray88, Red86, RW88, Rin88, RP83, Say87, SM83, SN88, SKB82, Szt85, Vin87, AH83]. Efficacy [GM87, Gil89]. Efficiencies [Joh87, WT83]. Efficiency [Dim85, JS82, Lyo86, PPB89, SGO88, WMG85]. Efficient [Mor82]. Effort [FB82, JD88, LUBO85, Mor85b, Par89a, Ser82a, Ser83]. Egg [Cam85, Enn85, HS85, HMK89, Ser82b]. Egg-per-Recruit [Cam85]. Eggs [DOFJ81, HP82, Hol87, MBRG83, RW88]. Eight [Don87, LR82]. Electric [PB88]. Electrically [SC87]. Electrofishing [BFB85b, BG88, BKSM89, Ebb87, Hal86, HHHH83, Hud85, JS82, JLZM87, Par89a, RKSC88, Ser82a, Ser83, SH86b, SC87, SC88, WT83]. Electronic [PN83]. Electroshocking [GLC86, GA87]. Elimination [CCS89]. Elk [Bea83]. Embryo [HP82, TB83]. Emergence [FGC86, HP82, Scr88]. Emigration [CK88, KPR^{+}89]. Emphasis [Bru84, Laa81]. Endangered [Mef83]. England [Alm87]. Enhancement [CSP88, HB85b, OB87, PR85, Ray88, TW87, WKC83]. Enrichment [GP88]. Entanglement [BJ84]. Entanglement-Gear [BJ84]. Entrainment [SJT^{+}82, STP83, WBJR87]. Entry [Tow85]. Environment [Hud85]. Environmental [JD88, MM85, SB87]. Equation [Wil89]. Equations [TP84]. Erie [FJ87, IR89, MW86, SJT^{+}82, Szt85, Wol81]. Errata [Ano82a, Ano82b, Ano82c, Ano86, Ano87]. Escambia [Yea85]. Escanaba [Ser82b, Ser87]. Escape [SMS88]. Escapement [PB88, RHT88, SS88]. Established [SWH85]. Establishment [Fra89]. Estimate [HGH82, Mat86, PB88, SRH^{+}83, SHR83]. Estimated [EH86, SB88a]. Estimates [Atk89, Bar84, BR88b, CA87, EC85, PB87, Whi89, WC83, ZF88]. Estimating [FNN82, FB82, GS87, HLS81, HN82b, KM82, MM88, SS88]. Estimation [CJH84, CRK88, DRM84, HS85, Hil89, JRK88, KS89, MM85, PC84]. Estimations [JS82]. Estuarine [Heg85]. Estuary [Bre87, Scr88]. ethanol [GB83]. European [WKC83]. Evaluating [Hau83, KB85, PN83]. Evaluation [AO88, BG89, Box86b, BJ84, BR86, Bul83, CAG81, DA89, DRG82, DP84, ES86, FGM84, GP88, HBS83, HB85b, LAA83, MM85, MS82, Nov87, PM87, POB85, PN88, RP85, RT86, SHP87, SWH85, Tow85, Whi89]. Even [RC88]. Even-Odd [RC88]. Evidence [CB83, DD89, KT84, ME82, NP84]. Examination [ST87]. Example [BP86b, GSO89, Hil89]. Exotic [HD87]. Expenditures [CSH^{+}87, GD86]. Experience [Bar89]. Experiment [CP87, POB85]. Experimental [MWTO89, SMS88]. Experimentally [Tho85]. Exploitation [BG82, Com82, DA89, Enn85, WC85]. Exploited [BR88a, HT89]. Exploits [BOS87, OB87]. Express [ES87]. Extraction [SD87].

F [KH82]. Facilities [Cus85]. Facility [NG87, PM87]. Factors [CRS87, JD88, Kal87, MM85]. Failure [CRK88]. Fall [EOS88, Ser82b]. Falls [CM86]. Farm [BR86, KNR88, KH82, PH85]. Fauna [EGT^{+}84, HN82a]. Feather [MVG83]. Features [SB87]. Fecundity [BR88a, POS87, Ser82b]. Federal [MG86]. Feeding [Ber82, HM85, PB82]. Female [BR86, Gow86, KH82]. Fencing [KB82, Nol87, PW84]. Field [CP87, DP84, HBDZ89, Jen89, PM87]. Field-Collected [HBDZ89]. Fillet [Ueb86]. Fin [BOS87, Box82, CB83, QR82]. Fin-Clipping [BOS87, Box82]. Fin-Ray [CB83]. Financial [HRDM84, Pof85]. Fine [WBJR87, WRH^{+}89]. Fine-Mesh [WBJR87]. Fingerling [Box82, FM87, Hau83, HC88, KP86, Ser82a, SSB85]. Fingerlings [EOS88, Laa81, MWD86]. Fins [HFG85]. First [CB83, KT84, NP84, RH82, WL84]. First-Year [WL84]. Fish [AK86, AS87, BK86, Bar84, BA88, BMJ85, BJ83, BWCQ88, CJH84, CK85, CCW87, Cla83, CM88, CM83, CSH^{+}87, DHBZ89, Dim85, Eva81, FK87, Gab84, Gar82, GH83, Gil82, Gus88, Har86, HN82b, HN82a, HBDZ89, Hol87, JS82, JS83, JSN87, Jac89, Jen89, KS89, KDP86, KP84, Laa81, LR82, LMW87, MCS85, MB81, MBRG83, MMS^{+}83, McK87, MS86, Mil83, MS82, MGS87, MH88, MWTO89, Mos83, MVG83, NT87, NG87, Nie82, PBP83, PCC81, PJ82, Pof85, PDA82, Qui88, RH82, RP85, SSHP85, SA84, SGS88, SJ83, SB83, Tap86, TPM85, TS82, TG85, THHS89, TH86a, Tsu87, WOH89, WGWP84, WJ86]. Fish-food [PJ82]. Fisheries [Bai87, BNHF88, Bru84, Cla83, Eas82, FNW82, HB85a, HWPS84, Hil89, MG86, Mor85a, Mun82, MR88, OC89, Ove85, PBB88, Ray88, RT85, Sch84, Sta81, WKC83, Wri81]. Fisherman [Mil83]. Fishermen [Fer86, GD86, Hea85, LD87, MSG86, Sco83]. Fishery [And89, Axo81, BM85, BMM81, BJ83, BJ84, BP86b, Cam85, CW85, Dew89, DRM84, ES87, Eva81, FB82, GSDK89, HB85a, Hea85, HEJ87, Jen86, Jon82, KMP86, LUBO85, MT87, MSC81, MR86, MHK^{+}87, Nic86, Nie82, NO88, PN83, Qui88, RHT88, SMU84, SHR83, Szt85, Tay81, TP84, Wol81, Yea85, Zub81]. Fishes [BWCQ88, CB83, Gil89, GA81, HC88, HR89, KMR89, KM82, LO88, Lyo86, ME82, MM82, PPB89, Rin88, SH87, THHS89, Wil84]. Fishing [Alc81, Bai87, Bar89, CH85, Cla85, DW81, DLP89, EH86, FNN82, FK87, FB82, GD86, HM86, Heg87, HRDM84, KDP86, LD88, Mac82, Mat83, Nes86, OC89, WT84]. Fishing-Success [FNN82]. Five [Mor85b, Roh87, SB88a]. Fixed [KPR^{+}89]. Fixed-Crest [KPR^{+}89]. Flaming [SH87]. Float [Sch86]. Float-Angler [Sch86]. Flood [SJ83]. Flood-Control [SJ83]. Floodwater [Mit81]. Florida [CSH^{+}87, CCS89, CCP89, IB84, SSHP85, SHB85, SHP87, TP84, WC83, WC85, Yea85, Zal87]. Flounder [HL87, Wal88]. Flow [CA87, FMW88, Gis85, LO88, McK87, MWTO89, SM83]. Flows [Cus85, JSN87]. Floy [Bul83, EC82, FM87, TC82]. Fluctuating [BMJ85, Gis85, JT85, TS82]. Fluctuation [RT85, SKB82]. Fluctuations [PN88]. Fluorescent [EE85, HBDZ89]. Fluoride [DD89]. Flushing [HN82a]. Flyscreen [KP84]. Followed [Box82]. following [SD87, Wor81]. Food [HM85, LJS82, VHJB83, PJ82]. Foods [JSC88]. Forecasting [Jen85]. Foreign [FNW82]. Forest [KH89, MG86]. Forested [PN85]. Forgone [JRK88]. Fork [MVG83, NBH85, Roh86, Roh87]. Formation [CCP89, QR82]. Formulatiug [Mor85a]. Four [BNA83, BK82, DP84, HLS81, HL87, HR89, Kal87, KD87, MB81, SMJR89]. Fraser [Bre87]. Freeze [Box82, FP85]. Freeze-Cauterization [Box82]. Freezing [Say87]. Frequency [FF82, Jac89, RP83]. Fresh [LKL87]. Freshwater [FJ87, Kwa87, Lyo86, MGS87, WKC83]. Fringing [FK87]. Fry [BZM84, Bil86, Dom87, FGC86, HP82, HP88, LM87, MS85a, QD83, Scr88, WL84]. Fullness [KM82]. Fundy [Cam85]. Future [HM86, LD88]. Fyke [MS85a].

Gabion [KN88]. Gabions [Tap86]. Galeorhinus [MSK89]. galeus [MSK89]. Galveston [GD86]. Game [Qui88]. Garrison [MBRG83]. Gastric [LAA83]. Gatewells [BJFF87]. GD [GB83]. GD-174 [GB83]. Gear [BJ84, Dot82, SHR83]. Generalist [SM86]. Generalized [BMJ85, Tay81]. Generation [MKM89]. Genetic [Win89]. Geographical [AGN81]. Georges [BP86b, GSO89, Ove85]. Georgia [CRS87, GH83, HGH82, MSB84, WE86]. Giant [HS85]. Gill [BG82, CCW87, HS83, IR89, LGJR89, MH88, Neg82, Nes86, Qui88, Red86, WMG85, Wil87b]. Gill-Net [HS83, IR89, LGJR89, Nes86, Qui88, Red86, WMG85, Wil87b]. Gilling [SRJ82]. Gizzard [KDP86, TH86a, Wil87a]. Gold [Har86, Tho85]. Gorge [SH87]. Grand [Wal88]. Grass [HM85, Hil86, SB88b]. Gravel [Hof86, TB83]. Grazing [KB82, PW84, PN85, Rin88]. Great [DS88, DLP89, Jac89, MRL83]. Green [BR86, GA87, KH82, LT89]. Greenland [BP86a]. Grinding [Mac88]. Groundfish [MR88]. Grounds [NP87]. Groundwater [GSDK89]. Group [CM83]. Growth [BC87, BK82, CK88, DA89, Eck83, GLC86, GA87, Gow86, HM85, HT89, HP88, JSC88, KPR^{+}89, LD88, LJS82, MSB84, MS85b, MW86, PW87a, PT84, Rin82, SD87, SM82, TC82, WL84, Whi89, WS89]. Guilds [LO88]. Gulf [HL87, Pof85, WC85].

Habitat [And89, BFB85a, BTB85, CSP88, CK88, Cra87, EPNB89, HR81, HMK86, HB86, HR89, JSN87, JSC88, KB82, KN88, KH89, LMW87, LO88, Par89b, PH88, Pit89, PN83, Rin88, SM85, TBM84, VHJB83, WGH87]. Habitation [Zal87]. Habitats [DHBZ89, Llo87, PW84, Rin82]. Habits [LJS82, VHJB83]. Haddock [BP86b, GSO89]. Hake [Alm87]. Half [Har83]. Half-Log [Har83]. Halibut [BP86a]. Handline [LUBO85]. Hard [And89]. Hard-Shell [And89]. Hardening [TMO89]. Harmful [HEJ87]. Harvest [GM85, Har83, JSN87, JT85, KNR88, Mit81, Nes86, NH88, Par82, PW87b, PB87, RM82, Rei87, Roh87, SGM89, SSHP85, SN88, SRH^{+}83]. Harvestability [TMCS81]. Harvestable [CSH^{+}87]. Harvested [Ser87]. Harvesting [CRS87, FTCF84, HEJ87, HRDM84, RP83]. Harvesting/Processing [HRDM84]. Harvests [JSC88, KNR88, SHR83]. Hatchery [BSS82, CM88, DP84, Eck83, EOS88, HLS81, HFG85, HES86, Mor82, PH89, Rei88, RC89, See87a, Vin87]. Hatchery- [PH89]. Hatchery-Raised [See87a]. Hatchery-Reared [Eck83, EOS88, HLS81, RC89, BSS82]. Haul [Chi89]. Haul-Seine [Chi89]. Hawaiian [MT87]. Hazards [PDA82]. Head [BK85]. Headwater [JS83, SJ83]. Health [PDA82]. Heated [GM85]. Helens [MWD86]. Hellgrammite [NO88]. Hellgrammite-Crayfish [NO88]. Henry [Roh86, RT86, Roh87]. Herring [BTPM89, Bul83, CS87, HS85, MEB84, PB81, SS88]. Hidrostal [PM87, RP85]. High [HR89, Hud85, SH86b]. High-Plains [HR89]. High-Voltage [SH86b]. Higher [Eas82]. Highway [BG89]. Histologic [MZ87]. Historical [NP87]. History [AFL83, CAG81, POS87]. Holding [FGC86, Nol87]. Holston [NBH85]. Holyoke [BK86]. Home [Rin82]. Hook [BML^{+}86]. Hook-and-Line [BML^{+}86]. Hooking [Bar85, CC89, Fle87, PPP89, Sch89a, SGG86, Wer88]. Hoop [GH89, PCC81]. Horizontal [Hof86, Nes87]. Hour [Hal86]. Hudson [Dew88, DRWM87]. Huron [NP84, NP87]. Hybrid [DJ87, Gow86, HM85, JDTK87, RT86, Yea85]. Hybrids [BR86, Fra89, GA87, GM85, JSC88, KH82, LJS82, RT86, TMCS81]. Hydrilla [CSH^{+}87]. Hydro [HN82a]. Hydroacoustic [BMM87]. Hydroelectric [BK85, Cus85, GSZ^{+}88, HG86, NG87, PM87, Ray88]. Hydrogen [Mir86].

Ice [Nes86]. Ice-Fishing [Nes86]. Ichthyoplankton [Nes87, WBJR87]. Idaho [BMW86, GA81, KB82, PN85, RC89, Roh86, RT86, Roh87]. Identification [Gar82, MKM89]. Identifying [FP85]. II [AH83]. Illinois [AS87, DJ87, JDTK87, SH88, WGWP84]. Images [BTPM89]. Imaging [MMS^{+}83]. Impact [BM81, Enn85, FTCF84, GEB86, HRDM84, KMP86, MMC87, SJT^{+}82]. Impacts [Car82b, PN85, Sch84, Tho85]. Impingement [SJT^{+}82]. Implanted [FHB87, Tyu88]. Implementation [Dew89, HWPS84]. Implications [Atk89, BP86a, Cob88, FTCF84, MSC81, MR86, MK89, PN88, PJ82, ST85, SS88, TGPS86]. Importance [SMJR89]. Impoundment [Par82, RH82]. Impoundments [BA88, Box86b, Box87, CJH84, Hal86]. Imprinted [HK88]. Improve [KN88, Tap86, Tsu87]. Improved [Rin82, WOH89]. Improving [DA89]. Incentive [MA84]. Inch [CM83, Nov87, Par82, Par84]. Inch-Group [CM83]. Incidence [THHS89]. Incidental [MR88, Qui88]. Inclement [SGO88]. Inclined [MK88, SLA85]. Inclined-Plane [MK88]. Inclined-Screen [SLA85]. Incorporating [TP84]. Increase [Dim85, MA84]. Increasing [KNR88, SGO88]. Increments [BC87]. Incubation [Bam85, Scr88]. Incubators [Dom87]. Index [AK86, Cra87, HR89, KB85, WGH87, Wil87b]. Indicator [Hal86]. Indices [Gus88, Kim88, WGS84, WMG85]. Individual [Dew89]. Induced [Dot82]. Industry [DLP89, Nie82, Tow85]. Ineffectiveness [GB83]. Inexpensive [HMK89, KP84, TPM85]. Infectious [Fli83]. Influence [Box87, CSH^{+}87, GC84, Hea85, MWD86, Mit81, PH88, SC88, SJ83]. Influences [LJ89]. Influencing [JD88]. Information [Nic86]. Ingested [Gar82]. Inhabiting [FK87]. Injections [FP85]. Injuries [SC88]. Injury [RKSC88]. Inland [SH88, Tay81]. Inlet [PB82]. Input [MK89]. Insect [Gis85, HN82a]. Inshore [Dew89]. Instantaneous [GLC86]. Instream [Bam85, CA87, HB85b, HB86, LO88, TW87]. Intake [BJFF87, MCS85]. Integrated [HRDM84, LP82]. Integrity [AK86, AS87]. Intensities [Heg87]. Intensity [EH86]. Interactions [HD87, MS85b, PJ82]. Intermountain [PN89]. International [FL85]. Interpretation [AK86, CJH84]. Interstice [LJ89]. Interstrain [Fra89]. Intertidal [QD83]. Intervals [Gus88]. Interview [PB87]. Intragravel [Hof86, MM88]. Introduced [HS86, WC83]. Introductions [KKM85]. Invertebrate [BNHF88, Mac82]. Invertebrates [GA81, Har86]. Investment [TGPS86]. Iowa [Hil86, HS83, Mit81, Par82, Par84, PW87a]. IPN [Fli83]. Island [Hau83, PH89]. Islands [MT87].

Japanese [MHK^{+}87]. Joaquin [SM83]. Jocassee [Bar85]. John [BJFF87, DD89]. Jordan [JD88]. Juvenile [BB82, BJFF87, BMW86, CSP88, GS87, HMK86, KD87, Mit81, POB85, PC84, RC89, SHS85, Wal88].

Kansas [Bea83, Gab87, WJ86, WS87, Wil87a]. Kaplan [BK85]. Kayak [Roh86]. Kelp [HS85]. Kentucky [BJ84, Axo81, BJ83, BJ84, BKSM89, HT89, JS82, Jon82, SJ83, THHS89]. Kentucky-Tennessee [HT89]. Keystone [CP82]. Kokanee [HR81, STP83]. Kokanees [PH88]. Kootenai [CM86].

Labelling [BC87]. Labrador [BP86a, Bru84]. Lacustrine [POB85]. Ladder [MWTO89]. Lahontan [KKM85]. Lake [Axo81, BG82, BJ84, Bru84, BKSM89, CH85, CSH^{+}87, CCS89, DS88, DOFJ81, DJ87, EC82, EC85, ES87, EPNB89, ES86, EOS88, GP88, HBS83, Hau83, HLS81, HS86, HT89, HMK89, HS83, IR89, JT85, JDTK87, Jon82, Kal87, KS89, KDP86, KKM85, KK84, Lyo86, MSC81, MKM89, MS82, Mos83, MMC87, MW86, Nes86, NP84, NP87, Nov87, PB82, RH82, RT86, Sch89a, ST85, SRJ82, Ser82b, Ser87, SB88a, STP83, SH89, Szt85, TB86, TG85, THHS89, Wag81, BJ84, BMM87, CH85, EC85, EOS88, FJ87, HGH82, KT84, MKM89, Mit81, NP84, Nic84, RM82, Say87, SSHP85, See87a, SJ85, SJT^{+}82, Wag81, WE86, Wol81]. Lakes [Alc81, CCW87, Fra89, Gow86, Hil86, HS86, Kal87, MB89, MGS87, PBB88, SMJR89, SHB85, SHP87, Ser82a, SB88b, TS82, DLP89, Jac89]. Lamprey [Jac89, SH89]. Lampreys [HS89]. Lampricide [See87b]. Lampricides [HS89]. Land [LP82, PN88]. Land-Aquatic [LP82]. Land-Use [PN88]. Landings [LWL89]. Landlocked [NG87, PB81, SMJR89]. Lanier [WE86]. Large [BKJ89, HW83, RKSC88, SLA85, SM87, TPM85, ZF88]. Largemouth [Box82, Box86b, Box87, BML^{+}86, CCS89, CCP89, DPK84, Ebb87, FNN82, FK83, FHB87, Gab87, Hal86, HGH82, Hub88, MSB84, Mir86, MM87, MS85b, Nov87, Par82, PT84, PGR88, Qui89, SM85, SHB85, SHP87, SWH85, SM82, TC82, WGS84, Whi89]. Larvae [Hol87, TH86a]. Larval [TH86a]. Latex [FP85]. Latitudinal [MS85b]. Laurel [Jon82]. Lavage [LAA83]. Lawrence [DRG82]. Leading [EH83]. Length [AGN81, AO88, BML^{+}86, FF82, Gab84, LUBO85, MH88, Nov87, Par82, Par84, Say87, Ueb86, Wil89]. Length-Categorization [Gab84]. Length-Frequency [FF82]. Letter [Sta81]. Level [Bea83, MSB84]. Levels [BMJ85, Kal87, Sch86, TS82]. Lewis [WL84]. Liberty [MSK89]. License [Ret84]. Licensing [Sco83]. Life [See87b, Ser87]. Lifetime [POS87]. Lifts [BK86]. Light [SHS85]. Limit [AO88, BML^{+}86, Enn85, Nov87, Par82, Par84, SGM89]. Limitation [Ret84]. Limitations [Tap86]. Limiting [BWCQ88]. Limits [DRG82]. Line [BML^{+}86]. Linear [Kim88]. Live [CC89, Jen89, MMS^{+}83, Mir86, PM87, PPP89, PGR88, SC87]. Lives [Bar88]. Livestock [PW84]. Living [Hil87]. Load [AH86]. Lobster [Cam85, Enn85, Jen86, VJL82]. Lobsters [Enn83, MT87]. Log [Har83, Kim88]. Log-Linear [Kim88]. Logged [KN88]. Logging [GEB86, HMK86]. Long [DS88, EE85, MB86, Tyu88, Szt85]. Long-Term [DS88, EE85, MB86, Tyu88]. Longfin [SM83]. Longitudinal [BKSM89]. Longline [Hil89]. Loons [Fli83]. Loss [EC82, Kal89]. Losses [SA84, STP83]. Lost [Bre87]. Louis [Alc81]. Louisiana [KPR^{+}89, NT87]. Low [BK85, CSP88, Heg87, McK87]. Low-Cost [CSP88]. Low-Flow [McK87]. Low-Head [BK85]. Lower [BNC^{+}86, PBP83].

m [TH86a]. Macrophytes [WGWP84]. Mad [HK88]. Madison [Sch86, Vin87]. Maine [BJM82, HLS81, HL87, Jen86, SMJR89, Tow85]. Maintain [BTB85]. Making [Bai87]. Male [BR86, Gow86, KH82]. Manage [Mor85a]. Management [Bai87, Bea83, BTPM89, Box86b, Bry83, BWCQ88, BKJ89, Cla85, Cob88, EC85, FTCF84, FL85, FNW82, Gus88, HB85a, Hea85, HG86, Hil87, Hub88, Jen89, Low81, LGJR89, MB89, Mef83, MSK89, Mun82, MK89, PR85, PJ82, POS87, ST85, Sch84, SS88, Tay81, TB86, VJL82, WKC83, Whi89, WW89, WC85, Wri81, Ken89]. Managers [MG86, Zub81]. Managing [CM88, Eas82, Ove85, Rei87]. Manipulation [MSB84]. Manitoban [Eri83]. Many [Mil83]. Mapping [EPNB89]. Maquoketa [PW87a]. Marina [IB84, FTCF84]. Marine [CB83, DW81, FB82, LD88, MSG86, MG86, Sco83]. Mark [Bul83, CRK88, LM87, PC84, ZF88]. Mark-Recapture [Bul83, PC84]. Marked [HBDZ89]. Marker [AGN81]. Marking [Bil86, Box82, CK88, ES86, HC88, KP86, TS82]. Marks [EE85, FP85]. Marsh [KPR^{+}89]. Marshes [NT87, Rog85]. Maryland [SRH^{+}83, SHR83, WT83]. Mask [SGS88]. Mask-Bar [SGS88]. Mass [Box82]. Matched [Gow86]. Material [Nol87]. Mathematical [BG82]. Mature [MMC87]. Maturity [BTPM89, MW86]. Maximum [Bar88]. Mayfly [RC88]. McMichaels [Har83]. McNary [FMW88, GS87]. Meadows [FTCF84]. Means [DA89]. Measure [BM85]. Measurement [SGS88]. Measurements [CM83, LMW87]. Measures [MS82]. Measuring [Jac89, WRH^{+}89]. Mechanical [HBS83]. Mechanism [EH83]. Mechanisms [SMS88]. Medicine [KH89]. Megawatt [BK85]. Members [KDP86]. Menhaden [CW85, Jen85, RC86, SHS85]. Mesh [HRDM84, Red86, RC86, WBJR87, WMG85]. Mesh-Size [RC86]. Metamorphosed [HS89]. Meter [PN83]. Method [BFB85b, CB83, FNN82, GEB86, HMK89, KS89, RC86, TB83, TPM85]. Methods [Box86a, CS87, HP82, Hau83, HLS81, KB85, Nol87, Ove85, PC84]. Metropolitan [MM82]. Mexico [Pof85, WC85]. Michigan [AH83, Bai87, Car82a, CAG81, CH85, CM88, DOFJ81, EC85, EPNB89, Gow86, HAD83, HS89, Laa81, Say87, ST85, See87a, Wag81]. Michipicoten [KK84]. Microfiche [Tsu87]. Microhabitat [WC85]. Microhabitats [BFB85b]. Microscopic [ST87]. Microtaggants [KP86]. Middle [Alm87]. Midwater [EH89]. Midway [MT87]. Migrating [PH89, WOH89]. Migration [HS89, HG86, NG87]. Migratory [BM85, MM85, MR86, Mun82]. Minimum [AO88, BML^{+}86, DRG82, JSN87, Nov87, Par82, Par84]. Minimum-Length [AO88]. Minnesota [Ebb87, OC89]. Minnows [CC89]. Miramichi [CR86]. Mississippi [PBP83, PCC81, Pit89, SGM89, SB83, BR86, RT85, TG85]. Missouri [HN82b, Alc81, ES87, New89, Nov87, RR89]. Mitigating [EGT^{+}84]. Mitigation [CK85]. Mixed [MKM89]. Mixed-Stock [MKM89]. mm [AO88, BML^{+}86]. Model [BMJ85, BG82, Cam85, CA87, LWL89, PJ82, SS88, SB88b, WGH87, WGWP84]. Models [HWPS84, HR89, Jen86, Kim88, KH89]. Modified [Neg82, SLA85, WGH87]. Monofilament [Joh87]. Montana [BG89, CM86, Sch86, Tho85, Vin87, Wor81]. Moose [TB86]. Morpholine [HK88]. Morpholine-Imprinted [HK88]. Mortalities [Dot82]. Mortality [Bar85, BK85, Bre87, CC89, CK88, DRWM87, DRM^{+}89, EH83, Fle87, Hil86, HBDZ89, Hud85, Jew86, KPR^{+}89, Low81, MSB84, PPP89, Sch89a, SGG86, SHP87, SH86b, SD87, Wer88]. Mothership [MHK^{+}87]. Motivations [CH88, DW81]. Motive [LD87]. Mottled [BG88]. Mount [MWD86]. Mountain [Don87]. Mountains [MRL83, Rin85, PR85]. Movement [BWCQ88, CM86, Git86, MK88, Rin82, WC83, WC85]. Movements [Atk89, MB86, TP84]. Multi [SHR83]. Multi-gear [SHR83]. Multiattribute [BNHF88]. Multifilament [Joh87]. Multiple [BML^{+}86]. Multispecies [Ove85]. Musculature [FHB87]. Muskellunge [Axo81, Dom87]. Mysis [MB89].

Names [Ken88]. Natal [Rei88]. National [Don87, Kal87, KH89, MRL83, SGG86]. Native [WC83]. Natural [Car82a, Hau83, NP84, PBP83, WS87]. Naturally [MKM89, PH89, RC89]. Navigation [Hol87]. Navigation-Related [Hol87]. near [GM85]. Nebraska [HN82a, New89]. Necrosis [Fli83]. Neosho [Com82]. Net [BJFF87, BJ83, CCW87, Chi89, HS83, IR89, LGJR89, MH88, Nes86, Nes87, Qui88, Red86, RHT88, SRJ82, STP83, TH86a, WMG85, Wil87b]. Net-Pens [RHT88]. Nets [BG82, DHBZ89, GH89, LR82, MS85a, MH88, Neg82, PCC81]. Nevada [WW89]. Newcastle [HD87]. Newfoundland [BOS87, Enn83, Enn85, OB87, Red86]. Newly [SM82]. Night [Par89a]. Nine [BMM81]. Niobrara [HN82a]. Nongame [HR89, ME82]. Nonlethal [BTPM89]. Nonvegetated [DHBZ89]. Nonwooded [WJ86]. North [Lyo86, MBRG83, MM85, BM85, HEJ87, Ken88, MVG83, Zal87]. North-Central [Zal87]. North-Temperate [Lyo86]. Northeast [FNW82]. Northeastern [HS86, ST85]. Northern [BR88b, BG82, DRG82, FMW88, Mar87, MMC87, Wil89, WS89, EPNB89, MSC81, Ser82a]. Northwest [HRDM84]. Northwestern [Sch89a]. Norway [HG86]. Norwegian [Heg87]. Notes [KNR88]. Number [Mat86]. Numbers [BKJ89, HS85, Mor85b].

Nursery [Bry83]. Nutritional [EH83].

Oahe [BMM87]. Observations [Fli83, HR81, NP87]. Observer [FNW82]. Obtained [Whi89]. Obtaining [TBM84]. Occurrence [BNC^{+}86, BJ83, LMW87]. Ocean [Mac82, AFL83, FNW82, WG83]. Oconee [HGH82]. Ocqueoc [HS89]. Odd [Nic84, RC88]. Odd-Year [Nic84]. O’Dell [Vin87]. off [BP86a, Low81, MSK89, WG83]. Offs [Mur85]. Offshore [GD86]. Ohio [CJH84, CK85, Hal86, IR89, Nie82]. Okeechobee [SSHP85]. Oklahoma [Box86a, Box87, Com82, CP82]. Old [KK84, Nic84]. Omak [KKM85]. Onalaska [RM82]. One [BNA83]. One-Time [BNA83]. Ontario [BTB85, EOS88, HS86, MKM89, MMC87, Sch89a, SJ85, TB86]. Open [NT87, Sta81]. Opening [Nes87]. Openings [TH86a]. Operated [JLZM87]. Operation [HRDM84, LRLM84]. Operational [HHHH83, Pof85]. Opercula [BC87]. opilio [TMO89]. Optimal [CW85]. Optimization [Ove85]. Orange [CSH^{+}87]. Orders [Cla85]. Oregon [HB85b, HB86, Mor82, Mor85b, WG83]. Origin [Mat83]. Origins [MKM89, MHK^{+}87, MR88]. Orkla [HG86]. Other [CRS87, MSC81]. Otolith [Box86a, CS87, SH86a]. Otoliths [CCP89, Eri83, LM87, Mac88, PT84, Sch89b, ST87, TPM85]. Otter [DRM84, Ove85, Rog85]. Outcome [Don87]. Outlet [GM85]. Overlap [SM85, SH87]. Overview [KRJ86]. Overwintering [SH88]. Oxygen [DJ87]. Oxytetracycline [Bil86]. Oysters [Eck83]. Ozark [MR87, RR89].

Pacific [FNW82, AFL83, BTPM89, CS87, FF82, HS85, SS88, WG83, Wri81]. Paddlefish [Com82, Cra87, HT89, SGM89]. Pamlico [MM85]. Pancreatic [Fli83]. Parental [HM86, MKM89, MM87]. Parentals [KH82]. Park [Kal87, MRL83, SGG86]. Parks [Don87]. Part [AH83, HAD83]. Partial [DRM84]. Participation [LD87]. Passage [BG89, DD89, GS87, MWTO89, Tap86, TH86a]. Passing [BK85]. past [CM86, PH89]. Patterns [Gis85, GD86, HS83]. Pectoral [HFG85, QR82, SD87]. Penaeus [KPR^{+}89]. Pennsylvania [SA84, Har83, SBB88]. Pens [RHT88]. Percent [Ser87]. Perceptions [Sch86]. Perch [AFL83, HS83, Kal87, SJT^{+}82, WG83]. Performance [BB89, HFG85, HB83, HRDM84]. Period [MM88, OC89]. Periodicity [Zal87]. Permanganate [Heg85]. Peroxide [Mir86]. Persistence [GAD86]. Perspective [WKC83]. PHABSIM [CA87]. Phases [BNA83]. Phoenix [MM82]. Physical [HR89, MR87]. Physicochemical [LMW87]. Piedmont [WT83]. Pigment [EE85, HBDZ89]. Pike [DRG82, Mar87, MMC87, Wil89, WS89]. Pink [Git86, KN88, KMP86, KK84, Kwa87, Mur85, Nic86, Nic84, Nol87, Rob86]. Pipe [Hof86]. Placement [TW87]. Plains [HR89]. Plane [MK88]. Planning [HB85a]. Plant [Gil82]. Planted [Bar85, Hil86, NP84, Roh87, See87a, Wag81]. Planting [HP82]. Plantings [Gow86, SSB85]. Plants [KMR89]. Plugged [NT87]. Point [MSB84, Szt85]. Poisoning [Wor81]. Policy [TGPS86]. Pollock [WT84]. Polluted [PDA82].

Polychlorinated [BMM81]. Pond [Gab87, GB83, KNR88, KH82]. Ponds [BHN82, BR86, CK88, CRK88, GH83, GAD86, HM85, LJS82, MS85b, PH85, SM82, WGWP84]. Pony [ES87]. Pool [Pit89, SB83]. Pools [EGT^{+}84]. Pop [DHBZ89]. Population [AH83, AN84, Bea83, BR88b, Ber82, CRK88, DA89, Ebb87, EC85, EH86, Enn83, Enn85, Gus88, HBS83, HT89, HW83, JS82, JS83, JT85, KS89, KH82, Laa81, LT89, MSC81, MR87, NBH85, PC84, RR89, ST85, SSHP85, WGS84, WS87, WC83]. Populations [BR88a, Box87, BM81, CJH84, CK85, Cob88, CM83, CSH^{+}87, DS88, FK87, Fra89, GH83, HLS81, Hil89, HB86, Mac82, MCS85, MS82, MGS87, MVG83, PBP83, PN88, Rin85, SM87, SJ83, WS89, Wor81, ZF88]. Porcupine [Fle87]. Portable [McK87, PB88]. Portions [MEB84]. Possible [EH83, Fli83]. Post [BB89, HFG85]. Post-Stocking [HFG85, BB89]. Posteruption [MWD86]. Postrelease [SHP87]. Pot [KB85]. Potassium [Heg85]. Potential [Bru84, Cla83, DW81, Eva81, MMS^{+}83, PDA82, RT86, Ser82b]. Potentialities [HB85a]. Potomac [KMR89]. Pots [SRJ82]. Pound [Chi89]. Pound-Net [Chi89]. Powell [PB82]. Practices [GEB86]. Pragmatic [BP86b]. Prairie [Mos83]. Pre [Mur85]. Pre-Spawn [Mur85]. Precambrian [Fra89]. Precision [FGM84, SB88a]. Predation [BZM84, Mat89]. Predator [MS85b]. Predators [Fli83, LJ89, PH85]. Predict [SRJ82]. Predicting [BMJ85, MR87]. Prediction [EH86, Nes86]. Preference [KB82, TBM84]. Preferences [BJM82, DW81, DJ87, Hea85, JSC88, VHJB83]. Preferred [SHS85]. Preliminary [BC87, MB86]. Premolt [CRS87]. Preparation [ST87]. Prerelease [SHP87]. Previous [Nes86]. Prey [Gar82, MS85b, PB81]. Prints [Tsu87]. Probability [SJ85]. Probability-of-Use [SJ85]. Problem [MCS85]. Problems [HB85a]. Procedure [Mac88]. Procedures [Ken89]. Processing [HRDM84]. Produce [KH82]. Produced [MKM89, PH89]. Production [Enn85, Jen86, JRK88, MMC87, NBH85, RC88, SB87, WGWP84]. Productivity [MS82]. Professional [MG86]. Professionals [Sta81]. Profiles [Mor85a]. Program [FNW82, MA84]. Programs [BNHF88]. Project [Car82b]. Projecting [LD88]. Projects [TW87]. Property [Eas82, TGPS86]. Proportional [CJH84, WS89]. Proportions [PH89]. Proposed [MB89, Wil89]. Provide [Wil87b]. Public [KDP86]. Puget [RHT88, Cla85]. Pulse [SC88]. Pump [PM87, RP85]. Pumping [GSDK89]. Purse [BKJ89, RC86]. Push [Rog85]. Push-Otter [Rog85]. Put [GSDK89]. Put-and-Take [GSDK89].

Quahogs [Mac82]. Quality [Alc81, HN82a, Hof86, Llo87, Nes86]. Quality-Size [Nes86]. Quantifying [BFB85a]. Quantitative [BM85, BFB85b, HMK89]. Queen [AFL83]. Quickly [LGJR89]. Quotas [Dew89].

R [BMM81, BHN82]. Radio [Mar87]. Radio-Tagged [Mar87].

Radiotelemetry [Par89b, TBM84]. Radiotracking [DJ87]. Railroad [WW89]. Rainbow [BB89, Bar85, BMM81, BK82, BMM87, CM86, DS88, DP84, FP85, GM87, HFG85, HD87, HB83, Hud85, Jon82, MRL83, Mor82, NBH85, RKSC88, Roh87, SMJR89, SC88, SSB85, SH89, Vin87, Wor81]. Raised [See87a]. Raising [Dom87]. Random [Atk89, Heg87, SRH^{+}83]. Randomized [SHR83]. Range [Rin82]. Rapid [BTPM89, SH86a]. Rapidly [Cus85]. Rare [Rin82]. Rate [Enn85, Mat83, Rei88]. Rates [BKSM89, EH86, IR89, Low81, Mor85b, Par82, Qui89, SRJ82]. Rathbun [Mit81]. Ratio [WS87]. Rationale [SGM89]. Ray [CB83]. Rays [QR82]. Razorback [LT89, Tyu88]. Reaches [Roh87]. Reared [Eck83, EOS88, HLS81, RHT88, RC89, BSS82]. Rearing [HB85b, POB85]. Recapture [Bul83, BML^{+}86, CRK88, PC84, Qui89, ZF88]. Recently [HS89]. Reciprocal [TMCS81]. Reclamation [Car82b]. Reconstruction [AFL83]. Record [Kwa87, MSK89]. Recovery [Rei88, Wor81]. Recreational [CH88, Cla83, FB82, LD88, MT87, Sco83]. Recruit [Cam85, DRM84, GSO89, Nic86]. Recruitment [CR86, DRM84, GC84, HM86, MM87, WS87, Wil87a, Wil87b]. Red [FGM84, Fer86, Mat86]. Redbelly [BZM84]. Redds [BNA83]. Redear [KH82]. Redfish [Atk89, DRM84, HRDM84]. Reduce [STP83]. Reducing [AH83, DRM^{+}89, Qui88]. Reduction [HBS83]. Reductions [Mor85b, WBJR87]. Reef [FK87, Pof85]. Reef-Fish [Pof85]. Reefs [IB84]. Reference [Nie82]. Refuge [WW89]. Regime [HM86]. Regimes [Hea85, WT84]. Regulated [Gis85, Kal87]. Regulation [AN84, MR86]. Regulations [Bai87, CAG81]. Rehabilitation [AFL83, CH85, MW86]. Reinforcement [KD87]. Related [Hol87, LMW87, RC88, SB87]. Relating [TB83]. Relation [Git86, HG86, JSN87, MM87, Mor85b, Rei88]. Relations [HR89, IR89, WS89]. Relationship [AGN81, CR86, DPK84, PN89, Ser82a, Ser83, Ser87, WGWP84]. Relationships [WS87]. Relative [Kim88, LR82, SB83, TMCS81, WT83]. Relay [Eas82]. Release [AN84, Bar89, Cla83, HP88, SGG86]. Released [FJ87, HES86, PGR88, Qui89, Wer88]. Relevance [MM85, Mun82]. Reliability [QR82]. relicta [MB89]. Removal [Gab87, KS89, KDP86, Mir86, MRL83, PC84, Sch89b, SH86a]. Removing [Dim85, HAD83]. Renovation [Mef83]. Repeated [GLC86, GA87]. Reported [MSG86]. Reproducing [Fra89]. Reproduction [Car82a, HS86, NP84, PB81, Wag81, Wil87a]. Required [BTB85]. Requirements [LO88, PBB88]. Research [PR85, Rin88]. Reservoir [BB89, Bar84, Bar85, Bea83, BR88b, Bru84, BWCQ88, CP82, HD87, HB83, JS83, ME82, MSB84, PT84, PJ82, SM85, SA84, SKB82, SSB85, SJ83, VHJB83, SH87]. Reservoirs [BMJ85, DPK84, MB89, PB81, WJ86, WS87, Wil87a]. Resident [CP87]. Residual [TG85]. Residues [BMM81]. Resource [BP86a, Hil87, MG86, Sch84, TGPS86, Wal88, WG83]. Resources [VJL82]. Response [KDP86]. Responses [CK85, Gab87, MVG83]. Rest [PN85].

Rest-Rotation [PN85]. Restoration [And89]. Restore [GSDK89]. Restricted [Qui88, Tow85]. Results [BC87, MB86, Par82, WG83]. Retention [DRWM87, EE85, FHB87, Mat83, Tyu88]. Return [KK84, RP85, WE86]. Returns [Eas82, FGM84, HK88, HES86, MA84, SSB85]. Revetted [PBP83]. Review [Bar89, BM81, Cus85, WGW88, Wil84]. Rewards [WE86]. Riffles [EGT^{+}84]. Rights [Eas82]. Riparian [BTB85, KB82, MWD86, PN83, PW84, PR85, WGF87]. Riparian-Fishery [PN83]. Riprap [KD87]. Risk [BP86b]. River [AO88, Bai87, BOS87, Bre87, CR86, Com82, HS89, HK88, HN82a, Jon82, KK84, LT89, MEB84, MVG83, Mun82, NBH85, New89, OB87, PW87a, Par89b, PB82, PH89, Ray88, Roh86, SGG86, Sch86, SBB88, SM83, TB86, Win89, WC83, WC85, Yea85, BK86, BR88b, BNC^{+}86, CH88, DD89, Dew88, DRG82, DRWM87, FMW88, GS87, HN82b, HG86, KMR89, MK88, MR86, Nie82, NO88, PB88, PBP83, PCC81, Pit89, RT85, SGM89, SB83, Vin87]. Riverine [BFB85b]. Rivers [BJM82, Heg87, SLA85, TH86b, WOH89]. Rock [MR87, Nov87, PH89]. Rocky [PR85]. Role [Bar84, Bry83, FNW82]. Roles [SM86]. Rotation [PN85]. Rotenone [BA88, GAD86, Heg85, JS82]. Rough [BWCQ88]. Run [Axo81, BKSM89, Nic84, Zal87].

Sable [Bai87]. Sablefish [KB85, SMS88]. Sacramento [SM83]. Sagitta [Sch89b]. Sahnonid [Gil89]. Sale [NH88]. Salmon [Bam85, BSS82, BNA83, BJM82, Bil86, BB82, BOS87, BMW86, Car82a, CSP88, CR86, Cla85, DD89, EE85, FL85, GSZ^{+}88, GP88, HK88, HES86, HG86, KN88, KT84, KK84, Kwa87, MM88, MWD86, MK88, MR86, MS85a, MZ87, Mun82, Mur85, MHK^{+}87, MR88, NG87, Nic84, Nol87, OB87, PB88, POB85, PC84, QD83, Ray88, Red86, Rei87, Rei88, RW88, RHT88, RC89, SMJR89, SJ85, SM83, Wer88, Win89, Wri81]. Salmonid [BM81, Bry83, FGC86, GEB86, Hof86, HB85b, HB86, Llo87, PN89, Scr88, SLA85, TB83, TH86b]. Salmonids [BJFF87, GS87, HMK86, KD87, MWTO89, SH86a]. Saltwater [LD87]. Sample [PBB88, SRH^{+}83, SHR83]. Samples [Whi89]. Sampling [AK86, BFB85b, BA88, BNHF88, CW85, Heg85, HGH82, Hof86, JS82, KP84, MCS85, MS86, MS85a, Nes86, SGO88, Wil84]. San [SM83]. Sanctuary [Szt85, Git86, KRJ86, KMP86, Rob86]. Sand [AH83, AH86, HAD83]. Sanger [JSC88]. Santee [MEB84]. sapidus [CRS87]. Sauger [LJS82]. Saugeye [LJS82]. Scale [Box86a, CS87, Gar82, NG87, Whi89]. Scales [Eri83, Jen89, Tsu87]. Scallop [FTCF84, HEJ87]. Scan [EPNB89]. School [MSK89]. Screen [McK87, SLA85]. Screens [MCS85, WBJR87]. Sculpin [BG88]. Sea [Cla85, HS89, Jac89, Low81, SH89, FNW82, HW83, MR88, WT84]. Seagrass [And89]. Seasonal [CP82, GH89, Par89b]. Seatrout [Fer86]. Second [MKM89]. Secour [Eck83]. Section [Mac88]. Sectioning [TPM85]. Sediment [AH83, BB82, HAD83, RW88, WRH^{+}89]. Seeding [SKB82].

Segment [SGG86]. Seine [Chi89, Lyo86, PPB89]. Seines [DHBZ89, DRM^{+}89, FJ87, RC86, WT83]. Seining [BKJ89]. Selected [MB81, MRL83, RH82]. Selection [PH88]. Selective [GB83]. Selectivity [BR88b, LR82, MCS85, RC86]. Self [McK87, NBH85]. Self-Cleaning [McK87]. Self-Sustaining [NBH85]. Sensitivity [BMM81, MB81]. Sequential [KNR88]. Series [Jen85]. Set [Eck83, MH88, MSK89]. Seven [Lyo86, Mat83, Hau83]. Several [Ser87]. Sex [BTPM89, MMS^{+}83]. Shad [BK86, BK85, BKJ89, KDP86, LM87, MWTO89, SM83, TH86a, Wil87a]. Shade [LJ89]. Shallow [Rog85]. Shape [SC88]. Shark [MSK89]. Shell [And89, TMO89, Tow85]. Shelled [TMO89]. Shellfish [Eas82]. Shield [Fra89]. Shielded [SC87]. Shock [SBB88]. Shore [FJ87]. Shoreline [HGH82]. Short [Bul83]. Should [Mil83]. Shown [OC89]. Shrimp [BM85, Git86, Jew86, KMP86, KPR^{+}89, LWL89, MM85, Nic86, Rob86]. Side [EPNB89, MWD86]. Side-Scan [EPNB89]. Silver [Alm87, FH82]. Silverside [SH88]. Similar [HES86]. Simple [EH89, Mac88, Mor85a, Sch89b, Tsu87, WGWP84]. Simulated [BMW86]. Simulation [JT85]. Simulations [HM86]. Simulator [Tay81]. Site [PH88]. Six [BG89, CRS87, Gow86]. Sixteenmile [Wor81]. Size [BR88b, Box87, CM83, DRG82, Enn85, FGM84, Gus88, HFG85, HP88, HRDM84, JS82, Kwa87, LR82, LT89, LJ89, MCS85, MGS87, MM87, Nes86, Par89a, PBB88, RW88, Roh87, RC86, SS88, Ser87, SN88, TB83, Vin87, WMG85]. Size-Classes [RW88]. Size-Selectivity [MCS85]. Sizes [HES86, Red86, SWH85]. Slat [PW87b]. Sling [TW87]. Slot [SGM89]. Small [AS87, Box86b, Box87, CK85, CCW87, EH89, Gow86, GA81, HBS83, KS89, KN88, KH89, KP84, LRLM84, Mos83, MMC87, NG87, PBB88, PPB89, PC84, Rog85, SB87, ST87, Tho85, TH86b, WGF87]. Small-Scale [NG87]. Smallmouth [AO88, BR88b, CC89, Hub88, MSC81, MR87, Par84, PW87a, Par89a, RR89, Szt85, WGS84]. Smallwood [Bru84]. Smelt [BMM87, SMJR89, SM83]. Smoky [MRL83]. Smolt [TH86b]. Smolting [See87a]. Smolts [BSS82, BOS87, GSZ^{+}88, HG86, MK88, MS85a, NG87, PH89, SLA85]. Snake [Roh86]. Snorkel [ZF88]. Snow [TMO89]. Social [Car82b]. Sockeye [GP88, MR86]. Sodium [Wil84]. Soft [TMO89, Tow85]. Soft-Shell [Tow85]. Soft-Shelled [TMO89]. Sole [Ueb86]. Some [Car82a, KDP86, MM85, PBB88, Sch84]. Sonar [EPNB89]. Sonoran [Mef83]. Sound [AFL83, MM85, RHT88, Cla85]. Sources [Chi89]. South [Bar85, LWL89, HM85, LUBO85, NBH85, Pof85, PB81, SMU84, SM82]. Southeast [Low81]. Southeastern [DOFJ81, MS85a, Mur85, WGH87]. Southern [HM86, BTB85, IB84, MSK89, RHT88, SH88]. Southwestern [Rin82]. Spatially [Hil89]. Spawn [Mur85, SS88]. Spawned [RC89]. Spawning [BJM82, BMJ85, DOFJ81, EPNB89, GSO89, GC84, HR81, Hof86, HB85b, KN88, KT84, MEB84, NP87, PH88, PB82, Pit89, RHT88, TB83]. Spawns [HS85]. Special [Bru84, Laa81, Nie82]. Specialist [SM86]. Specializations [CH88]. Species [HL87, LR82, MB81, Mat89, RH82, SKB82, Vin87]. Spencer [HN82a]. Spinal [SC88]. Spine [SD87]. Spines [Eri83]. Spinners [CC89]. Spiny [MT87, MB86, VJL82]. Sport [BJ83, CSH^{+}87, HB85a, JSN87, LD87, MSG86, Nie82, OC89, SSHP85, THHS89, WGWP84]. Sport-Boat [MSG86]. Sport-Fish [CSH^{+}87, JSN87]. Sport-Fisheries [OC89]. Sportfishing [CH85]. Spot [Jew86]. Spotted [Fer86, LMW87, Nov87]. Spring [EOS88, EE85, Jen89, KT84, Ray88, Zal87, DJ87, JDTK87]. Springbrook [Mef83]. Springfish [WW89]. Squawfish [BR88b, BM81, FMW88, TBM84, Tyu88]. St [MWD86]. St. [Alc81, DRG82]. Stability [JS83]. Stable [Gis85]. Stage [CW85]. Stages [See87b]. Stampede [ME82]. Standard [Llo87, Wil89]. Standing [Bar84, Enn83, GEB86, Har83, Hil86, MRL83, SB87, WJ86]. States [FNW82, Low81, Ret84]. Station [HG86]. Statistical [CRK88]. Statistics [BR88b, Ebb87, EC85, Laa81, Mun82, PBB88]. Status [FH82, GH83, LT89]. Steelhead [EE85, HP88, PH89, Ray88, RW88, See87a, SJ85, WL84]. Stenothermal [Zal87]. Sterile [RT86]. Stock [Alm87, AFL83, CJH84, CR86, Enn83, GSO89, GC84, MKM89, MM87, MW86, MHK^{+}87, MR88, PH85, ST85, SS88, TP84, WGW88, WKC83, WMG85, WS89]. Stock-Recruitment [CR86]. Stocked [EOS88, EH83, HD87, HP88, Mat89, Mor85b, PB81, SM82, WL84]. Stocking [Bar85, Box86b, BR86, Don87, EH83, Fra89, GH83, Hau83, HFG85, JDTK87, KH82, Laa81, Mor82, SWH85, SN88, SB88b, Vin87, WS87, BB89]. Stocks [AGN81, EC85, Eva81, Gab84, Hil86, Jac89, MR86, Win89]. Stomach [Dim85, KM82, LAA83, Mir86]. Storage [Mit81, WS87]. Story [Jon82]. Strain [Gow86, Mor82]. Strains [BB89, BK82, CK88, DP84, HB83, RT86]. Strategies [AFL83, GP88, JT85, Low81, POS87, VJL82]. Stratified [Atk89, Heg87, HS83]. Stream [AH83, AH86, AK86, AS87, BFB85a, BWCQ88, Car82b, CP87, EGT^{+}84, EH83, GSDK89, Gis85, HAD83, HB85b, JS83, MR87, MS85a, PC84, PN85, PR85, PN89, RR89, Rei88, RC89, Rin88, SM87, Tho85, Wil84]. Stream-Fish [AK86]. Stream-Resident [CP87]. Stream-Stocked [EH83]. Streambed [Scr88]. Streamer [Jew86]. Streamflows [PW87a]. Streams [BTB85, Bry83, Car82a, CK85, CAG81, CA87, FGC86, GA81, HAD83, Har86, HMK86, HR89, HP88, KD87, KH89, MWD86, MRL83, Mor82, Mor85b, MK89, Par84, PH88, PW84, SB87, SJ85, WGF87, WRH^{+}89, WT83, ZF88]. Streamside [MK89]. Strength [RC88]. Stress [Dot82]. Striped [CP82, Dew88, DJ87, DRWM87, DRM^{+}89, GM85, JDTK87, KP86, PH85, PB82, POS87, VHJB83, WGW88, WC83, Yea85]. Strips [BTB85]. Structure [Box87, Gus88, MGS87, MM87, NG87, Par89a, WGS84]. Structured [Bai87]. Structures [CK85, HB85b, HB86, LJ89, SB88a]. Studies [BFB85a, Bul83, CA87, KRJ86, SWH85]. Study [GD86, HWPS84, MK88]. Studying [CCW87]. Stumps [TG85]. Stunted [DA89]. Sturgeon [SMU84, TB86, WC85]. Styles [MG86]. Submandibular [FP85]. Submerged [DPK84, SC87]. Submersed [CCS89, KMR89].

Subsequent [Mit81]. Substrate [BFB85a, DOFJ81, SJ85]. Subyearling [HES86, SJ85]. Success [BMJ85, FNN82, Hau83, Hea85, Mos83, See87a, WS87]. Successful [HS86, NP84]. Sucker [LT89, Tyu88]. Suckers [FM87, QR82]. Suction [GA81, Har86, Tho85]. Suitability [Cra87, HR89, JSN87, LMW87, PB81, WGH87]. Suitable [MB86]. Summer [BMW86, EE85, JD88, Ray88]. Summit [KB82]. Sunfish [BR86, GA87, KH82]. Superior [KT84, KK84, Nic84]. Supplemental [Ber82, Box86b, KNR88]. Supplied [Ebb87]. Surface [Hil89]. Surplus [Gab87, Jen86]. Survey [CW85, HB85a, PB87, SS88, SRH^{+}83, WG83]. Surveys [Heg87, KB85, Roh86, SGO88, SHR83]. Survival [BSS82, BK82, CK88, Eck83, FJ87, Gow86, HM85, HP82, HGH82, HP88, JSC88, LJS82, MM88, MWD86, MK88, PGR88, RW88, SHB85, Scr88, Ser82b, SM82, SN88, SH89, TB83, TC82, WL84]. Susceptibility [GSZ^{+}88, Mat89, SBB88]. Suspended [BB82, Gil82]. Susquehanna [SBB88]. Sustainable [Bar88]. Sustaining [NBH85]. Swim [Dom87]. Swim-Up [Dom87]. Symposium [Bar89]. Synopsis [KMP86]. System [CM88, Gab84, LP82, LGJR89, Neg82, Nes87, Par89b, RP85, MEB84, SM83]. Systems [GSZ^{+}88, HHHH83, Mat83].

Table [Nov87]. Tag [DRWM87, MA84, MB86]. Tagged [FGM84, Mar87, MSK89, TMO89, TC82]. Tagging [DRWM87, FM87, MZ87]. Tags [Bul83, EC82, ES86, FHB87, FM87, HC88, Jew86, Kal89, KP86, TC82, WE86]. Tailwater [JD88, JS83, JSN87, SJ83]. Tailwaters [JS82]. Take [GSDK89]. Tank [SC87]. Tasmania [DS88]. Tax [TGPS86]. Taxonomic [Ken88]. Technique [CSP88, Cra87, HAD83, LD88, SH86a, Tsu87, WRH^{+}89, Zub81]. Techniques [Bam85, HEJ87, WGW88]. Temperate [Lyo86]. Temperature [DJ87, HBDZ89]. Temperatures [GAD86]. Tennessee [HT89, THHS89, VHJB83]. Term [Bul83, DS88, EE85, MB86, Tyu88]. Test [CA87, HR89]. Testing [HEJ87, KH89]. Tests [CRK88]. Tetracycline [BC87, LM87]. Texas [DPK84, Fer86, Mat83, Mat86, SM85]. TFM [See87b]. Their [DW81, Mat83, Mun82, PB87, PN88, TMCS81, ST85]. Them [QR82]. Theory [BNHF88]. Thermally [PT84]. These [Atk89]. Threadfin [BKJ89, TH86a]. Threatened [WW89]. Three [BB89, Bam85, DP84, Eck83, Heg87, HHHH83, HB83, Mat89, RB87]. Three-year [DP84]. Tidal [CRS87]. Tilapia [BZM84, KP86, KNR88, SBB88, Zal87]. Time [BNA83, CCP89, HP88, Jen85, MH88, MSK89, PB82]. Time-at-Liberty [MSK89]. Times [HES86]. Timing [BM85, MM85, MR86, Mun82]. Tobe [HB86]. Tolerance [BNA83]. Tomah [HLS81]. Tool [Box86b, MMS^{+}83, PN83, Zub81]. Topminnow [Mef83]. Tortugas [Git86, KRJ86, KMP86, Nic86, Rob86]. Total [BML^{+}86, SHP87]. Total-Length [BML^{+}86]. Tournament [FNN82, LD87, SHB85].

Tournament-Caught [SHB85]. Tournaments [SHP87]. Tow [Nes87]. Tow-Net [Nes87]. Toxicants [Gil82, MB81, MVG83]. Toxicity [See87b]. Traditional [DOFJ81]. Transfer [PM87]. Transferable [Dew89]. Transferred [Rei88]. Transmitters [Fli83, Tyu88]. Transportation [CM88]. Trap [CRS87, FK87, FGC86, KP84, LR82, QD83, RP83, SRJ82, SLA85, TH86b, WOH89]. Trap-Net [SRJ82]. Trapping [Nol87, RB87]. Traps [Bre87, MK88, Mil83, PW87b, SMS88]. Trawl [Atk89, BP86a, DRM84, EH89, Ove85, Rog85, Wal88, WG83]. Trawler [HL87]. Trawling [LRLM84]. Trawls [DRM^{+}89]. Treatment [BHN82, HS89]. Treatments [MVG83]. Trends [Bea83, OC89]. Trials [GB83]. Tributaries [BNC^{+}86, NO88, RT85]. Tributary [BWCQ88, See87a, SJ85]. Trip [MSG86]. Triploid [RT86]. Trollers [Wer88]. Trotlines [Joh87]. Trout [AH83, AH86, AN84, BB89, Bai87, BTB85, Bar85, BG89, BMM81, BNC^{+}86, BK82, CM86, CAG81, CH85, CA87, CK88, CRK88, CP87, DS88, Don87, DA89, DOFJ81, Dot82, DP84, EPNB89, ES86, EOS88, EH86, EH83, FP85, Fli83, Fra89, GSDK89, GLC86, GM87, Gow86, HAD83, HP82, Har83, HLS81, HFG85, HD87, HS86, HMK89, HB83, Hud85, Jon82, KB82, Ken88, KH89, KKM85, ME82, MKM89, Mat89, MRL83, Mor82, Mor85a, Mor85b, Nes86, NP84, NP87, NBH85, PBB88, PN88, RKSC88, Rin82, Rin85, RT86, Roh87, SB87, SGG86, SC88, SB88a, SM87, SSB85, SH89, Vin87, Wag81, WGF87, WGH87, Wor81, ZF88]. Tubes [Dim85]. Tuna [HM86, Hil89]. Turbid [PB88]. Turbidity [LKL87, Llo87]. Turbine [BK85, BJFF87]. Turf [Dom87]. Two [Alc81, BR88a, BNC^{+}86, CK85, CW85, CK88, EC85, HM85, HP82, Har86, Hau83, Hil86, JS82, Jon82, PB81, RW88, RT86, SHB85, SHP87, Scr88, TH86b, Vin87]. Two-Stage [CW85]. Two-Story [Jon82]. Type [Dot82]. Types [CRS87, Eck83].

Ultrasonic [BTPM89, MMS^{+}83]. Unbaited [PCC81]. Unbalanced [PH85]. Uncertainty [BP86b, Hil87]. Undercut [BMW86]. Underwater [SM87, SGS88]. Unexploited [RR89]. Unimproved [Rin82]. Unit [LUBO85, MBRG83, Par89a, Ser83]. United [FNW82, Low81, Ret84]. Unobserved [TP84]. Unobstructed [Nes87]. Unregulated [JS83]. Unsuccessful [PH85]. Updated [Nic86]. Upon [FK83]. Upper [HN82b, MR86, PCC81, Pit89, SGM89, SB83]. Urban [Alc81, Car82b]. Usable [CA87]. Use [AGN81, BTPM89, BNHF88, BMW86, CSH^{+}87, CK88, CP87, HB85a, Heg85, HC88, Jen89, KN88, LO88, LRLM84, LM87, LJ89, MB86, MK88, MS86, MGS87, PB88, Par89b, PN88, Rog85, RT86, SM85, SRJ82, Sch84, Sch86, SJ85, Tap86, TBM84, Wil84, Wil87b]. Used [Joh87, MS85a, TW87]. Useful [MR87]. Using [Eri83, FNN82, FF82, Mir86, POS87, SA84]. Utah [LT89, HD87, HB83, PB82, SH87]. Utility [BNHF88]. Utilization [MEB84].

Vaccinated [BSS82]. Validating [BC87]. Validity [CB83, KS89]. Valley [WW89]. Variability [BP86a, SM86, Wal88, Win89]. Variable [PW87a].

Variables [LMW87, MR87]. Variation [Chi89, FMW88, HS83, PPB89, Rin85]. Variations [Enn83, TH86b]. Various [PW87b, SWH85]. Varying [Cus85, HRDM84]. Vegetated [DHBZ89]. Vegetation [CCS89, DPK84, MWD86, WGF87]. Velocity [BJM82, SJ85]. Vermont [WL84]. versus [Ueb86]. Vertebrae [Bil86]. Vertical [CCW87, LGJR89, Neg82]. Vessels [Pof85]. Viable [PW84]. Vibrio [BSS82]. Violations [CRK88]. Virginia [And89, AO88, CH88, DW81, NBH85, Nie82, NO88, PH85]. Visible [SHS85]. Visual [NP87]. Vital [Ebb87, Laa81]. Voltage [Hud85, SH86b]. Voluntarily [Qui89]. Voluntary [Cla83]. Volunteered [FGM84]. Voyageurs [Kal87]. vs [CM83, WT83]. Vulnerability [BZM84, BML^{+}86, Mar87].

Walleye [BR88a, BR88b, Hau83, JSC88, Kal87, Laa81, LJS82, Par89b, Pit89, RC88, SM85, See87b, Ser82a, Ser82b, Ser83, Ser87, WT84, WS87]. Walleyes [BC87, Eri83, Fle87, IR89, JSC88, Kal89, MMC87, MW86, PPP89, Sch89a, Sch89b, Wil87b, Wol81]. Warmwater [BA88, EGT^{+}84, Par89b]. Washington [Fle87, KD87, KKM85, MWD86, PH89, RHT88, STP83, WG83]. Water [Bea83, BJM82, BMJ85, GM85, HN82a, Hof86, HBDZ89, Llo87, McK87, MSB84, Say87, TS82]. Waters [IR89, LKL87, MSC81, PDA82, SHR83, Wol81]. Watershed [GEB86, KN88]. Watersheds [PN85]. Watts [VHJB83]. Wavelengths [SHS85]. Weather [SGO88]. Wedge [WBJR87]. Wedge-Wire [WBJR87]. Weighing [Jen89]. Weight [Say87, Wil89]. Weighted [CA87]. Weir [KPR^{+}89, MWTO89]. Weirs [KN88]. Wells [PGR88]. West [MWD86, AO88, Bry83, MSB84, Nie82, NO88, PN89, SM85]. West-Side [MWD86]. Western [Bru84, IR89, KD87, MW86, NP87, VJL82]. While [Mat83]. White [Bea83, Box86a, FM87, GM85, LWL89, MG88, QR82, VHJB83, Rin85]. Whitefish [BG82, Bru84, EC82, EC85, JT85, ST85, SRJ82]. Wild [AN84, Ber82, EH86, FL85, Vin87]. Winter [CSP88, CP87, HMK86, New89, RM82, SMJR89]. Wire [ES86, FHB87, HC88, KP86, MZ87, PW87b, WBJR87]. Wisconsin [HBS83, Hau83, MSC81, RM82, Ser82a, Ser82b, Ser87]. without [HFG85]. Wooded [WJ86]. Woody [Bry83, MK89]. Wound [Box82]. Wounds [Jac89]. Wyoming [Hub88, KH89, NH88, SH87, WGH87].

x [KH82].

Year [Hau83, Kal87, KK84, MSB84, MM87, Nic84, OC89, RC88, Ser82b, Ser87, SKB82, VHJB83, WL84, DP84]. Year-Class [Hau83, RC88]. Yearling [EH86, HES86]. Yearlings [EOS88, Ser83]. Years [Cla85, GH83, RH82]. Yellow [HW83, HS83, Kal87, SJT^{+}82, VHJB83]. Yellowfin [Hil89]. Yellowstone [SGG86, SGG86]. Yellowtail [Wal88].

Yield [Bar88, Bru84, Cam85, DRM84, EC85, Hil89, MM85, MMC87, Nic86]. Yield-per-Recruit [Nic86]. Yields [KNR88, Ueb86, WT84]. York [DW81, SA84, Wol81]. Young [HGH82, Kal87, MSB84, MM87, Ser82b, SM83, SKB82, VHJB83]. Young-of-the-Year [Kal87, SKB82, VHJB83]. Young-of-Year [MSB84, MM87, Ser82b]. Yukon [Mun82].

Zealand [Dew89]. Zones [SKB82]. Zooplankton [MS82, MGS87]. Zooplanktophagous [SH87]. Zostera [FTCF84].

References

Archibald:1983:RSH

[AFL83] Chris P. Archibald, David Fournier, and Bruce M. Leaman. Reconstruction of stock history and development of rehabilitation strategies for Pacific Ocean perch in Queen Charlotte Sound, Canada. North American Journal of Fisheries Management, 3(3):283–294, July 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ashley:1981:UAL

[AGN81] Keith W. Ashley, Donald L. Garling, Jr., and John J. Ney. Use of the age–length relationship as a marker for differentiating geographical stocks of adult channel catfish. North American Journal of Fisheries Management, 1(1):77–79, January 1981. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Alexander:1983:SSM

[AH83] Gaylord R. Alexander and Edward A. Hansen. Sand sediment in a Michigan trout stream, Part II. Effects of reducing sand, bedload on a trout population. North American Journal of Fisheries Management, 3(4):365–372, October 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Alexander:1986:SBL

[AH86] Gaylord R. Alexander and Edward A. Hansen. Sand bed load in a brook trout stream. North American Journal of Fisheries Management, 6(1):9–23, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Angermeier:1986:AIB

[AK86] Paul L. Angermeier and James R. Karr. Applying an index of biotic integrity based on stream-fish communities: Considerations in sampling and interpretation. North American Journal of Fisheries Management, 6(3):418–429, July 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Alcorn:1981:FQT

[Alc81] Steven R. Alcorn. Fishing quality in two urban fishing lakes, St. Louis, Missouri. North American Journal of Fisheries Management, 1(1):80–84, January 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Almeida:1987:SDS

[Alm87] Frank P. Almeida. Stock definition of silver hake in the New England–Middle Atlantic area. North American Journal of Fisheries Management, 7(2):169–186, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:1984:ECR

[AN84] Richard M. Anderson and R. Barry Nehring. Effects of a catchand-release regulation on a wild trout population in Colorado and its acceptance by anglers. North American Journal of Fisheries Management, 4(3):257–265, July 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:1989:EBH

[And89] Eric E. Anderson. Economic benefits of habitat restoration: Seagrass and the Virginia hard-shell blue crab fishery. North American Journal of Fisheries Management, 9(2):140–149, May 1989. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:1982:Ea

[Ano82a] Anonymous. Errata. North American Journal of Fisheries Management, 2(2):208, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:1982:Eb

[Ano82b] Anonymous. Errata. North American Journal of Fisheries Management, 2(4):306, July 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:1982:Ec

[Ano82c] Anonymous. Errata. North American Journal of Fisheries Management, 2(4):414, October 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:1986:E

[Ano86] Anonymous. Errata. North American Journal of Fisheries Management, 6(3):451, July 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:1987:E

[Ano87] Anonymous. Errata. North American Journal of Fisheries Management, 7(1):161, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Austen:1988:EMM

[AO88] Douglas J. Austen and Donald J. Orth. Evaluation of a 305-mm minimum-length limit for smallmouth bass in the New River, Virginia and West Virginia. North American Journal of Fisheries Management, 8(2):231–239, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Angermeier:1987:ABI

[AS87] Paul L. Angermeier and Isaac J. Schlosser. Assessing biotic integrity of the fish community in a small Illinois stream. North American Journal of Fisheries Management, 7(3):331–338, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Atkinson:1989:DMB

[Atk89] D. B. Atkinson. Diel movements of beaked redfish and the implications of these for stratified random bottom trawl estimates of biomass and abundance. North American Journal of Fisheries Management, 9(2):163–170, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Axon:1981:DMF

[Axo81] James R. Axon. Development of a muskellunge fishery at Cave Run Lake, Kentucky, 1974–1979. North American Journal of Fisheries Management, 1(2):134–143, April 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bayley:1988:CDC

[BA88] Peter B. Bayley and Douglas J. Austen. Comparison of detonating cord and rotenone for sampling fish in warmwater impoundments. North American Journal of Fisheries Management, 8(3):310–316, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bain:1987:SDM

[Bai87] Mark B. Bain. Structured decision making in fisheries management: Trout fishing regulations on the Au Sable River, Michigan. North American Journal of Fisheries Management, 7(4):475–481, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bams:1985:CTI

[Bam85] Robert A. Bams. Comparison of three instream incubation techniques for Coho salmon. North American Journal of Fisheries Management, 5(2A):159–172, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barwick:1984:RFD

[Bar84] D. Hugh Barwick. Role of fish distribution on estimates of standing crop in a cooling reservoir. North American Journal of Fisheries Management, 4(3):308–313, July 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barwick:1985:SHM

[Bar85] D. Hugh Barwick. Stocking and hooking mortality of planted rainbow trout in Jocassee Reservoir, South Carolina. North American Journal of Fisheries Management, 5(4):580–583, October 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barber:1988:MSY

[Bar88] Willard E. Barber. Maximum sustainable yield lives on. North American Journal of Fisheries Management, 8(2):153–157, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barnhart:1989:SRC

[Bar89] Roger A. Barnhart. Symposium review: Catch-and-release fishing, a decade of experience. North American Journal of Fisheries Management, 9(1):74–80, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bisson:1982:ASS

[BB82] Peter A. Bisson and Robert E. Bilby. Avoidance of suspended sediment by juvenile Coho salmon. North American Journal of Fisheries Management, 2(4):371–374, October 1982. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Babey:1989:PSP

[BB89] George J. Babey and Charles R. Berry, Jr. Post-stocking performance of three strains of rainbow trout in a reservoir. North American Journal of Fisheries Management, 9(3):309–315, August 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Babaluk:1987:PRT

[BC87] John A. Babaluk and J. Scott Campbell. Preliminary results of tetracycline labelling for validating annual growth increments in opercula of walleyes. North American Journal of Fisheries Management, 7(1):138–141, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beam:1983:EAW

[Bea83] James H. Beam. The effect of annual water level management on population trends of white crappie in Elk City Reservoir, Kansas. North American Journal of Fisheries Management, 3(1):34–40, January 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Berger:1982:SFW

[Ber82] Tommie A. Berger. Supplemental feeding of a wild bluegill population. North American Journal of Fisheries Management, 2(2): 158–163, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bain:1985:QSS

[BFB85a] Mark B. Bain, John T. Finn, and Henry E. Booke. Quantifying stream substrate for habitat analysis studies. North American Journal of Fisheries Management, 5(3B):499–500, July 1985. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bain:1985:QMS

[BFB85b] Mark B. Bain, John T. Finn, and Henry E. Booke. A quantitative method for sampling riverine microhabitats by electrofishing. North American Journal of Fisheries Management, 5(3B):489–493, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Berkes:1982:MME

[BG82] Fikret Berkes and Türker Gönenc. A mathematical model on the exploitation of northern lake whitefish with gill nets. North American Journal of Fisheries Management, 2(2):176–183, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barrett:1988:EDC

[BG88] Jeffrey C. Barrett and Gary D. Grossman. Effects of direct current electrofishing on the mottled sculpin. North American Journal of Fisheries Management, 8(1):112–116, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Belford:1989:ETP

[BG89] David A. Belford and William R. Gould. An evaluation of trout passage through six highway culverts in Montana. North American Journal of Fisheries Management, 9(4):437–445, Fall 1989. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boyd:1982:ART

[BHN82] Claude E. Boyd, Md Hanapi, and Md Noor. Aquashade(R) treatment of channel catfish ponds. North American Journal of Fisheries Management, 2(2):193–196, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bilton:1986:MCS

[Bil86] H. T. Bilton. Marking chum salmon fry vertebrae with oxytetracycline. North American Journal of Fisheries Management, 6(1):126– 128, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bronte:1983:OSF

[BJ83] Charles R. Bronte and Donald W. Johnson. Occurrence of sport fish in a commercial net fishery in Kentucky. North American Journal of Fisheries Management, 3(3):239–242, July 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bronte:1984:ECE

[BJ84] Charles R. Bronte and Donald W. Johnson. Evaluation of the commercial entanglement-gear fishery in Lake Barkley and Kentucky Lake, Kentucky. North American Journal of Fisheries Management, 4(1):75–83, January 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brege:1987:BNC

[BJFF87] Dean A. Brege, Richard C. Johnsen, Richard W. Frazier, and Winston E. Farr. Box net for collection of juvenile salmonids from turbine intake gatewells at John Day Dam. North American Journal of Fisheries Management, 7(3):446–448, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beland:1982:WDV

[BJM82] Kenneth F. Beland, Richard M. Jordan, and Alfred L. Meister. Water depth and velocity preferences of spawning Atlantic salmon in Maine rivers. North American Journal of Fisheries Management, 2(1):11–13, January 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brauhn:1982:SGC

[BK82] James L. Brauhn and Harold Kincaid. Survival, growth, and catchability of rainbow trout of four strains. North American Journal of Fisheries Management, 2(1):1–10, January 1982. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bell:1985:MAA

[BK85] Charles E. Bell and Boyd Kynard. Mortality of adult American shad passing through a 17-megawatt kaplan turbine at a low-head hydroelectric dam. North American Journal of Fisheries Management, 5(1):33–38, January 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barry:1986:AAA

[BK86] Timothy Barry and Boyd Kynard. Attraction of adult American shad to fish lifts at Holyoke Dam, Connecticut River. North American Journal of Fisheries Management, 6(2):233–241, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buynak:1989:MBP

[BKJ89] Gerard L. Buynak, Benjamin T. Kinman, and Ralph V. Jackson. Management brief: Purse seining to capture large numbers of adult threadfin shad. North American Journal of Fisheries Management, 9(1):121–123, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buynak:1989:LDE

[BKSM89] Gerard L. Buynak, Lewis E. Kornman, Albert Surmont, and Bill Mitchell. Longitudinal differences in electrofishing catch rates and angler catches of black bass in Cave Run Lake, Kentucky. North American Journal of Fisheries Management, 9(2):226–230, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:1981:ISS

[BM81] Larry R. Brown and Peter B. Moyle. The impact of squawfish on salmonid populations: a review. North American Journal of Fisheries Management, 1(2):104–111, April 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Babcock:1985:QMM

[BM85] A. M. Babcock and P. R. Mundy. A quantitative measure of migratory timing applied to the commercial brown shrimp fishery in North Carolina. North American Journal of Fisheries Management, 5(2A):181–196, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bennett:1985:GMP

[BMJ85] David H. Bennett, O. Eugene Maughan, and Douglas B. Jester, Jr. Generalized model for predicting spawning success of fish in reservoirs with fluctuating water levels. North American Journal of Fisheries Management, 5(1):12–20, January 1985. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Burkett:1986:HLV

[BML^{+}86] Dale P. Burkett, Philip C. Mankin, George W. Lewis, William F. Childers, and David P. Philipp. Hook-and-line vulnerability and multiple recapture of largemouth bass under a minimum totallength limit of 457 mm. North American Journal of Fisheries Management, 6(1):109–112, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bills:1981:PBA

[BMM81] T. D. Bills, L. L. Marking, and W. L. Mauck. Polychlorinated biphenyl (Aroclor (R) 1254) residues in rainbow trout: Effects on sensitivity to nine fishery chemicals. North American Journal of Fisheries Management, 1(2):200–203, April 1981. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Burczynski:1987:HAA

[BMM87] Janusz J. Burczynski, Paul H. Michaletz, and Gary M. Marrone. Hydroacoustic assessment of the abundance and distribution of rainbow smelt in Lake Oahe. North American Journal of Fisheries Management, 7(1):106–116, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brusven:1986:SUS

[BMW86] Merlyn A. Brusven, William R. Meehan, and John F. Ward. Summer use of simulated undercut banks by juvenile Chinook salmon in an artificial Idaho channel. North American Journal of Fisheries Management, 6(1):32–37, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Becker:1983:EDC

[BNA83] C. Dale Becker, Duane A. Neitzel, and C. Scott Abernethy. Effects of dewatering on Chinook salmon redds: Tolerance of four development phases to one-time dewatering. North American Journal of Fisheries Management, 3(4):373–382, October 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bisson:1986:OAB

[BNC^{+}86] Peter A. Bisson, Jennifer L. Nielsen, Mark W. Chilcote, Bruce Crawford, and Steven A. Leider. Occurrence of anadromous brown trout in two Lower Columbia River tributaries. North American Journal of Fisheries Management, 6(2):290–292, April 1986. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boutillier:1988:UMU

[BNHF88] J. Boutillier, D. Noakes, D. Heritage, and J. Fulton. Use of multiattribute utility theory for designing invertebrate fisheries sampling programs. North American Journal of Fisheries Management, 8 (1):84–90, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bourgeois:1987:CBF

[BOS87] C. E. Bourgeois, M. F. O’Connell, and D. C. Scott. Cold-branding and fin-clipping Atlantic salmon smolts on the Exploits River, Newfoundland. North American Journal of Fisheries Management, 7 (1):154–156, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boxrucker:1982:MMF

[Box82] J. C. Boxrucker. Mass marking of fingerling largemouth bass by fin-clipping followed by freeze-cauterization of the wound. North American Journal of Fisheries Management, 2(1):94–96, January 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boxrucker:1986:COS

[Box86a] Jeff Boxrucker. A comparison of the otolith and scale methods for aging white crappies in Oklahoma. North American Journal of Fisheries Management, 6(1):122–125, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boxrucker:1986:ESS

[Box86b] Jeff Boxrucker. Evaluation of supplemental stocking of largemouth bass as a management tool in small impoundments. North American Journal of Fisheries Management, 6(3):391–396, July 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boxrucker:1987:LBI

[Box87] Jeff Boxrucker. Largemouth bass influence on size structure of crappie populations in small Oklahoma impoundments. North American Journal of Fisheries Management, 7(2):273–278, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bowering:1986:DVT

[BP86a] W. R. Bowering and D. G. Parsons. Diel variability in trawl catches of Greenland halibut from the channels off coastal Labrador and implications for resource assessment. North American Journal of Fisheries Management, 6(2):149–155, April 1986. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:1986:RAG

[BP86b] B. E. Brown and G. P. Patil. Risk analysis in the Georges Bank haddock fishery — a pragmatic example of dealing with uncertainty. North American Journal of Fisheries Management, 6(2): 183–191, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brunson:1986:EMB

[BR86] Martin W. Brunson and H. Randall Robinette. Evaluation of male bluegill × female green sunfish hybrids for stocking Mississippi farm ponds. North American Journal of Fisheries Management, 6(2): 156–167, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baccante:1988:FCT

[BR88a] Dominic A. Baccante and David M. Reid. Fecundity changes in two exploited walleye populations. North American Journal of Fisheries Management, 8(2):199–209, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beamesderfer:1988:SSB

[BR88b] Raymond C. Beamesderfer and Bruce E. Rieman. Size selectivity and bias in estimates of population statistics of smallmouth bass, walleye, and northern squawfish in a Columbia River reservoir. North American Journal of Fisheries Management, 8(4):505– 510, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Breen:1987:MDC

[Bre87] Paul A. Breen. Mortality of Dungeness crabs caused by lost traps in the Fraser River Estuary, British Columbia. North American Journal of Fisheries Management, 7(3):429–435, July 1987. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bruce:1984:PFY

[Bru84] W. J. Bruce. Potential fisheries yield from Smallwood Reservoir, Western Labrador, with special emphasis on lake whitefish. North American Journal of Fisheries Management, 4(1):48–66, January 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bryant:1983:RMW

[Bry83] Mason D. Bryant. The role and management of woody debris in West Coast salmonid nursery streams. North American Journal of Fisheries Management, 3(3):322–330, July 1983. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baum:1982:SHR

[BSS82] Edward T. Baum, Evelyn S. Sawyer, and Richard G. Strout. Survival of hatchery-reared Atlantic salmon smolts vaccinated with a Vibrio anguillarum bacterin. North American Journal of Fisheries Management, 2(4):409–411, October 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barton:1985:DRB

[BTB85] David R. Barton, William D. Taylor, and R. M. Biette. Dimensions of riparian buffer strips required to maintain trout habitat in Southern Ontario streams. North American Journal of Fisheries Management, 5(3A):364–378, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonar:1989:MBU

[BTPM89] Scott A. Bonar, G. L. Thomas, Gilbert B. Pauley, and Roy W. Martin. Management briefs: Use of ultrasonic images for rapid nonlethal determination of sex and maturity of Pacific herring. North American Journal of Fisheries Management, 9(3):364–366, August 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bulak:1983:EFA

[Bul83] James S. Bulak. Evaluation of floy anchor tags for short term markrecapture studies with blueback herring. North American Journal of Fisheries Management, 3(1):91–94, January 1983. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bulow:1988:MBE

[BWCQ88] Frank J. Bulow, Mark A. Webb, W. Dennis Crumby, and Scott S. Quisenberry. Management briefs: Effectiveness of a fish barrier dam in limiting movement of rough fishes from a reservoir into a tributary stream. North American Journal of Fisheries Management, 8(2):273–275, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bickerstaff:1984:VRT

[BZM84] Wesley B. Bickerstaff, Charles D. Ziebell, and William J. Matter. Vulnerability of redbelly tilapia fry to bluegill predation with changes in cover availability. North American Journal of Fisheries Management, 4(1):120–125, January 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Conder:1987:TWU

[CA87] Allen L. Conder and Thomas C. Annear. Test of weighted usable area estimates derived from a PHABSIM model for instream flow studies on trout streams. North American Journal of Fisheries Management, 7(3):339–350, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:1981:HER

[CAG81] Richard D. Clark, Jr., Gaylord R. Alexander, and Howard Gowing. A history and evaluation of regulations for brook trout and brown trout in Michigan streams. North American Journal of Fisheries Management, 1(1):1–14, January 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Campbell:1985:AYE

[Cam85] Alan Campbell. Application of a yield and egg-per-recruit model to the lobster fishery in the Bay of Fundy. North American Journal of Fisheries Management, 5(1):91–104, January 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carl:1982:NRC

[Car82a] L. M. Carl. Natural reproduction of Coho salmon and Chinook salmon in some Michigan streams. North American Journal of Fisheries Management, 2(4):375–380, October 1982. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carl:1982:SIS

[Car82b] L. M. Carl. Social impacts of a stream reclamation project on urban anglers. North American Journal of Fisheries Management, 2(2): 164–170, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cass:1983:FEV

[CB83] Alan J. Cass and Richard J. Beamish. First evidence of validity of the fin-ray method of age determination for marine fishes.

North American Journal of Fisheries Management, 3(2):182–188, April 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clapp:1989:HMS

[CC89] David F. Clapp and Richard D. Clark, Jr. Hooking mortality of smallmouth bass caught on live minnows and artificial spinners. North American Journal of Fisheries Management, 9(1):81– 85, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Crawford:1989:TAF

[CCP89] Steve Crawford, William S. Coleman, and Wesley F. Porak. Time of annulus formation in otoliths of Florida largemouth bass. North American Journal of Fisheries Management, 9(2):231–233, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Colle:1989:DFL

[CCS89] Douglas E. Colle, Richard L. Cailteux, and Jerome V. Shireman. Distribution of Florida largemouth bass in a lake after elimination of all submersed aquatic vegetation. North American Journal of Fisheries Management, 9(2):213–218, May 1989. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chadwick:1987:VGN

[CCW87] James W. Chadwick, Mark F. Cook, and David S. Winters. Vertical gill net for studying depth distribution of fish in small lakes. North American Journal of Fisheries Management, 7(4):593–594, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:1985:CBS

[CH85] Richard D. Clark, Jr. and Bin Huang. Conflict between sportfishing, commercial fishing, and rehabilitation of lake trout in Lake Michigan. North American Journal of Fisheries Management, 5 (2B):261–276, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chipman:1988:RSM

[CH88] Brian D. Chipman and Louis A. Helfrich. Recreational specializations and motivations of Virginia River anglers. North American Journal of Fisheries Management, 8(4):390–398, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chittenden:1989:SVC

[Chi89] Mark E. Chittenden, Jr. Sources of variation in Chesapeake Bay pound-net and haul-seine catch compositions. North American Journal of Fisheries Management, 9(3):367–371, August 1989. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carline:1984:EIP

[CJH84] Robert F. Carline, Barry L. Johnson, and Thomas J. Hall. Estimation and interpretation of proportional stock density for fish populations in Ohio impoundments. North American Journal of Fisheries Management, 4(2):139–154, April 1984. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carline:1985:RFP

[CK85] Robert F. Carline and Steven P. Klosiewski. Responses of fish populations to mitigation structures in two small channelized streams in Ohio. North American Journal of Fisheries Management, 5(1): 1–11, January 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cone:1988:CSE

[CK88] R. Scott Cone and Charles C. Krueger. Comparison of survival, emigration, habitat use, marking mortality, and growth between two strains of brook trout in Adirondack ponds. North American Journal of Fisheries Management, 8(4):497–504, November 1988. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:1983:PEV

[Cla83] Richard D. Clark, Jr. Potential effects of voluntary catch and release of fish on recreational fisheries. North American Journal of Fisheries Management, 3(3):306–314, July 1983. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:1985:FSC

[Cla85] William G. Clark. Fishing in a sea of court orders: Puget Sound salmon management 10 years after the boldt decision. North American Journal of Fisheries Management, 5(3B):417–434, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cofer:1983:CVI

[CM83] Larry M. Cofer and Stephen P. Malvestuto. Centimeter vs. inchgroup measurements for describing age and size composition of fish populations. North American Journal of Fisheries Management, 3 (4):445–450, October 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chapman:1986:DMR

[CM86] D. W. Chapman and B. May. Downstream movement of rainbow trout past Kootenai Falls, Montana. North American Journal of Fisheries Management, 6(1):47–51, January 1986. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:1988:CSM

[CM88] Richard D. Clark, Jr. and Roger A. Martin. Computerized system for managing hatchery fish transportation in Michigan. North American Journal of Fisheries Management, 8(1):102–111, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Coble:1988:EAB

[Cob88] Daniel W. Coble. Effects of angling on bluegill populations: Management implications. North American Journal of Fisheries Management, 8(3):277–283, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Combs:1982:AEP

[Com82] David L. Combs. Angler exploitation of paddlefish in the Neosho River, Oklahoma. North American Journal of Fisheries Management, 2(4):334–342, October 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Combs:1982:SDS

[CP82] David L. Combs and Lawrence R. Peltz. Seasonal distribution of striped bass in Keystone Reservoir, Oklahoma. North American Journal of Fisheries Management, 2(1):66–73, January 1982. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cunjak:1987:CUS

[CP87] Richard A. Cunjak and Geoffrey Power. Cover use by streamresident trout in winter: a field experiment. North American Journal of Fisheries Management, 7(4):539–544, October 1987. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chadwick:1986:SRR

[CR86] E. M. P. Chadwick and R. G. Randall. A stock-recruitment relationship for Atlantic salmon in the Miramichi River, New Brunswick. North American Journal of Fisheries Management, 6(2):200–203, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Crance:1987:HSI

[Cra87] Johnie H. Crance. Habitat suitability index curves for paddlefish, developed by the Delphi technique. North American Journal of Fisheries Management, 7(1):123–130, January 1987. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cone:1988:FST

[CRK88] R. Scott Cone, Douglas S. Robson, and Charles C. Krueger. Failure of statistical tests to detect assumption violations in the mark– recapture population estimation of brook trout in Adirondack ponds. North American Journal of Fisheries Management, 8(4): 489–496, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Christian:1987:EST

[CRS87] Paul A. Christian, Mac V. Rawson, Jr., and Thomas D. Shierling. Effectiveness of six trap types and other factors for harvesting premolt blue crabs (Callinectes sapidus) in a Georgia tidal creek. North American Journal of Fisheries Management, 7(1): 63–70, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chilton:1987:COS

[CS87] Doris E. Chilton and Max Stocker. A comparison of otolith and scale methods for aging Pacific herring. North American Journal of Fisheries Management, 7(2):202–206, April 1987. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Colle:1987:IHH

[CSH^{+}87] Douglas E. Colle, Jerome V. Shireman, William T. Haller, Joseph C. Joyce, and Daniel E. Canfield, Jr. Influence of hydrilla on harvestable sport-fish populations, angler use, and angler expenditures at Orange Lake, Florida. North American Journal of Fisheries Management, 7(3):410–417, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cederholm:1988:LCE

[CSP88] C. J. Cederholm, W. J. Scarlett, and N. P. Peterson. Low-cost enhancement technique for winter habitat of juvenile Coho salmon. North American Journal of Fisheries Management, 8(4):438–441, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cushman:1985:REE

[Cus85] Robert M. Cushman. Review of ecological effects of rapidly varying flows downstream from hydroelectric facilities. North American Journal of Fisheries Management, 5(3A):330–339, July 1985. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chester:1985:TSS

[CW85] Alexander J. Chester and James R. Waters. Two-stage sampling for age distribution in the Atlantic menhaden fishery, with comments on optimal survey design. North American Journal of Fisheries Management, 5(3B):449–456, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Donald:1989:EEM

[DA89] David B. Donald and David J. Alger. Evaluation of exploitation as a means of improving growth in a stunted population of brook trout. North American Journal of Fisheries Management, 9(2): 177–183, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Damkaer:1989:EFE

[DD89] David M. Damkaer and Douglas B. Dey. Evidence for fluoride effects on salmon passage at John Day Dam, Columbia River, 1982– 1986. North American Journal of Fisheries Management, 9(2): 154–162, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dew:1988:BCC

[Dew88] C. Braxton Dew. Biological characteristics of commercially caught Hudson River striped bass, 1973–1975. North American Journal of Fisheries Management, 8(1):75–83, February 1988. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dewees:1989:AII

[Dew89] Christopher M. Dewees. Assessment of the implementation of individual transferable quotas in New Zealand’s inshore fishery. North American Journal of Fisheries Management, 9(2):131–139, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dewey:1989:CFC

[DHBZ89] Michael R. Dewey, Leslie E. Holland-Bartels, and Steven J. Zigler. Comparison of fish catches with buoyant pop nets and seines in vegetated and nonvegetated habitats. North American Journal of Fisheries Management, 9(2):249–253, May 1989. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dimond:1985:DIE

[Dim85] William F. Dimond. Device to increase efficiency of acrylic tubes for removing stomach contents of fish. North American Journal of Fisheries Management, 5(2A):214, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Douglas:1987:RHS

[DJ87] David R. Douglas and Larry A. Jahn. Radiotracking hybrid striped bass in Spring Lake, Illinois, to determine temperature and oxygen preferences. North American Journal of Fisheries Management, 7 (4):531–534, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dawson:1989:CFI

[DLP89] Chad P. Dawson, Frank R. Lichtkoppler, and Charles Pistis. The charter fishing industry in the Great Lakes. North American Journal of Fisheries Management, 9(4):493–499, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dorr:1981:SCA

[DOFJ81] John A. Dorr III, Daniel V. O’Connor, Neal R. Foster, and David J. Jude. Substrate conditions and abundance of lake trout eggs in a traditional spawning area in Southeastern Lake Michigan. North American Journal of Fisheries Management, 1(2):165–172, April 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dombeck:1987:ATI

[Dom87] Michael P. Dombeck. Artificial turf incubators for raising muskellunge to swim-up fry. North American Journal of Fisheries Management, 7(3):425–428, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Donald:1987:AOE

[Don87] D. B. Donald. Assessment of the outcome of eight decades of trout stocking in the mountain national parks, Canada. North American Journal of Fisheries Management, 7(4):545–553, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dotson:1982:MTC

[Dot82] Thurston Dotson. Mortalities in trout caused by gear type and angler-induced stress. North American Journal of Fisheries Management, 2(1):60–65, January 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dwyer:1984:TYH

[DP84] William P. Dwyer and Robert G. Piper. Three-year hatchery and field evaluation of four strains of rainbow trout. North American Journal of Fisheries Management, 4(2):216–221, April 1984. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Durocher:1984:RBA

[DPK84] Philip P. Durocher, William C. Provine, and Joseph E. Kraai. Relationship between abundance of largemouth bass and submerged vegetation in Texas reservoirs. North American Journal of Fisheries Management, 4(1):84–88, January 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunning:1982:EMS

[DRG82] Dennis J. Dunning, Quentin Ross, and John Gladden. Evaluation of minimum size limits for St. Lawrence River northern pike. North American Journal of Fisheries Management, 2(2):171–175, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Doubleday:1984:EPR

[DRM84] W. G. Doubleday, D. Rivard, and W. D. McKone. Estimation of partial recruitment and yield per recruit for an otter trawl fishery for deepwater redfish. North American Journal of Fisheries Management, 4(1):15–31, January 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunning:1989:RMS

[DRM^{+}89] Dennis J. Dunning, Quentin E. Ross, Mark T. Mattson, Paul Geoghegan, and John R. Waldman. Reducing mortality of striped bass captured in seines and trawls. North American Journal of Fisheries Management, 9(2):171–176, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunning:1987:TRT

[DRWM87] Dennis J. Dunning, Quentin E. Ross, John R. Waldman, and Mark T. Mattson. Tag retention by, and tagging mortality of, Hudson River striped bass. North American Journal of Fisheries Management, 7(4):535–538, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Davies:1988:LTC

[DS88] P. E. Davies and R. D. Sloane. Long-term changes in brown trout and rainbow trout populations in Great Lake, Tasmania. North American Journal of Fisheries Management, 8(4):463–474, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dawson:1981:MNY

[DW81] Chad P. Dawson and Bruce T. Wilkins. Motivations of New York and Virginia marine boat anglers and their preferences for potential fishing constraints. North American Journal of Fisheries Management, 1(2):151–158, April 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Easley:1982:PRS

[Eas82] J. E. Easley, Jr. Property rights in shellfish relay: Managing fisheries for higher economic returns. North American Journal of Fisheries Management, 2(4):343–350, October 1982. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ebbers:1987:VSL

[Ebb87] Mark A. Ebbers. Vital statistics of a largemouth bass population in Minnesota from electrofishing and angler-supplied data. North American Journal of Fisheries Management, 7(2):252–259, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ebener:1982:LFA

[EC82] Mark P. Ebener and Frederick A. Copes. Loss of floy anchor tags from lake whitefish. North American Journal of Fisheries Management, 2(1):90–93, January 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ebener:1985:PSY

[EC85] Mark P. Ebener and Fredrick A. Copes. Population statistics, yield estimates, and management considerations for two lake whitefish stocks in Lake Michigan. North American Journal of Fisheries Management, 5(3B):435–448, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Eckmayer:1983:GSH

[Eck83] William J. Eckmayer. Growth and survival of hatchery-reared American oysters set on three types of cultch and in Bon Secour Bay, Alabama. North American Journal of Fisheries Management, 3(2):171–175, April 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Evenson:1985:LTR

[EE85] Michael D. Evenson and R. D. Ewing. Long-term retention of fluorescent pigment marks by spring Chinook salmon and summer steelhead. North American Journal of Fisheries Management, 5(1): 26–32, January 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Edwards:1984:MEA

[EGT^{+}84] Clayton J. Edwards, Bernard L. Griswold, Richard A. Tubb, Earl C. Weber, and Lewis C. Woods. Mitigating effects of artificial riffles and pools on the fauna of a channelized warmwater stream. North American Journal of Fisheries Management, 4(2): 194–203, April 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ersbak:1983:NDA

[EH83] Kenneth Ersbak and Bruce L. Haase. Nutritional deprivation after stocking as a possible mechanism leading to mortality in streamstocked brook trout. North American Journal of Fisheries Management, 3(2):142–151, April 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Engstrom-Heg:1986:PWB

[EH86] Robert Engstrom-Heg. Prediction of wild brown trout catch rates from estimated yearling population density and fishing intensity. North American Journal of Fisheries Management, 6(3):410–417, July 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Enzenhofer:1989:SCM

[EH89] Hermann J. Enzenhofer and Jeremy M. B. Hume. Simple closing midwater trawl for small boats. North American Journal of Fisheries Management, 9(3):372–377, August 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ennis:1983:AVS

[Enn83] G. P. Ennis. Annual variations in standing stock in a Newfoundland population of lobsters. North American Journal of Fisheries Management, 3(1):26–33, January 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ennis:1985:AIS

[Enn85] G. P. Ennis. An assessment of the impact of size limit and exploitation rate changes on egg production in a Newfoundland lobster population. North American Journal of Fisheries Management, 5(1): 86–90, January 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Elrod:1988:CHR

[EOS88] Joseph H. Elrod, David E. Ostergaard, and Clifford P. Schneider. Comparison of hatchery-reared lake trout stocked as fall fingerlings and as spring yearlings in Lake Ontario. North American Journal of Fisheries Management, 8(4):455–462, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Edsall:1989:SSS

[EPNB89] Thomas A. Edsall, Thomas P. Poe, Robert T. Nester, and Charles L. Brown. Side-scan sonar mapping of lake trout spawning habitat in Northern Lake Michigan. North American Journal of Fisheries Management, 9(3):269–279, August 1989. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Erickson:1983:ADM

[Eri83] C. M. Erickson. Age determination of Manitoban walleyes using otoliths, dorsal spines, and scales. North American Journal of Fisheries Management, 3(2):176–181, April 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Elrod:1986:ECW

[ES86] Joseph H. Elrod and Clifford P. Schneider. Evaluation of coded wire tags for marking lake trout. North American Journal of Fisheries Management, 6(2):264–271, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Eder:1987:CCF

[ES87] Stephen Eder and William E. Mcdannold Sr. The channel catfish fishery at Pony Express Lake, Missouri, 1963–1984. North American Journal of Fisheries Management, 7(4):502–512, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Evans:1981:PCF

[Eva81] Geoffrey T. Evans. The potential collapse of fish stocks in a developing fishery. North American Journal of Fisheries Management, 1(2):127–133, April 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fraidenburg:1982:EBB

[FB82] Michael E. Fraidenburg and Gregory G. Bargmann. Estimating boat-based fishing effort in a marine recreational fishery. North American Journal of Fisheries Management, 2(4):351–358, October 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ferguson:1986:CRD

[Fer86] Maury Osborn Ferguson. Characteristics of red drum and spotted seatrout commercial fishermen in Texas. North American Journal of Fisheries Management, 6(3):344–358, July 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Foucher:1982:DPC

[FF82] R. P. Foucher and D. Fournier. Derivation of Pacific cod age composition using length-frequency analysis. North American Journal of Fisheries Management, 2(3):276–284, July 1982. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fraley:1986:ETH

[FGC86] J. J. Fraley, M. A. Gaub, and J. R. Cavigli. Emergence trap and holding bottle for the capture of salmonid fry in streams. North American Journal of Fisheries Management, 6(1):119–121, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ferguson:1984:EAP

[FGM84] Maury O. Ferguson, Albert W. Green, and Gary C. Matlock. Evaluation of the accuracy and precision of volunteered size data from tagged red drum returns. North American Journal of Fisheries Management, 4(2):181–185, April 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Freeze:1982:DSB

[FH82] Mike Freeze and Scott Henderson. Distribution and status of the bighead carp and silver carp in Arkansas. North American Journal of Fisheries Management, 2(2):197–200, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fletcher:1987:RCW

[FHB87] Douglas H. Fletcher, Frank Haw, and Peter K. Bergman. Retention of coded wire tags implanted into cheek musculature of largemouth bass. North American Journal of Fisheries Management, 7(3):436– 439, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fritz:1987:SFD

[FJ87] Kenneth R. Fritz and David L. Johnson. Survival of freshwater drums released from Lake Erie commercial shore seines. North American Journal of Fisheries Management, 7(2):293–298, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Feathers:1983:EDU

[FK83] Mark G. Feathers and Anthony E. Knable. Effects of depressurization upon largemouth bass. North American Journal of Fisheries Management, 3(1):86–90, January 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ferry:1987:ETF

[FK87] Roland E. Ferry and Christopher C. Kohler. Effects of trap fishing on fish populations inhabiting a fringing coral reef. North American Journal of Fisheries Management, 7(4):580–588, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fraidenburg:1985:WCS

[FL85] Michael E. Fraidenburg and Richard H. Lincoln. Wild Chinook salmon management: an international conservation challenge. North American Journal of Fisheries Management, 5(3A):311–329, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fletcher:1987:HMW

[Fle87] Douglas H. Fletcher. Hooking mortality of walleyes captured in porcupine Bay, Washington. North American Journal of Fisheries Management, 7(4):594–596, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Flick:1983:OLP

[Fli83] William A. Flick. Observations on loons as predators on brook trout and as possible transmitters of infectious pancreatic necrosis (IPN). North American Journal of Fisheries Management, 3(1): 95–96, January 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Franzin:1987:CFA

[FM87] W. G. Franzin and G. A. McFarlane. Comparison of floy anchor tags and fingerling tags for tagging white suckers. North American Journal of Fisheries Management, 7(2):307–309, April 1987. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Faler:1988:EVF

[FMW88] Micheal P. Faler, Lawrence M. Miller, and Kurt I. Welke. Effects of variation in flow on distributions of northern squawfish in the Columbia River below McNary Dam. North American Journal of Fisheries Management, 8(1):30–35, February 1988. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Farman:1982:ELB

[FNN82] Richard S. Farman, Larry A. Nielsen, and Mitchell D. Norman. Estimating largemouth bass abundance using creel census and tournament data in the fishing-success method. North American Journal of Fisheries Management, 2(3):249–256, July 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

French:1982:RUS

[FNW82] Robert French, Russell Nelson, Jr., and Janet Wall. Role of the United States observer program in management of foreign fisheries in the Northeast Pacific Ocean and Eastern Bering Sea. North American Journal of Fisheries Management, 2(2):122–131, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fay:1985:FBS

[FP85] Clemon W. Fay and Garland B. Pardue. Freeze brands and submandibular latex injections as identifying marks on rainbow trout. North American Journal of Fisheries Management, 5(2B):248–251, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fraser:1989:ERP

[Fra89] J. M. Fraser. Establishment of reproducing populations of brook trout after stocking of interstrain hybrids in Precambrian shield lakes. North American Journal of Fisheries Management, 9(3): 352–363, August 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fonseca:1984:ISH

[FTCF84] Mark S. Fonseca, Gordon W. Thayer, Alexander J. Chester, and Cynthia Foltz. Impact of scallop harvesting on eelgrass (Zostera marina) meadows: Implications for management. North American Journal of Fisheries Management, 4(3):286–293, July 1984. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Griffith:1981:ESS

[GA81] J. S. Griffith and D. A. Andrews. Effects of a small suction dredge on fishes and aquatic invertebrates in Idaho streams. North American Journal of Fisheries Management, 1(1):21–28, January 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gatz:1987:ERE

[GA87] A. John Gatz, Jr. and S. Marshall Adams. Effects of repeated electroshocking on growth of bluegill × green sunfish hybrids. North American Journal of Fisheries Management, 7(3):448–450, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gabelhouse:1984:LCS

[Gab84] Donald W. Gabelhouse, Jr. A length-categorization system to assess fish stocks. North American Journal of Fisheries Management, 4(3):273–285, July 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gabelhouse:1987:RLB

[Gab87] Donald W. Gabelhouse, Jr. Responses of largemouth bass and bluegills to removal of surplus largemouth bass from a Kansas pond. North American Journal of Fisheries Management, 7(1): 81–90, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gilderhus:1986:PRP

[GAD86] P. A. Gilderhus, J. L. Allen, and V. K. Dawson. Persistence of rotenone in ponds at different temperatures. North American Journal of Fisheries Management, 6(1):129–130, January 1986. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Garman:1982:IIP

[Gar82] Greg C. Garman. Identification of ingested prey fish based on scale characteristics. North American Journal of Fisheries Management, 2(2):201–203, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gilderhus:1983:SCC

[GB83] P. A. Gilderhus and R. M. Burress. Selective control of common carp: Ineffectiveness of 2-(Digeranylamino)-ethanol (GD-174) in pond trials. North American Journal of Fisheries Management, 3(1):61–66, January 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goodyear:1984:ADI

[GC84] C. Phillip Goodyear and Sigurd W. Christensen. On the ability to detect the influence of spawning stock on recruitment. North American Journal of Fisheries Management, 4(2):186–193, April 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Graefe:1986:BOF

[GD86] Alan R. Graefe and Robert B. Ditton. Bay and offshore fishing in the Galveston Bay area: a comparative study of fishing patterns, fishermen characteristics, and expenditures. North American Journal of Fisheries Management, 6(2):192–199, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grant:1986:AMA

[GEB86] James W. A. Grant, John Englert, and Brian F. Bietz. Application of a method for assessing the impact of watershed practices: Effects of logging on salmonid standing crops. North American Journal of Fisheries Management, 6(1):24–31, January 1986. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Geihsler:1983:SFP

[GH83] Michael R. Geihsler and Daniel R. Holder. Status of fish populations in Georgia ponds 1-4 years after stocking. North American Journal of Fisheries Management, 3(2):189–196, April 1983. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gerhardt:1989:ECB

[GH89] David R. Gerhardt and Wayne A. Hubert. Effect of cheese bait on seasonal catches of channel catfish in hoop nets. North American Journal of Fisheries Management, 9(3):377–379, August 1989. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gilderhus:1982:EAP

[Gil82] Philip A. Gilderhus. Effects of an aquatic plant and suspended clay on the activity of fish toxicants. North American Journal of Fisheries Management, 2(4):301–306, July 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gilderhus:1989:EBA

[Gil89] Philip A. Gilderhus. Efficacy of benzocaine as an anesthetic for sahnonid fishes. North American Journal of Fisheries Management, 9(2):150–153, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gislason:1985:AIA

[Gis85] Jeffrey C. Gislason. Aquatic insect abundance in a regulated stream under fluctuating and stable diel flow patterns. North American Journal of Fisheries Management, 5(1):39–46, January 1985. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gitschlag:1986:MPS

[Git86] Gregg R. Gitschlag. Movement of pink shrimp in relation to the Tortugas Sanctuary. North American Journal of Fisheries Management, 6(3):328–338, July 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gatz:1986:ERE

[GLC86] A. John Gatz, Jr., James M. Loar, and Glenn F. Cada. Effects of repeated electroshocking on instantaneous growth of trout. North American Journal of Fisheries Management, 6(2):176–182, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Glass:1985:CHS

[GM85] Randal D. Glass and O. Eugene Maughan. Concentrated harvest of striped bass × white bass hybrids near a heated water outlet. North American Journal of Fisheries Management, 5(1):105–107, January 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gilderhus:1987:CEA

[GM87] Philip A. Gilderhus and Leif L. Marking. Comparative efficacy of 16 anesthetic chemicals on rainbow trout. North American Journal of Fisheries Management, 7(2):288–292, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gowing:1986:SGM

[Gow86] Howard Gowing. Survival and growth of matched plantings of Assinica strain brook trout and hybrid brook trout (Assinica male × domestic female) in six small Michigan lakes. North American Journal of Fisheries Management, 6(2):242–251, April 1986. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Guthrie:1988:EEL

[GP88] Ian C. Guthrie and Randall M. Peterman. Economic evaluation of lake enrichment strategies for British Columbia sockeye salmon.

North American Journal of Fisheries Management, 8(4):442–454, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Giorgi:1987:EDP

[GS87] Albert E. Giorgi and Carl W. Sims. Estimating the daily passage of juvenile salmonids at McNary Dam on the Columbia River. North American Journal of Fisheries Management, 7(2):215–222, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gagen:1989:PAG

[GSDK89] Charles J. Gagen, William E. Sharpe, David R. Dewalle, and William G. Kimmel. Pumping alkaline groundwater to Restore a put-and-take trout fishery in a stream acidified by atmospheric deposition. North American Journal of Fisheries Management, 9 (1):92–100, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gabriel:1989:ASS

[GSO89] W. L. Gabriel, M. P. Sissenwine, and W. J. Overholtz. Analysis of spawning stock biomass per recruit: An example for Georges Bank haddock. North American Journal of Fisheries Management, 9(4): 383–391, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Giorgi:1988:SCS

[GSZ^{+}88] Albert E. Giorgi, George A. Swan, Waldo S. Zaugg, Travis Coley, and Theresa Y. Barila. Susceptibility of Chinook salmon smolts to bypass systems at hydroelectric dams. North American Journal of Fisheries Management, 8(1):25–29, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gustafson:1988:MBA

[Gus88] Kent A. Gustafson. Management briefs: Approximating confidence intervals for indices of fish population size structure. North American Journal of Fisheries Management, 8(1):139–141, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:1983:SSM

[HAD83] Edward A. Hansen, Gaylord R. Alexander, and William H. Dunn. Sand sediment in a Michigan trout stream. Part I. A technique for removing sand bedload from streams. North American Journal of Fisheries Management, 3(4):355–364, October 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hall:1986:ECP

[Hal86] Thomas J. Hall. Electrofishing catch per hour as an indicator of largemouth bass density in Ohio impoundments. North American Journal of Fisheries Management, 6(3):397–400, July 1986. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hartzler:1983:EHL

[Har83] James R. Hartzler. The effects of half-log covers on angler harvest and standing crop of brown trout in McMichaels Creek, Pennsylvania. North American Journal of Fisheries Management, 3(3): 228–238, July 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harvey:1986:ESG

[Har86] Bret C. Harvey. Effects of suction gold dredging on fish and invertebrates in two California streams. North American Journal of Fisheries Management, 6(3):401–409, July 1986. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hauber:1983:TME

[Hau83] Alan B. Hauber. Two methods for evaluating fingerling walleye stocking success and natural year-class densities in Seven Island Lake, Wisconsin, 1977–1981. North American Journal of Fisheries Management, 3(2):152–155, April 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hudy:1983:PTS

[HB83] Mark Hudy and Charles R. Berry, Jr. Performance of three strains of rainbow trout in a Utah reservoir. North American Journal of Fisheries Management, 3(2):136–141, April 1983. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harris:1985:SDS

[HB85a] Charles C. Harris and Eric P. Bergersen. Survey on demand for sport fisheries: Problems and potentialities for its use in fishery management planning. North American Journal of Fisheries Management, 5(3A):400–410, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

House:1985:EIE

[HB85b] Robert A. House and Paul L. Boehne. Evaluation of instream enhancement structures for salmonid spawning and rearing in a coastal Oregon stream. North American Journal of Fisheries Management, 5(2B):283–295, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

House:1986:EIS

[HB86] Robert A. House and Paul L. Boehne. Effects of instream structures on salmonid habitat and populations in Tobe Creek, Oregon. North American Journal of Fisheries Management, 6(1):38–46, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holland-Bartels:1989:EWT

[HBDZ89] Leslie E. Holland-Bartels, Michael R. Dewey, and Steven J. Zigler. Effects of water temperature on the mortality of field-collected fish marked with fluorescent pigment. North American Journal of Fisheries Management, 9(3):341–344, August 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hanson:1983:EMP

[HBS83] David A. Hanson, Brian J. Belonger, and Dean L. Schoenike. Evaluation of a mechanical population reduction of black crappie and black bullheads in a small Wisconsin lake. North American Journal of Fisheries Management, 3(1):41–47, January 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heidinger:1988:UCW

[HC88] Roy C. Heidinger and S. Bradford Cook. Use of coded wire tags for marking fingerling fishes. North American Journal of Fisheries Management, 8(2):268–272, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hepworth:1987:IBE

[HD87] Dale K. Hepworth and Daniel J. Duffield. Interactions between an exotic crayfish and stocked rainbow trout in Newcastle Reservoir, Utah. North American Journal of Fisheries Management, 7 (4):554–561, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Healey:1985:IFP

[Hea85] M. C. Healey. Influence of fishermen’s preferences on the success of commercial fishery management regimes. North American Journal of Fisheries Management, 5(2A):173–180, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hegen:1985:URP

[Heg85] H. E. Hegen. Use of rotenone and potassium permanganate in estuarine sampling. North American Journal of Fisheries Management, 5(3B):500–502, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heggenes:1987:RSC

[Heg87] Jan Heggenes. Random, stratified creel surveys in three Norwegian rivers with low fishing intensities. North American Journal of Fisheries Management, 7(3):363–368, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hsiao:1987:THE

[HEJ87] Yu-Mong Hsiao, J. E. Easley, Jr., and Thomas Johnson. Testing for harmful effects of clam and scallop harvesting techniques in the North Carolina bay scallop fishery. North American Journal of Fisheries Management, 7(2):187–193, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hemmingsen:1986:CAR

[HES86] A. R. Hemmingsen, R. D. Ewing, and R. G. Sheldon. Comparison of adult returns to a hatchery from subyearling and yearling Coho salmon released at similar sizes and different times. North American Journal of Fisheries Management, 6(2):204–208, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heimer:1985:PSP

[HFG85] John T. Heimer, W. M. Frazier, and J. S. Griffith. Post-stocking performance of catchable-size hatchery rainbow trout with and without pectoral fins. North American Journal of Fisheries Management, 5(1):21–25, January 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hesthagen:1986:MAS

[HG86] Trygve Hesthagen and Erik Garnås. Migration of Atlantic salmon smolts in River Orkla of Central Norway in relation to management of a hydroelectric station. North American Journal of Fisheries Management, 6(3):376–382, July 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hightower:1982:SCS

[HGH82] Joseph E. Hightower, Ronnie J. Gilbert, and Timmy B. Hess. Shoreline and cove sampling to estimate survival of young largemouth bass in Lake Oconee, Georgia. North American Journal of Fisheries Management, 2(3):257–261, July 1982. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heidinger:1983:OCT

[HHHH83] R. C. Heidinger, D. R. Helms, T. I. Hiebert, and P. H. Howe. Operational comparison of three electrofishing systems. North American Journal of Fisheries Management, 3(3):254–257, July 1983. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hill:1986:MSS

[Hil86] Kay R. Hill. Mortality and standing stocks of grass carp planted in two Iowa lakes. North American Journal of Fisheries Management, 6(3):449–451, July 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hilborn:1987:LUR

[Hil87] Ray Hilborn. Living with uncertainty in resource management. North American Journal of Fisheries Management, 7(1):1–5, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hilborn:1989:YES

[Hil89] Ray Hilborn. Yield estimation for spatially connected populations: An example of surface and longline fisheries for yellowfin tuna. North American Journal of Fisheries Management, 9(4):402–410, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hassler:1988:RMI

[HK88] Thomas J. Hassler and Stephen T. Kucas. Returns of morpholineimprinted Coho salmon to the Mad River, California. North American Journal of Fisheries Management, 8(3):356–358, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Howell:1987:CTD

[HL87] W. Huntting Howell and Richard Langan. Commercial trawler discards of four flounder species in the Gulf of Maine. North American Journal of Fisheries Management, 7(1):6–17, January 1987. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Havey:1981:CFM

[HLS81] Keith A. Havey, David O. Locke, and C. Leslie Smith. A comparison of four methods of estimating populations of hatchery-reared brook trout at Tomah Lake, Maine. North American Journal of Fisheries Management, 1(2):186–192, April 1981. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harberg:1985:FBF

[HM85] Mark C. Harberg and Timothy Modde. Feeding behavior, food consumption, growth, and survival of hybrid grass carp in two South Dakota ponds. North American Journal of Fisheries Management, 5(3B):457–464, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hampton:1986:CSF

[HM86] J. Hampton and J. Majkowski. Computer simulations of future southern bluefin tuna parental biomass, recruitment, and catches under the 1982 fishing regime. North American Journal of Fisheries Management, 6(1):77–87, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heifetz:1986:ELW

[HMK86] Jonathan Heifetz, Michael L. Murphy, and K. V. Koski. Effects of logging on winter habitat of juvenile salmonids in Alaskan streams. North American Journal of Fisheries Management, 6(1):52–58, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Horns:1989:IMQ

[HMK89] William H. Horns, J. Ellen Marsden, and Charles C. Krueger. Inexpensive method for quantitative assessment of lake trout egg deposition. North American Journal of Fisheries Management, 9(3):

280–286, August 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hesse:1982:EFS

[HN82a] Larry W. Hesse and Brad A. Newcomb. Effects of flushing Spencer hydro on water quality, fish, and insect fauna in the Niobrara River, Nebraska. North American Journal of Fisheries Management, 2(1): 45–52, January 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hesse:1982:EAF

[HN82b] Larry W. Hesse and Brad A. Newcomb. On estimating the abundance of fish in the Upper Channelized Missouri River. North American Journal of Fisheries Management, 2(1):80–83, January 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoffman:1986:HIP

[Hof86] Ray J. Hoffman. A horizontal intragravel pipe for sampling water quality in salmonid spawning gravel. North American Journal of Fisheries Management, 6(3):445–448, July 1986. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holland:1987:EBN

[Hol87] Leslie E. Holland. Effect of brief navigation-related dewaterings on fish eggs and larvae. North American Journal of Fisheries Management, 7(1):145–147, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harshbarger:1982:ESF

[HP82] T. J. Harshbarger and P. E. Porter. Embryo survival and fry emergence from two methods of planting brown trout eggs. North American Journal of Fisheries Management, 2(1):84–89, January 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hume:1988:EST

[HP88] Jeremy M. B. Hume and Eric A. Parkinson. Effects of size at and time of release on the survival and growth of steelhead fry stocked in streams. North American Journal of Fisheries Management, 8 (1):50–57, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hassemer:1981:ODS

[HR81] P. F. Hassemer and B. E. Rieman. Observations of deep-spawning Kokanee on artificially created spawning habitat. North American Journal of Fisheries Management, 1(2):173–176, April 1981. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hubert:1989:RPH

[HR89] Wayne A. Hubert and Frank J. Rahel. Relations of physical habitat to abundance of four nongame fishes in high-plains streams: a test of habitat suitability index models. North American Journal of Fisheries Management, 9(3):332–340, August 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Huson:1984:IVM

[HRDM84] R. M. Huson, D. Rivard, W. G. Doubleday, and W. D. McKone. Impact of varying mesh size and depth of fishing on the financial performance of an integrated harvesting/processing operation for redfish in the Northwest Atlantic. North American Journal of Fisheries Management, 4(1):32–47, January 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hubert:1983:PVG

[HS83] Wayne A. Hubert and Mark B. Sandheinrich. Patterns of variation in gill-net catch and diet of yellow perch in a stratified Iowa lake. North American Journal of Fisheries Management, 3(2):156–162, April 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haegele:1985:EEN

[HS85] C. W. Haegele and J. F. Schweigert. Estimation of egg numbers in Pacific herring spawns on giant kelp. North American Journal of Fisheries Management, 5(1):65–71, January 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hitchins:1986:SRI

[HS86] John R. Hitchins and Wilson G. A. Samis. Successful reproduction by introduced lake trout in 10 Northeastern Ontario lakes. North American Journal of Fisheries Management, 6(3):372–375, July 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hanson:1989:DMR

[HS89] Lee H. Hanson and William D. Swink. Downstream migration of recently metamorphosed sea lampreys in the Ocqueoc River, Michigan, before and after treatment with lampricides. North American Journal of Fisheries Management, 9(3):327–331, August 1989. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoffnagle:1989:AGC

[HT89] Timothy L. Hoffnagle and Tom J. Timmons. Age, growth, and catch analysis of the commercially exploited paddlefish population in Kentucky Lake, Kentucky-Tennessee. North American Journal of Fisheries Management, 9(3):316–326, August 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hubert:1988:MBA

[Hub88] Wayne A. Hubert. Management briefs: Altitude as the determinant of distribution of largemouth bass and smallmouth bass in Wyoming. North American Journal of Fisheries Management, 8 (3):386–387, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hudy:1985:RTB

[Hud85] Mark Hudy. Rainbow trout and brook trout mortality from high voltage AC electrofishing in a controlled environment. North American Journal of Fisheries Management, 5(3B):475–479, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Huang:1983:CAP

[HW83] Bin Huang and Carl J. Walters. Cohort analysis and population dynamics of large yellow croaker in the China Sea. North American Journal of Fisheries Management, 3(3):295–305, July 1983. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hilborn:1984:MFC

[HWPS84] Ray Hilborn, Carl J. Walters, Randall M. Peterman, and Michael J. Staley. Models and fisheries: a case study in implementation. North American Journal of Fisheries Management, 4(1):9–14, January 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Iversen:1984:ARU

[IB84] Edwin S. Iversen and Scott P. Bannerot. Artificial reefs under marina docks in Southern Florida. North American Journal of Fisheries Management, 4(3):294–299, July 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isbell:1989:RGN

[IR89] Gary L. Isbell and Michael R. Rawson. Relations of gill-net catches of walleyes and angler catch rates in Ohio waters of Western Lake Erie. North American Journal of Fisheries Management, 9(1):41– 46, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jacobson:1989:NAM

[Jac89] Lawrence D. Jacobson. New approach to measuring the frequency of sea lamprey wounds in fish stocks in the Great Lakes. North American Journal of Fisheries Management, 9(1):23–34, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jackson:1988:EFI

[JD88] Donald C. Jackson and William D. Davies. Environmental factors influencing summer angler effort on the Jordan Dam tailwater, Alabama. North American Journal of Fisheries Management, 8(3): 305–309, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jahn:1987:ESH

[JDTK87] Larry A. Jahn, David R. Douglas, Monte J. Terharr, and Glen W. Kruse. Effects of stocking hybrid striped bass in Spring Lake, Illinois. North American Journal of Fisheries Management, 7(4):522– 530, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jensen:1985:TSA

[Jen85] A. L. Jensen. Time series analysis and the forecasting of menhaden catch and CPUE. North American Journal of Fisheries Management, 5(1):78–85, January 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jensen:1986:AML

[Jen86] A. L. Jensen. Assessment of the Maine lobster fishery with surplus production models. North American Journal of Fisheries Management, 6(1):63–68, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jennings:1989:MBU

[Jen89] Mark R. Jennings. Management briefs: Use of spring scales for weighing live fish in the field. North American Journal of Fisheries Management, 9(4):509–511, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jewett:1986:MSS

[Jew86] Stephen C. Jewett. Mortality of spot shrimp associated with streamer tags. North American Journal of Fisheries Management, 6(2):260–263, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

James:1987:DOE

[JLZM87] Paul W. James, Stuart C. Leon, Alexander V. Zale, and O. Eugene Maughan. Diver-operated electrofishing device. North American Journal of Fisheries Management, 7(4):597–598, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1987:CEM

[Joh87] Russell J. Johnson. Comparative efficiencies of monofilament and multifilament drops used on trotlines for catfish. North American Journal of Fisheries Management, 7(1):156–158, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jones:1982:TSR

[Jon82] Albert R. Jones. The two-story rainbow trout fishery at Laurel River Lake, Kentucky. North American Journal of Fisheries Management, 2(2):132–137, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jensen:1988:EPF

[JRK88] A. L. Jensen, R. H. Reider, and W. P. Kovalak. Estimation of production forgone. North American Journal of Fisheries Management, 8(2):191–198, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jacobs:1982:EFP

[JS82] Kenneth E. Jacobs and William D. Swink. Estimations of fish population size and sampling efficiency of electrofishing and rotenone in two Kentucky tailwaters. North American Journal of Fisheries Management, 2(3):239–248, July 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jacobs:1983:FAP

[JS83] Kenneth E. Jacobs and William D. Swink. Fish abundance and population stability in a reservoir tailwater and an unregulated headwater stream. North American Journal of Fisheries Management, 3(4):395–402, October 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:1988:HPS

[JSC88] Barry L. Johnson, David L. Smith, and Robert F. Carline. Habitat preferences, survival, growth, foods, and harvests of walleyes and walleye × sanger hybrids. North American Journal of Fisheries Management, 8(3):292–304, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jacobs:1987:MTF

[JSN87] Kenneth E. Jacobs, William D. Swink, and Jerry F. Novotny. Minimum tailwater flows in relation to habitat suitability and sportfish harvest. North American Journal of Fisheries Management, 7 (4):569–574, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jacobson:1985:SHS

[JT85] Peter C. Jacobson and William W. Taylor. Simulation of harvest strategies for a fluctuating population of lake whitefish. North American Journal of Fisheries Management, 5(4):537–546, October 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kallemeyn:1987:CRL

[Kal87] Larry W. Kallemeyn. Correlations of regulated lake levels and climatic factors with abundance of young-of-the-year walleye and yellow perch in four lakes in Voyageurs National Park. North American Journal of Fisheries Management, 7(4):513–521, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kallemeyn:1989:LCT

[Kal89] Larry W. Kallemeyn. Loss of carlin tags from walleyes. North American Journal of Fisheries Management, 9(1):112–115, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keller:1982:RFG

[KB82] Chares R. Keller and Kenneth P. Burnham. Riparian fencing, grazing, and trout habitat preference on Summit Creek, Idaho. North American Journal of Fisheries Management, 2(1):53–59, January 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kimura:1985:BME

[KB85] Daniel K. Kimura and James W. Balsiger. Bootstrap methods for evaluating sablefish pot index surveys. North American Journal of Fisheries Management, 5(1):47–56, January 1985. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Knudsen:1987:ERB

[KD87] E. Eric Knudsen and Stephen J. Dilley. Effects of riprap bank reinforcement on juvenile salmonids in four Western Washington streams. North American Journal of Fisheries Management, 7(3): 351–356, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kirk:1986:RSM

[KDP86] James P. Kirk, William D. Davies, and Kyuseok Park. Response of some members of the fish community to gizzard shad removal from Chambers County Public Fishing Lake, Alabama. North American Journal of Fisheries Management, 6(2):252–255, April 1986. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kendall:1988:TCN

[Ken88] Robert L. Kendall. Taxonomic changes in North American trout names. North American Journal of Fisheries Management, 8(4): 389, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kendall:1989:NEP

[Ken89] Robert L. Kendall. New editorial procedures for Management. North American Journal of Fisheries Management, 9(4):381–382, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kurzawski:1982:CSP

[KH82] Kenneth F. Kurzawski and Roy C. Heidinger. The cyclic stocking of parentals in a farm pond to produce a population of male bluegill × female green sunfish F 1 hybrids and male redear sunfish x female green sunfish F_{1} hybrids. North American Journal of Fisheries Management, 2(2):188–192, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kozel:1989:THA

[KH89] Steven J. Kozel and Wayne A. Hubert. Testing of habitat assessment models for small trout streams in the Medicine Bow National Forest, Wyoming. North American Journal of Fisheries Management, 9(4):458–464, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kimura:1988:ARA

[Kim88] Daniel K. Kimura. Analyzing relative abundance indices with loglinear models. North American Journal of Fisheries Management, 8(2):175–180, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kwain:1984:RYO

[KK84] W. Kwain and S. J. Kerr. Return of 1-year-old pink salmon in Michipicoten River, Eastern Lake Superior. North American Journal of Fisheries Management, 4(3):335–337, July 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kucera:1985:ILC

[KKM85] Paul A. Kucera, David L. Koch, and Gerald F. Marco. Introductions of Lahontan cutthroat trout into Omak Lake, Washington. North American Journal of Fisheries Management, 5(2B):296–301, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Knight:1982:ESF

[KM82] Roger L. Knight and F. Joseph Margraf. Estimating stomach fullness in fishes. North American Journal of Fisheries Management, 2(4):413–414, October 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Klima:1986:STP

[KMP86] Edward F. Klima, Geoffrey A. Matthews, and Frank J. Patella. Synopsis of the Tortugas pink shrimp fishery, 1960–1983, and the impact of the Tortugas Sanctuary. North American Journal of Fisheries Management, 6(3):301–310, July 1986. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Killgore:1989:DAF

[KMR89] K. Jack Killgore, Raymond P. Morgan II, and Nancy B. Rybicki. Distribution and abundance of fishes associated with submersed aquatic plants in the Potomac River. North American Journal of Fisheries Management, 9(1):101–111, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Klassen:1988:UGW

[KN88] H. D. Klassen and T. G. Northcote. Use of gabion weirs to improve spawning habitat for pink salmon in a small logged watershed. North American Journal of Fisheries Management, 8(1):36– 44, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Konikoff:1988:IYC

[KNR88] Mark Konikoff, Scott H. Newton, and Walter R. Robison. Increasing yields of caged channel catfish in a farm pond by sequential harvests, with notes on a supplemental harvest of blue tilapia. North American Journal of Fisheries Management, 8(1):91–97, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kushneriuk:1984:IFT

[KP84] Robert S. Kushneriuk and Jyri E. Paloheimo. An inexpensive flyscreen trap for sampling small fish. North American Journal of Fisheries Management, 4(2):222–226, April 1984. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Klar:1986:MFS

[KP86] Gerald T. Klar and Nick C. Parker. Marking fingerling striped bass and blue tilapia with coded wire tags and microtaggants. North American Journal of Fisheries Management, 6(3):439–444, July 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Knudsen:1989:EFC

[KPR^{+}89] E. Eric Knudsen, Ronald F. Paille, Barton D. Rogers, William H. Herke, and James P. Geaghan. Effects of a fixed-crest weir on brown shrimp Penaeus aztecus growth, mortality, and emigration in a Louisiana coastal marsh. North American Journal of Fisheries Management, 9(4):411–419, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Klima:1986:OTS

[KRJ86] Edward F. Klima, Terrell W. Roberts, and Albert C. Jones. Overview of the Tortugas Sanctuary studies. North American Journal of Fisheries Management, 6(3):297–300, July 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kelso:1989:VRM

[KS89] J. R. M. Kelso and B. J. Shuter. Validity of the removal method for fish population estimation in a small lake. North American Journal of Fisheries Management, 9(4):471–476, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kwain:1984:FES

[KT84] Wen-Hwa Kwain and Evan Thomas. The first evidence of spring spawning by Chinook salmon in Lake Superior. North American Journal of Fisheries Management, 4(2):227–228, April 1984. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kwain:1987:RSF

[Kwa87] W. Kwain. Record size of freshwater pink salmon. North American Journal of Fisheries Management, 7(2):302–303, April 1987. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Laarman:1981:VSM

[Laa81] Percy W. Laarman. Vital statistics of a Michigan fish population, with special emphasis on the effectiveness of stocking 15-cm walleye fingerlings. North American Journal of Fisheries Management, 1(2):177–185, April 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Light:1983:EGL

[LAA83] Robert W. Light, Peter H. Adler, and Dean E. Arnold. Evaluation of gastric lavage for stomach analyses. North American Journal of Fisheries Management, 3(1):81–85, January 1983. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Loomis:1987:AMP

[LD87] David K. Loomis and Robert B. Ditton. Analysis of motive and participation differences between saltwater sport and tournament fishermen. North American Journal of Fisheries Management, 7 (4):482–487, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Loomis:1988:TPF

[LD88] David K. Loomis and Robert B. Ditton. Technique for projecting the future growth and distribution of marine recreational fishing demand. North American Journal of Fisheries Management, 8(2): 259–263, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lynch:1989:MBQ

[LGJR89] William E. Lynch, Jr., Jerald M. Gerber, David L. Johnson, and Jeffrey R. Reutter. Management brief: a quickly deployed vertical gill-net system. North American Journal of Fisheries Management, 9(1):119–121, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lynch:1989:IIS

[LJ89] William E. Lynch, Jr. and David L. Johnson. Influences of interstice size, shade, and predators on the use of artificial structures by bluegills. North American Journal of Fisheries Management, 9(2): 219–225, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lynch:1982:SGF

[LJS82] William E. Lynch, Jr., David L. Johnson, and Scott A. Schell. Survival, growth, and food habits of walleye × sauger hybrids (saugeye) in ponds. North American Journal of Fisheries Management, 2(4):381–387, October 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lloyd:1987:ETF

[LKL87] Denby S. Lloyd, Jeffrey P. Koenings, and Jacqueline D. Laperriere. Effects of turbidity in fresh waters of Alaska. North American Journal of Fisheries Management, 7(1):18–33, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lloyd:1987:TWQ

[Llo87] Denby S. Lloyd. Turbidity as a water quality standard for salmonid habitats in Alaska. North American Journal of Fisheries Management, 7(1):34–45, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lorson:1987:UTM

[LM87] Richard D. Lorson and Vincent A. Mudrak. Use of tetracycline to mark otoliths of American shad fry. North American Journal of Fisheries Management, 7(3):453–455, July 1987. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Layher:1987:SBH

[LMW87] William G. Layher, O. Eugene Maughan, and W. D. Warde. Spotted bass habitat suitability related to fish occurrence and biomass and measurements of physicochemical variables. North American Journal of Fisheries Management, 7(2):238–251, April 1987. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leonard:1988:UHG

[LO88] Paul M. Leonard and Donald J. Orth. Use of habitat guilds of fishes to determine instream flow requirements. North American Journal of Fisheries Management, 8(4):399–409, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Low:1981:MRM

[Low81] R. A. Low, Jr. Mortality rates and management strategies for black sea bass off the Southeast Coast of the United States. North American Journal of Fisheries Management, 1(2):93–103, April 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lotspeich:1982:ILA

[LP82] Frederick B. Lotspeich and William S. Platts. An integrated landaquatic classification system. North American Journal of Fisheries Management, 2(2):138–149, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Laarman:1982:RSS

[LR82] Percy W. Laarman and James R. Ryckman. Relative size selectivity of trap nets for eight species of fish. North American Journal of Fisheries Management, 2(1):33–37, January 1982. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lopez-Rojas:1984:DOS

[LRLM84] Hector Lopez-Rojas, John G. Lundberg, and Edie Marsh. Design and operation of a small trawling apparatus for use with dugout canoes. North American Journal of Fisheries Management, 4(3): 331–334, July 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lanigan:1989:PSS

[LT89] Steven H. Lanigan and Harold M. Tyus. Population size and status of the razorback sucker in the Green River Basin, Utah and Colorado. North American Journal of Fisheries Management, 9 (1):68–73, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Low:1985:ACP

[LUBO85] R. A. Low, Jr., G. F. Ulrich, C. A. Barans, and D. A. Oakley. Analysis of catch per unit of effort and length composition in the South Carolina commercial handline fishery, 1976–1982. North American Journal of Fisheries Management, 5(3A):340–363, July 1985. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lam:1989:MWS

[LWL89] C. F. Lam, J. D. Whitaker, and F. S. Lee. Model for white shrimp landings for the Central Coast of South Carolina. North American Journal of Fisheries Management, 9(1):12–22, February 1989. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lyons:1986:CEB

[Lyo86] John Lyons. Capture efficiency of a beach seine for seven freshwater fishes in a north-temperate lake. North American Journal of Fisheries Management, 6(2):288–289, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Macritchie:1984:IPI

[MA84] Ian C. Macritchie and Edward R. Armstrong. An incentive program to increase angler tag returns. North American Journal of Fisheries Management, 4(1):129–130, January 1984. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mackenzie:1982:CIP

[Mac82] Clyde L. Mackenzie, Jr. Compatibility of invertebrate populations and commercial fishing for ocean quahogs. North American Journal of Fisheries Management, 2(3):270–275, July 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:1988:SGP

[Mac88] Michael J. Maceina. Simple grinding procedure to section otoliths. North American Journal of Fisheries Management, 8(1):141–143, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Margenau:1987:VRT

[Mar87] Terry L. Margenau. Vulnerability of radio-tagged northern pike to angling. North American Journal of Fisheries Management, 7 (1):158–159, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Matlock:1983:AOT

[Mat83] Gary C. Matlock. Angler origin and their catch retention rate while fishing on seven Texas Bay systems. North American Journal of Fisheries Management, 3(4):438–444, October 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Matlock:1986:ENR

[Mat86] Gary C. Matlock. Estimate of the number of red drum anglers in Texas. North American Journal of Fisheries Management, 6(2): 292–294, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Matkowski:1989:DST

[Mat89] Shelley M. D. Matkowski. Differential susceptibility of three species of stocked trout to bird predation. North American Journal of Fisheries Management, 9(2):184–187, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marking:1981:SFS

[MB81] Leif L. Marking and Terry D. Bills. Sensitivity of four species of carp to selected fish toxicants. North American Journal of Fisheries Management, 1(1):51–54, January 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McFarlane:1986:TSA

[MB86] Gordon A. McFarlane and Richard J. Beamish. A tag suitable for assessing long-term movements of spiny dogfish and preliminary results from use of this tag. North American Journal of Fisheries Management, 6(1):69–76, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martinez:1989:PBM

[MB89] Patrick J. Martinez and Eric P. Bergersen. Proposed biological management of Mysis relicta in Colorado lakes and reservoirs. North American Journal of Fisheries Management, 9(1): 1–11, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marking:1983:CCF

[MBRG83] L. L. Marking, T. D. Bills, J. J. Rach, and S. J. Grabowski. Chemical control of fish and fish eggs in the Garrison Diversion Unit, North Dakota. North American Journal of Fisheries Management, 3(4):410–418, October 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McKay:1987:PSC

[McK87] Edward H. McKay. Portable, self-cleaning fish screen for low-flow water diversions. North American Journal of Fisheries Management, 7(4):603–605, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Margraf:1985:ISS

[MCS85] F. Joseph Margraf, David M. Chase, and Kirk Strawn. Intake screens for sampling fish populations: The size-selectivity problem. North American Journal of Fisheries Management, 5(2A):210–213, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marrin:1982:EAC

[ME82] Donn L. Marrin and Don C. Erman. Evidence against competition between trout and nongame fishes in Stampede Reservoir, California. North American Journal of Fisheries Management, 2(3): 262–269, July 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meador:1984:UPS

[MEB84] M. R. Meador, A. G. Eversole, and J. S. Bulak. Utilization of portions of the Santee River System by spawning blueback herring. North American Journal of Fisheries Management, 4(2):155–163, April 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meffe:1983:ACR

[Mef83] Gary K. Meffe. Attempted chemical renovation of an Arizona springbrook for management of the endangered Sonoran topminnow. North American Journal of Fisheries Management, 3(3):315– 321, July 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miller:1986:PSF

[MG86] Marc L. Miller and Richard P. Gale. Professional styles of Federal forest and marine fisheries resource managers. North American Journal of Fisheries Management, 6(2):141–148, April 1986. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:1988:ADB

[MG88] Michael J. Maceina and Ira F. Greenbaum. Allozymic differences between black and white crappies. North American Journal of Fisheries Management, 8(1):123–126, February 1988. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mills:1987:UZS

[MGS87] Edward L. Mills, David M. Green, and Albert Schiavone, Jr. Use of zooplankton size to assess the community structure of fish populations in freshwater lakes. North American Journal of Fisheries Management, 7(3):369–378, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Minns:1988:ENL

[MH88] Charles K. Minns and Donal A. Hurley. Effects of net length and set time on fish catches in gill nets. North American Journal of Fisheries Management, 8(2):216–223, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Myers:1987:SOC

[MHK^{+}87] Katherine W. Myers, Colin K. Harris, Curtis M. Knudsen, Robert V. Walker, Nancy D. Davis, and Donald E. Rogers. Stock origins of Chinook salmon in the area of the Japanese mothership salmon fishery. North American Journal of Fisheries Management, 7(4):459–474, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miller:1983:HMT

[Mil83] Robert J. Miller. How many traps should a crab fisherman fish? North American Journal of Fisheries Management, 3(1):1–8, January 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:1986:RSC

[Mir86] L. Esteban Miranda. Removal of stomach contents from live largemouth bass using hydrogen peroxide. North American Journal of Fisheries Management, 6(2):285–286, April 1986. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mitzner:1981:IFS

[Mit81] Larry Mitzner. Influence of floodwater storage on abundance of juvenile crappie and subsequent harvest at Lake Rathbun, Iowa. North American Journal of Fisheries Management, 1(1):46–50, January 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McMenemy:1988:UIP

[MK88] James R. McMenemy and Boyd Kynard. Use of inclined-plane traps to study movement and survival of Atlantic salmon smolts in the Connecticut River. North American Journal of Fisheries Management, 8(4):481–488, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murphy:1989:IDW

[MK89] Michael L. Murphy and K V. Koski. Input and depletion of woody debris in Alaska streams and implications for streamside management. North American Journal of Fisheries Management, 9(4): 427–436, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marsden:1989:IPO

[MKM89] J. Ellen Marsden, Charles C. Krueger, and Beanie May. Identification of parental origins of naturally produced lake trout in Lake Ontario: Application of mixed-stock analysis to a second generation. North American Journal of Fisheries Management, 9(3):257– 268, August 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marsh:1982:FPM

[MM82] Paul C. Marsh and W. L. Minckley. Fishes of the Phoenix metropolitan area in Central Arizona. North American Journal of Fisheries Management, 2(4):395–402, October 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Matylewich:1985:ERS

[MM85] M. A. Matylewich and P. R. Mundy. Evaluation of the relevance of some environmental factors to the estimation of migratory timing and yield for the brown shrimp of Pamlico Sound, North Carolina. North American Journal of Fisheries Management, 5(2A):197–209, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:1987:RYY

[MM87] L. Esteban Miranda and Robert J. Muncy. Recruitment of youngof-year largemouth bass in relation to size structure of parental stock. North American Journal of Fisheries Management, 7(1): 131–137, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mackenzie:1988:ESA

[MM88] Chet Mackenzie and John R. Moring. Estimating survival of Atlantic salmon during the intragravel period. North American Journal of Fisheries Management, 8(1):45–49, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mosindy:1987:IAP

[MMC87] Thomas E. Mosindy, Walter T. Momot, and Peter J. Colby. Impact of angling on the production and yield of mature walleyes and northern pike in a small boreal lake in Ontario. North American Journal of Fisheries Management, 7(4):493–501, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martin:1983:UIP

[MMS^{+}83] Roy W. Martin, John Myers, Stacia A. Sower, David J. Phillips, and Carlin McAuley. Ultrasonic imaging, a potential tool for sex determination of live fish. North American Journal of Fisheries Management, 3(3):258–264, July 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moring:1982:EHS

[Mor82] John R. Moring. An efficient hatchery strain of rainbow trout for stocking Oregon streams. North American Journal of Fisheries Management, 2(3):209–215, July 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moring:1985:FSA

[Mor85a] John R. Moring. Formulatiug simple angler profiles to better manage catchable trout fisheries. North American Journal of Fisheries Management, 5(3A):413–414, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moring:1985:RAC

[Mor85b] John R. Moring. Relation of angler catch rates and effort to reductions in numbers of trout stocked in five Oregon streams. North American Journal of Fisheries Management, 5(4):575–579, October 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mosher:1983:EAC

[Mos83] Thomas D. Mosher. Effects of artificial circulation on fish distribution and angling success for channel catfish in a small prairie lake. North American Journal of Fisheries Management, 3(4):403– 409, October 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Merritt:1986:MTU

[MR86] M. F. Merritt and K. Roberson. Migratory timing of Upper Copper River sockeye salmon stocks and its implications for the regulation of the commercial fishery. North American Journal of Fisheries Management, 6(2):216–225, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McClendon:1987:PBV

[MR87] Danny D. McClendon and Charles F. Rabeni. Physical and biological variables useful for predicting population characteristics of smallmouth bass and rock bass in an Ozark stream. North American Journal of Fisheries Management, 7(1):46–56, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Myers:1988:SOC

[MR88] Katherine W. Myers and Donald E. Rogers. Stock origins of Chinook salmon in incidental catches by groundfish fisheries in the Eastern Bering Sea. North American Journal of Fisheries Management, 8(2):162–171, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moore:1983:SCB

[MRL83] Stephen E. Moore, Bromfield Ridley, and Gary L. Larson. Standing crops of brook trout concurrent with removal of rainbow trout from selected streams in Great Smoky Mountains National Park. North American Journal of Fisheries Management, 3(1):72–80, January 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mills:1982:EFC

[MS82] Edward L. Mills and Albert Schiavone, Jr. Evaluation of fish communities through assessment of zooplankton populations and measures of lake productivity. North American Journal of Fisheries Management, 2(1):14–27, January 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Milner:1985:FNU

[MS85a] Alexander Milner and Leigh Smith. Fyke nets used in a Southeastern Alaskan stream for sampling salmon fry and smolts. North American Journal of Fisheries Management, 5(3B):502–506, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Modde:1985:LGE

[MS85b] Timothy Modde and Charles G. Scalet. Latitudinal growth effects on predator–prey interactions between largemouth bass and bluegills in ponds. North American Journal of Fisheries Management, 5(2B):227–232, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Metzger:1986:UDC

[MS86] Randy J. Metzger and Paul L. Shafland. Use of detonating cord for sampling fish. North American Journal of Fisheries Management, 6(1):113–118, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:1984:EWL

[MSB84] L. Esteban Miranda, William L. Shelton, and Thomas D. Bryce. Effects of water level manipulation on abundance, mortality, and growth of young-of-year largemouth bass in West Point Reservoir, Alabama–Georgia. North American Journal of Fisheries Management, 4(3):314–320, July 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marinac-Sanders:1981:SBP

[MSC81] Patricia Marinac-Sanders and Daniel W. Coble. The smallmouth bass population and fishery in a Northern Wisconsin lake, with implications for other waters. North American Journal of Fisheries Management, 1(1):15–20, January 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McEachron:1986:ATD

[MSG86] Lawrence W. McEachron, Gary E. Saul, and Steven J. Gutreuter. Accuracy of trip durations reported by marine sport-boat fishermen. North American Journal of Fisheries Management, 6(1):59– 62, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moulton:1989:MBN

[MSK89] Peter L. Moulton, Stephen R. Saddlier, and Ian A. Knuckey. Management briefs: New time-at-liberty record set by tagged school shark Galeorhinus galeus caught off Southern Australia. North American Journal of Fisheries Management, 9(2):254–255, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Macdonald:1987:CRF

[MT87] Craig D. Macdonald and Bruce E. Thompson. Characteristics of a recreational fishery for Hawaiian spiny lobsters at the Midway Islands. North American Journal of Fisheries Management, 7(2): 194–201, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mundy:1982:CMT

[Mun82] P. R. Mundy. Computation of migratory timing statistics for adult Chinook salmon in the Yukon River, Alaska, and their relevance to fisheries management. North American Journal of Fisheries Management, 2(4):359–370, October 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murphy:1985:OPS

[Mur85] Michael L. Murphy. Die-offs of pre-spawn adult pink salmon and chum salmon in Southeastern Alaska. North American Journal of Fisheries Management, 5(2B):302–308, April 1985. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moyle:1983:RFP

[MVG83] Peter B. Moyle, Bruce Vondracek, and Gary D. Grossman. Responses of fish populations in the North Fork of the Feather River, California, to treatments with fish toxicants. North American Journal of Fisheries Management, 3(1):48–60, January 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muth:1986:CGM

[MW86] Kenneth M. Muth and David R. Wolfert. Changes in growth and maturity of walleyes associated with stock rehabilitation in Western Lake Erie, 1964–1983. North American Journal of Fisheries Management, 6(2):168–175, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martin:1986:IRV

[MWD86] Douglas J. Martin, Lawrence J. Wasserman, and Virginia H. Dale. Influence of riparian vegetation on posteruption survival of Coho salmon fingerlings on the west-side streams of Mount St. Helens, Washington. North American Journal of Fisheries Management, 6 (1):1–8, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Monk:1989:EFW

[MWTO89] Bruce Monk, Dick Weaver, Clark Thompson, and Frank Ossiander. Effects of flow and weir design on the passage behavior of American shad and salmonids in an experimental fish ladder. North American Journal of Fisheries Management, 9(1):60–67, February 1989. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morrison:1987:HEC

[MZ87] John Morrison and David Zajac. Histologic effect of coded wire tagging in chum salmon. North American Journal of Fisheries Management, 7(3):439–441, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neves:1985:APS

[NBH85] Richard J. Neves, Steven L. Brayton, and Louis A. Helfrich. Abundance and production of a self-sustaining population of rainbow trout in the South Fork Holston River, Virginia. North American Journal of Fisheries Management, 5(4):584–589, October 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Negus:1982:MAS

[Neg82] Mary T. Negus. Modified anchoring system for vertical gill nets. North American Journal of Fisheries Management, 2(4):412–413, October 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nesler:1986:PIF

[Nes86] Thomas P. Nesler. Prediction of ice-fishing harvest of quality-size lake trout from previous gill-net sampling. North American Journal of Fisheries Management, 6(2):277–281, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nester:1987:HIT

[Nes87] R. T. Nester. Horizontal ichthyoplankton tow-net system with unobstructed net opening. North American Journal of Fisheries Management, 7(1):148–150, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Newcomb:1989:WAC

[New89] Brad A. Newcomb. Winter abundance of channel catfish in the channelized Missouri River, Nebraska. North American Journal of Fisheries Management, 9(2):195–202, May 1989. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nettles:1987:MLA

[NG87] D. C. Nettles and S. P. Gloss. Migration of landlocked Atlantic salmon smolts and effectiveness of a fish bypass structure at a smallscale hydroelectric facility. North American Journal of Fisheries Management, 7(4):562–568, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Noel:1988:HSB

[NH88] Lynn E. Noel and Wayne A. Hubert. Harvest and sale of baitfish in Wyoming. North American Journal of Fisheries Management, 8 (4):511–515, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nicolette:1984:YOP

[Nic84] Joseph P. Nicolette. A 3-year-old pink salmon in an odd-year run in Lake Superior. North American Journal of Fisheries Management, 4(1):130–132, January 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nichols:1986:UYP

[Nic86] Scott Nichols. Updated yield-per-recruit information on the Tortugas pink shrimp fishery. North American Journal of Fisheries Management, 6(3):339–343, July 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nielsen:1982:BFI

[Nie82] Larry A. Nielsen. The bait-fish industry in Ohio and West Virginia, with special reference to the Ohio River sport fishery. North American Journal of Fisheries Management, 2(3):232–238, July 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nielsen:1988:HCB

[NO88] Larry A. Nielsen and Donald J. Orth. The hellgrammite-crayfish bait fishery of the New River and its tributaries, West Virginia. North American Journal of Fisheries Management, 8(3):317–324, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Noltie:1987:FMM

[Nol87] Douglas B. Noltie. Fencing material and methods for effectively trapping and holding pink salmon. North American Journal of Fisheries Management, 7(1):159–161, January 1987. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Novinger:1987:EIM

[Nov87] Gary D. Novinger. Evaluation of a 15.0-inch minimum length limit on largemouth bass and spotted bass catches at Table Rock Lake, Missouri. North American Journal of Fisheries Management, 7(2): 260–272, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nester:1984:FES

[NP84] Robert T. Nester and Thomas P. Poe. First evidence of successful natural reproduction of planted lake trout in Lake Huron. North American Journal of Fisheries Management, 4(1):126–128, January 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nester:1987:VOH

[NP87] R. T. Nester and T. P. Poe. Visual observations of historical lake trout spawning grounds in Western Lake Huron. North American Journal of Fisheries Management, 7(3):418–424, July 1987. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neill:1987:CFC

[NT87] Christopher Neill and R. Eugene Turner. Comparison of fish communities in open and plugged backfilled canals in Louisiana coastal marshes. North American Journal of Fisheries Management, 7(1): 57–62, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

OConnell:1987:ASE

[OB87] M. F. O’Connell and C. E. Bourgeois. Atlantic salmon enhancement in the Exploits River, Newfoundland, 1957–1984. North American Journal of Fisheries Management, 7(2):207–214, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Olson:1989:SFT

[OC89] Donald E. Olson and Paul K. Cunningham. Sport-fisheries trends shown by an annual Minnesota fishing contest over a 58-year period. North American Journal of Fisheries Management, 9(3):287–297, August 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Overholtz:1985:MMO

[Ove85] William J. Overholtz. Managing the multispecies otter trawl fisheries of Georges Bank with catch optimization methods. North American Journal of Fisheries Management, 5(2B):252–260, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paragamian:1982:CRH

[Par82] Vaughn L. Paragamian. Catch rates and harvest results under a 14.0-inch minimum length limit for largemouth bass in a New Iowa impoundment. North American Journal of Fisheries Management, 2(3):224–231, July 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paragamian:1984:ACI

[Par84] Vaughn L. Paragamian. Angler compliance with a 12.0-inch minimum length limit for smallmouth bass in Iowa streams. North American Journal of Fisheries Management, 4(2):228–229, April 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paragamian:1989:CDN

[Par89a] Vaughn L. Paragamian. A comparison of day and night electrofishing: Size structure and catch per unit effort for smallmouth bass. North American Journal of Fisheries Management, 9(4):500–503, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paragamian:1989:SHU

[Par89b] Vaughn L. Paragamian. Seasonal habitat use by walleye in a warmwater river system, as determined by radiotelemetry. North American Journal of Fisheries Management, 9(4):392–401, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Prince:1981:LBH

[PB81] Eric D. Prince and D. Hugh Barwick. Landlocked blueback herring in two South Carolina reservoirs: Reproduction and suitability as stocked prey. North American Journal of Fisheries Management, 1(1):41–45, January 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Persons:1982:FAS

[PB82] W. R. Persons and R. V. Bulkley. Feeding activity and spawning time of striped bass in the Colorado River Inlet, Lake Powell, Utah. North American Journal of Fisheries Management, 2(4):403–408, October 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Phippen:1987:ADT

[PB87] Kenneth W. Phippen and Eric P. Bergersen. Angling definitions and their effects on the accuracy of count-interview creel survey harvest estimates. North American Journal of Fisheries Management, 7(4):488–492, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Palmisano:1988:UPE

[PB88] Aldo N. Palmisano and Carl V. Burger. Use of a portable electric barrier to estimate Chinook salmon escapement in a turbid Alaskan River. North American Journal of Fisheries Management, 8(4): 475–480, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parkinson:1988:SSR

[PBB88] Eric A. Parkinson, Jonathan Berkowitz, and Christopher J. Bull. Sample size requirements for detecting changes in some fisheries statistics from small trout lakes. North American Journal of Fisheries Management, 8(2):181–190, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pennington:1983:FPA

[PBP83] C. H. Pennington, John A. Baker, and Michael E. Potter. Fish populations along natural and revetted banks on the Lower Mississippi River. North American Journal of Fisheries Management, 3(2): 204–211, April 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:1984:CRM

[PC84] N. P. Peterson and C. J. Cederholm. A comparison of the removal and mark-recapture methods of population estimation for juvenile Coho salmon in a small stream. North American Journal of Fisheries Management, 4(1):99–102, January 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:1981:FCB

[PCC81] Rodney B. Pierce, Daniel W. Coble, and Scott Corley. Fish catches in baited and unbaited hoop nets in the Upper Mississippi River. North American Journal of Fisheries Management, 1(2):204–206, April 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Puffer:1982:PHH

[PDA82] Harold W. Puffer, Marysia J. Duda, and Stanley P. Azen. Potential health hazards from consumption of fish caught in polluted coastal waters of Los Angeles County. North American Journal of Fisheries Management, 2(1):74–79, January 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Plumb:1988:SCR

[PGR88] J. A. Plumb, John M. Grizzle, and W. A. Rogers. Survival of caught and released largemouth bass after containment in live Wells. North American Journal of Fisheries Management, 8(3):325–328, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pardue:1985:UAS

[PH85] Garland B. Pardue and Kenneth B. Huntington. An unsuccessful attempt to stock striped bass as predators in new and unbalanced Virginia farm ponds. North American Journal of Fisheries Management, 5(2B):242–247, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parsons:1988:IHA

[PH88] Bradford G. M. Parsons and Wayne A. Hubert. Influence of habitat availability on spawning site selection by Kokanees in streams. North American Journal of Fisheries Management, 8(4):426–431, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peven:1989:PHN

[PH89] Charles M. Peven and Steven G. Hays. Proportions of hatcheryand naturally produced steelhead smolts migrating past rock Island Dam, Columbia River, Washington. North American Journal of Fisheries Management, 9(1):53–59, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pitlo:1989:WSH

[Pit89] John Pitlo, Jr. Walleye spawning habitat in Pool 13 of the Upper Mississippi River. North American Journal of Fisheries Management, 9(3):303–308, August 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ploskey:1982:BMR

[PJ82] Gene R. Ploskey and Robert M. Jenkins. Biomass model of reservoir fish and fish-food interactions, with implications for management. North American Journal of Fisheries Management, 2(2): 105–121, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Patrick:1987:FEH

[PM87] P. H. Patrick and R. S. McKinley. Field evaluation of a hidrostal pump for live transfer of American eels at a hydroelectric facility. North American Journal of Fisheries Management, 7(2):303–305, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Platts:1983:ECM

[PN83] William S. Platts and Rodger Loren Nelson. The electronic capacitance meter: a tool for evaluating riparian-fishery habitat. North American Journal of Fisheries Management, 3(3):219–227, July 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Platts:1985:IRR

[PN85] William S. Platts and Rodger Loren Nelson. Impacts of restrotation grazing on stream banks in forested watersheds in Idaho. North American Journal of Fisheries Management, 5(4):547–556, October 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Platts:1988:FTP

[PN88] William S. Platts and Rodger L. Nelson. Fluctuations in trout populations and their implications for land-use evaluation. North American Journal of Fisheries Management, 8(3):333–345, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Platts:1989:SCR

[PN89] William S. Platts and Rodger L. Nelson. Stream canopy and its relationship to salmonid biomass in the Intermountain West. North American Journal of Fisheries Management, 9(4):446–457, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pepper:1985:EEL

[POB85] V. A. Pepper, N. P. Oliver, and R. Blundon. Evaluation of an experiment in Lacustrine rearing of juvenile anadromous Atlantic salmon. North American Journal of Fisheries Management, 5 (4):507–525, October 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Poffenberger:1985:OFC

[Pof85] John R. Poffenberger. Operational and financial characteristics of reef-fish vessels in the South Atlantic and Gulf of Mexico areas. North American Journal of Fisheries Management, 5(3A):379–388, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Prager:1987:ULF

[POS87] Michael H. Prager, John F. O’Brien, and Saul B. Saila. Using lifetime fecundity to compare management strategies: a case history for striped bass. North American Journal of Fisheries Management, 7(3):403–409, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parsley:1989:VCE

[PPB89] Michael J. Parsley, Douglas E. Palmer, and Randy W. Burkhardt. Variation in capture efficiency of a beach seine for small fishes. North American Journal of Fisheries Management, 9(2):239–244, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Payer:1989:HMW

[PPP89] Ronald D. Payer, Rodney B. Pierce, and Donald L. Pereira. Hooking mortality of walleyes caught on live and artificial baits. North American Journal of Fisheries Management, 9(2):188–192, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Platts:1985:RSE

[PR85] William S. Platts and John N. Rinne. Riparian and stream enhancement management and research in the Rocky Mountains. North American Journal of Fisheries Management, 5(2A):115–125, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perry:1984:AGL

[PT84] Lance G. Perry and John A. Tranquilli. Age and growth of largemouth bass in a thermally altered reservoir, as determined from otoliths. North American Journal of Fisheries Management, 4(3): 321–330, July 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Platts:1984:FCL

[PW84] William S. Platts and Fred J. Wagstaff. Fencing to control livestock grazing on riparian habitats along streams: Is it a viable alternative? North American Journal of Fisheries Management, 4(3): 266–272, July 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paragamian:1987:EVS

[PW87a] Vaughn L. Paragamian and Michael J. Wiley. Effects of variable streamflows on growth of smallmouth bass in the Maquoketa River, Iowa. North American Journal of Fisheries Management, 7(3): 357–362, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perry:1987:CST

[PW87b] W. Guthrie Perry and Arthur Williams. Comparison of slat traps, wire cages, and various baits for commercial harvest of catfish. North American Journal of Fisheries Management, 7(2):283–287, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quimby:1983:ISF

[QD83] Alan J. Quimby and Nick C. Dudiak. An intertidal salmon fry trap. North American Journal of Fisheries Management, 3(2):212–215, April 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quinn:1982:AFW

[QR82] Stephen P. Quinn and Michael R. Ross. Annulus formation by white suckers and the reliability of pectoral fin rays for ageing them. North American Journal of Fisheries Management, 2(2):204–208, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quinn:1988:ERA

[Qui88] Stephen P. Quinn. Effectiveness of restricted areas in reducing incidental catches of game fish in a gill-net fishery. North American Journal of Fisheries Management, 8(2):224–230, May 1988. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quinn:1989:RRV

[Qui89] Stephen P. Quinn. Recapture rates of voluntarily released largemouth bass. North American Journal of Fisheries Management, 9(1):86–91, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Raymond:1988:EHD

[Ray88] Howard L. Raymond. Effects of hydroelectric development and fisheries enhancement on spring and summer Chinook salmon and steelhead in the Columbia River Basin. North American Journal of Fisheries Management, 8(1):1–24, February 1988. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rach:1987:CTB

[RB87] J. J. Rach and T. D. Bills. Comparison of three baits for trapping crayfish. North American Journal of Fisheries Management, 7(4):601–603, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rulifson:1986:MDM

[RC86] Roger A. Rulifson and John E. Cooper. A method to determine mesh-size selectivity in commercial menhaden purse seines. North American Journal of Fisheries Management, 6(3):359–366, July 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ritchie:1988:EOY

[RC88] Beverlee J. Ritchie and Peter J. Colby. Even-odd year differences in walleye year-class strength related to mayfly production. North American Journal of Fisheries Management, 8(2):210–215, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richards:1989:DAH

[RC89] Carl Richards and Phillip J. Cernera. Dispersal and abundance of hatchery-reared and naturally spawned juvenile Chinook salmon in an Idaho stream. North American Journal of Fisheries Management, 9(3):345–351, August 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reddin:1986:EDM

[Red86] David G. Reddin. Effects of different mesh sizes on gill-net catches of Atlantic salmon in Newfoundland. North American Journal of Fisheries Management, 6(2):209–215, April 1986. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reisenbichler:1987:BMH

[Rei87] Reginald R. Reisenbichler. Basis for managing the harvest of Chinook salmon. North American Journal of Fisheries Management, 7(4):589–591, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reisenbichler:1988:RBD

[Rei88] Reginald R. Reisenbichler. Relation between distance transferred from natal stream and recovery rate for hatchery Coho salmon. North American Journal of Fisheries Management, 8(2):172–174, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rettig:1984:LLU

[Ret84] R. Bruce Rettig. License limitation in the United States and Canada: an assessment. North American Journal of Fisheries Management, 4(3):231–248, July 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rainwater:1982:SCB

[RH82] William C. Rainwater and Alfred Houser. Species composition and biomass of fish in selected coves in Beaver Lake, Arkansas, during the first 18 years of impoundment (1963–1980). North American Journal of Fisheries Management, 2(4):316–325, October 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rensel:1988:FCS

[RHT88] John E. Rensel, Richard P. Harris, and Timothy J. Tynan. Fishery contribution and spawning escapement of Coho salmon reared in net-pens in Southern Puget Sound, Washington. North American Journal of Fisheries Management, 8(3):359–366, August 1988. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rinne:1982:MHR

[Rin82] John N. Rinne. Movement, home range, and growth of a rare southwestern trout in improved and unimproved habitats. North American Journal of Fisheries Management, 2(2):150–157, April 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rinne:1985:VAT

[Rin85] John N. Rinne. Variation in Apache trout populations in the White Mountains, Arizona. North American Journal of Fisheries Management, 5(2A):146–158, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rinne:1988:GES

[Rin88] John N. Rinne. Grazing effects on stream habitat and fishes: Research design considerations. North American Journal of Fisheries Management, 8(2):240–247, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reynolds:1988:EIL

[RKSC88] James B. Reynolds, A. Lawrence Kolz, Norman G. Sharber, and Steven W. Carothers. Electrofishing injury to large rainbow trout. North American Journal of Fisheries Management, 8(4):516–518, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rach:1982:WHB

[RM82] Jeffrey J. Rach and Fred P. Meyer. Winter harvest of bluegills in 1976–1977 from Lake Onalaska, Wisconsin. North American Journal of Fisheries Management, 2(1):28–32, January 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roberts:1986:ADP

[Rob86] Terrell W. Roberts. Abundance and distribution of pink shrimp in and around the Tortugas Sanctuary, 1981–1983. North American Journal of Fisheries Management, 6(3):311–327, July 1986. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rogers:1985:SPO

[Rog85] Barton D. Rogers. A small push-otter trawl for use in shallow marshes. North American Journal of Fisheries Management, 5 (3A):411–413, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rohrer:1986:CSK

[Roh86] Robert L. Rohrer. Creel surveys from a kayak on the Henry’s Fork of the Snake River, Idaho. North American Journal of Fisheries Management, 6(2):294–295, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rohrer:1987:HPR

[Roh87] Robert L. Rohrer. Harvest of planted rainbow trout of catchable size from five reaches of Henry’s Fork, Idaho. North American Journal of Fisheries Management, 7(1):142–144, January 1987. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Romaire:1983:ETD

[RP83] Robert P. Romaire and Vernon A. Pfister. Effects of trap density and diel harvesting frequency on catch of crawfish. North American Journal of Fisheries Management, 3(4):419–424, October 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rodgers:1985:EHP

[RP85] David W. Rodgers and Paul H. Patrick. Evaluation of a hidrostal pump fish return system. North American Journal of Fisheries Management, 5(3A):393–399, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reed:1989:CUS

[RR89] Michael S. Reed and Charles F. Rabeni. Characteristics of an unexploited smallmouth bass population in a Missouri Ozark stream. North American Journal of Fisheries Management, 9(4):420–426, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Risotto:1985:AFA

[RT85] Stephen P. Risotto and R. Eugene Turner. Annual fluctuation in abundance of the commercial fisheries of the Mississippi River and tributaries. North American Journal of Fisheries Management, 5 (4):557–574, October 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rohrer:1986:ETH

[RT86] Robert L. Rohrer and Gary H. Thorgaard. Evaluation of two hybrid trout strains in Henry’s Lake, Idaho, and comments on the potential use of sterile triploid hybrids. North American Journal of Fisheries Management, 6(3):367–371, July 1986. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reiser:1988:ETS

[RW88] Dudley W. Reiser and Robert G. White. Effects of two sediment size-classes on survival of steelhead and Chinook salmon eggs. North American Journal of Fisheries Management, 8(4):432–437, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:1984:AAF

[SA84] Edward J. Smith and John K. Andersen. Attempts to alleviate fish losses from Allegheny Reservoir, Pennsylvania and New York, using acoustics. North American Journal of Fisheries Management, 4(3): 300–307, July 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sayers:1987:EFW

[Say87] Richard E. Sayers, Jr. Effects of freezing in and out of water on length and weight of Lake Michigan bloaters. North American Journal of Fisheries Management, 7(2):299–301, April 1987. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sylvester:1983:DRA

[SB83] J. R. Sylvester and J. D. Broughton. Distribution and relative abundance of fish in Pool 7 of the Upper Mississippi River. North American Journal of Fisheries Management, 3(1):67–71, January 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scarnecchia:1987:TPS

[SB87] Dennis L. Scarnecchia and Eric P. Bergersen. Trout production and standing crop in Colorado’s small streams, as related to environmental features. North American Journal of Fisheries Management, 7(3):315–330, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sharp:1988:PEA

[SB88a] Dan Sharp and David R. Bernard. Precision of estimated ages of lake trout from five calcified structures. North American Journal of Fisheries Management, 8(3):367–372, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Swanson:1988:GCS

[SB88b] Eric D. Swanson and Eric P. Bergersen. Grass carp stocking model for coldwater lakes. North American Journal of Fisheries Management, 8(3):284–291, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stauffer:1988:CSS

[SBB88] Jay R. Stauffer, Jr., Suzanne E. Boltz, and Jeffrey M. Boltz. Cold shock susceptibility of blue tilapia from the Susquehanna River, Pennsylvania. North American Journal of Fisheries Management, 8(3):329–332, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sharber:1987:SES

[SC87] Norman G. Sharber and Steven W. Carothers. Submerged, electrically shielded live tank for electrofishing boats. North American Journal of Fisheries Management, 7(3):450–453, July 1987. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sharber:1988:IEP

[SC88] N. G. Sharber and S. W. Carothers. Influence of electrofishing pulse shape on spinal injuries in adult rainbow trout. North American Journal of Fisheries Management, 8(1):117–122, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schoning:1984:SIR

[Sch84] Robert W. Schoning. Some impacts of resource data use in fisheries management. North American Journal of Fisheries Management, 4 (1):1–8, January 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schoolmaster:1986:BFA

[Sch86] F. Andrew Schoolmaster. Bank- and float-angler perceptions of use levels on the Madison River, Montana. North American Journal of Fisheries Management, 6(3):430–438, July 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schaefer:1989:HMW

[Sch89a] W. F. Schaefer. Hooking mortality of walleyes in a Northwestern Ontario lake. North American Journal of Fisheries Management, 9(2):193–194, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schlueter:1989:SRS

[Sch89b] Lynn R. Schlueter. Simple removal of Sagitta otoliths from walleyes. North American Journal of Fisheries Management, 9(1): 123, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scogin:1983:LMR

[Sco83] William Malon Scogin, Jr. The licensing of marine recreational fishermen. North American Journal of Fisheries Management, 3(3): 276–282, July 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scrivener:1988:TDA

[Scr88] J. C. Scrivener. Two devices to assess incubation survival and emergence of salmonid fry in an estuary streambed. North American Journal of Fisheries Management, 8(2):248–258, May 1988. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stevenson:1987:MGC

[SD87] Frederick Stevenson and Richard E. Day. Mortality and growth of channel catfish following extraction of a pectoral spine. North American Journal of Fisheries Management, 7(3):443–445, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Seelbach:1987:SSH

[See87a] Paul W. Seelbach. Smolting success of hatchery-raised steelhead planted in a Michigan tributary of Lake Michigan. North American Journal of Fisheries Management, 7(2):223–231, April 1987. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Seelye:1987:TTL

[See87b] James G. Seelye. Toxicity of TFM lampricide to early life stages of walleye. North American Journal of Fisheries Management, 7 (4):598–601, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Serns:1982:RWF

[Ser82a] Steven L. Serns. Relationship of walleye fingerling density and electrofishing catch per effort in Northern Wisconsin lakes. North American Journal of Fisheries Management, 2(1):38–44, January 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Serns:1982:WFP

[Ser82b] Steven L. Serns. Walleye fecundity, potential egg deposition, and survival from egg to fall young-of-year in Escanaba Lake, Wisconsin, 1979–1981. North American Journal of Fisheries Management, 2(4):388–394, October 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Serns:1983:RBE

[Ser83] Steven L. Serns. Relationship between electrofishing catch per unit effort and density of walleye yearlings. North American Journal of Fisheries Management, 3(4):451–452, October 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Serns:1987:RBS

[Ser87] Steven L. Serns. Relationship between the size of several walleye year classes and the percent harvested over the life of each cohort in Escanaba Lake, Wisconsin. North American Journal of Fisheries Management, 7(2):305–306, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schill:1986:HMC

[SGG86] Daniel J. Schill, J. S. Griffith, and Robert E. Gresswell. Hooking mortality of cutthroat trout in a catch-and-release segment of the Yellowstone River, Yellowstone National Park. North American Journal of Fisheries Management, 6(2):226–232, April 1986. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scarnecchia:1989:RHS

[SGM89] Dennis L. Scarnecchia, Thomas W. Gengerke, and Clark T. Moen. Rationale for a harvest slot limit for paddlefish in the Upper Mississippi River. North American Journal of Fisheries Management, 9 (4):477–487, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Spiller:1988:IEA

[SGO88] Kyle W. Spiller, Albert W. Green, and Hal R. Osburn. Increasing the efficiency of angler surveys by canceling sampling during inclement weather. North American Journal of Fisheries Management, 8(1):132–138, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Swenson:1988:CMB

[SGS88] William A. Swenson, William P. Gobin, and Timothy D. Simonson. Calibrated mask-bar for underwater measurement of fish. North American Journal of Fisheries Management, 8(3):382–385, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schneidervin:1986:RTO

[SH86a] Roger W. Schneidervin and Wayne A. Hubert. A rapid technique for otolith removal from salmonids and catostomids. North American Journal of Fisheries Management, 6(2):287, April 1986. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sharber:1986:CMH

[SH86b] Norman G. Sharber and Mark Hudy. Comments: Mortality from high-voltage AC electrofishing. North American Journal of Fisheries Management, 6(1):134, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schneidervin:1987:DOA

[SH87] Roger W. Schneidervin and Wayne A. Hubert. Diet overlap among zooplanktophagous fishes in Flaming Gorge Reservoir, Wyoming– Utah. North American Journal of Fisheries Management, 7(3): 379–385, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stoeckel:1988:OIS

[SH88] Joseph N. Stoeckel and Roy C. Heidinger. Overwintering of the inland silverside in Southern Illinois. North American Journal of Fisheries Management, 8(1):127–131, February 1988. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Swink:1989:SRT

[SH89] William D. Swink and Lee H. Hanson. Survival of rainbow trout and lake trout after sea lamprey attack. North American Journal of Fisheries Management, 9(1):35–40, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schramm:1985:STC

[SHB85] Harold L. Schramm, Jr., Paul J. Haydt, and Nicholas A. Bruno. Survival of tournament-caught largemouth bass in two Florida lakes. North American Journal of Fisheries Management, 5(4):606– 611, October 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schramm:1987:EPP

[SHP87] Harold L. Schramm, Jr., Paul J. Haydt, and Kenneth M. Portier. Evaluation of prerelease, postrelease, and total mortality, of largemouth bass caught during tournaments, in two Florida lakes. North American Journal of Fisheries Management, 7(3):394–402, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Summers:1983:RSS

[SHR83] J. K. Summers, H. W. Hoffman, and W. A. Richkus. Randomized sample surveys to estimate annual blue crab harvests by a multigear fishery in the Maryland waters of Chesapeake Bay. North American Journal of Fisheries Management, 3(1):9–20, January 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sager:1985:PWV

[SHS85] David R. Sager, Charles H. Hocutt, and Jay R. Stauffer, Jr. Preferred wavelengths of visible light for juvenile Atlantic menhaden. North American Journal of Fisheries Management, 5(1):72–77, January 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Swink:1983:IKF

[SJ83] William D. Swink and Kenneth E. Jacobs. Influence of a Kentucky flood-control reservoir on the tailwater and headwater fish populations. North American Journal of Fisheries Management, 3(2): 197–203, April 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sheppard:1985:PUD

[SJ85] J. Douglas Sheppard and James H. Johnson. Probability-of-use for depth, velocity, and substrate by subyearling Coho salmon and steelhead in Lake Ontario tributary streams. North American Journal of Fisheries Management, 5(2B):277–282, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stanford:1982:BII

[SJT^{+}82] Richard M. Stanford, Scott W. Jordan, Daniel R. Talhelm, Charles R. Liston, Charles Korson, and Milton H. Steinmueller. The bioeconomic impact of impingement and entrainment on yellow perch in Lake Erie. North American Journal of Fisheries Management, 2(4):285–293, July 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Strange:1982:ESR

[SKB82] Richard J. Strange, William B. Kittrell, and Timothy D. Broadbent. Effects of seeding reservoir fluctuation zones on young-of-the-year black bass and associated species. North American Journal of Fisheries Management, 2(4):307–315, October 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Seelbach:1985:MIS

[SLA85] Paul W. Seelbach, Roger N. Lockwood, and Gaylord R. Alexander. A modified inclined-screen trap for catching salmonid smolts in large rivers. North American Journal of Fisheries Management, 5(3B):494–498, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stone:1982:GSL

[SM82] Clifton C. Stone and Timothy Modde. Growth and survival of largemouth bass in newly stocked South Dakota ponds. North American Journal of Fisheries Management, 2(4):326–333, October 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stevens:1983:ERF

[SM83] Donald E. Stevens and Lee W. Miller. Effects of river flow on abundance of young Chinook salmon, American shad, longfin smelt, and delta smelt in the Sacramento–San Joaquin River System. North American Journal of Fisheries Management, 3(4):425–437, October 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schlagenhaft:1985:HUO

[SM85] Timothy W. Schlagenhaft and Brian R. Murphy. Habitat use and overlap between adult largemouth bass and walleye in a West Texas reservoir. North American Journal of Fisheries Management, 5 (3B):465–470, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:1986:SGR

[SM86] Courtland L. Smith and Robert McKelvey. Specialist and generalist: Roles for coping with variability. North American Journal of Fisheries Management, 6(1):88–99, January 1986. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Slaney:1987:AUC

[SM87] P. A. Slaney and A. D. Martin. Accuracy of underwater census of trout populations in a large stream in British Columbia. North American Journal of Fisheries Management, 7(1):117–122, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sayers:1989:IRS

[SMJR89] Richard E. Sayers, Jr., John R. Moring, Paul R. Johnson, and Scott A. Roy. Importance of rainbow smelt in the winter diet of landlocked Atlantic salmon in four Maine lakes. North American Journal of Fisheries Management, 9(3):298–302, August 1989. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scarsbrook:1988:EEE

[SMS88] J. R. Scarsbrook, G. A. McFarlane, and W. Shaw. Effectiveness of experimental escape mechanisms in sablefish traps. North American Journal of Fisheries Management, 8(2):158–161, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:1984:ASF

[SMU84] Theodore I. J. Smith, Donald E. Marchette, and Glen F. Ulrich. The Atlantic sturgeon fishery in South Carolina. North American Journal of Fisheries Management, 4(2):164–176, April 1984. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Storck:1988:ESS

[SN88] Ted Storck and Dennis Newman. Effects of size at stocking on survival and harvest of channel catfish. North American Journal of Fisheries Management, 8(1):98–101, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Summers:1983:ARS

[SRH^{+}83] J. Kevin Summers, William A. Richkus, Howard W. Hoffman, Christopher Bonzek, Howard H. King, and Michael Burch. Application of random sample survey designs to estimate the commercial blue crab harvest in Maryland. North American Journal of Fisheries Management, 3(1):21–25, January 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schneeberger:1982:GTN

[SRJ82] Philip J. Schneeberger, Thomas L. Rutecki, and David J. Jude. Gilling in trap-net pots and use of catch data to predict lake whitefish gilling rates. North American Journal of Fisheries Management, 2(4):294–300, July 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schweigert:1988:EME

[SS88] J. F. Schweigert and M. Stocker. Escapement model for estimating Pacific herring stock size from spawn survey data and its management implications. North American Journal of Fisheries Management, 8(1):63–74, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stuber:1985:RTR

[SSB85] Robert J. Stuber, Clee Sealing, and Eric P. Bergersen. Rainbow trout returns from fingerling plantings in Dillon Reservoir, Colorado, 1975–1979. North American Journal of Fisheries Management, 5(3B):471–474, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schramm:1985:ECH

[SSHP85] Harold L. Schramm, Jr., Jerome V. Shireman, Dennis E. Hammond, and Duncan M. Powell. Effect of commercial harvest of sport fish on the black crappie population in Lake Okeechobee, Florida. North American Journal of Fisheries Management, 5(2B):217–226, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scheerer:1985:PDS

[ST85] Paul D. Scheerer and William W. Taylor. Population dynamics and stock differentiation of lake whitefish in Northeastern Lake Michigan with implications for their management. North American Journal of Fisheries Management, 5(4):526–536, October 1985. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schultz:1987:PSO

[ST87] David L. Schultz and R. Scott Taylor. Preparation of small otoliths for microscopic examination. North American Journal of Fisheries Management, 7(2):309–311, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stange:1981:OLA

[Sta81] Doug Stange. An open letter to anglers and fisheries professionals. North American Journal of Fisheries Management, 1(2):193–199, April 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stober:1983:BNR

[STP83] Q. J. Stober, R. W. Tyler, and C. E. Petrosky. Barrier net to reduce entrainment losses of adult Kokanee from Banks Lake, Washington. North American Journal of Fisheries Management, 3(4):331–354, October 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Spinelli:1985:ASE

[SWH85] Anthony J. Spinelli, B. G. Whiteside, and D. G. Huffman. Aquarium studies on the evaluation of stocking various sizes of channel catfish with established largemouth bass. North American Journal of Fisheries Management, 5(2A):138–145, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sztramko:1985:ESS

[Szt85] L. Kazimierz Sztramko. Effects of a sanctuary on the smallmouth bass fishery of Long Point Bay, Lake Erie. North American Journal of Fisheries Management, 5(2B):233–241, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tappel:1986:LUG

[Tap86] Paul D. Tappel. Limitations on the use of gabions to improve fish passage. North American Journal of Fisheries Management, 6(1):131–132, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:1981:GIF

[Tay81] Melvin W. Taylor. A generalized inland fishery simulator for management biologists. North American Journal of Fisheries Management, 1(1):60–72, January 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tappel:1983:NMR

[TB83] Paul D. Tappel and Ted C. Bjornn. A new method of relating size of spawning gravel to salmonid embryo survival. North American Journal of Fisheries Management, 3(2):123–135, April 1983. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Threader:1986:BML

[TB86] R. W. Threader and C. S. Broussaeu. Biology and management of the lake sturgeon in the Moose River, Ontario. North American Journal of Fisheries Management, 6(3):383–390, July 1986. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tyus:1984:URO

[TBM84] Harold M. Tyus, Bob D. Burdick, and C. W. McAda. Use of radiotelemetry for obtaining habitat preference data on Colorado squawfish. North American Journal of Fisheries Management, 4(2):

177–180, April 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tranquilli:1982:GSL

[TC82] John A. Tranquilli and William F. Childers. Growth and survival of largemouth bass tagged with floy anchor tags. North American Journal of Fisheries Management, 2(2):184–187, April 1982. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Timmons:1985:ERS

[TG85] Tom J. Timmons and William E. Garrett. Effectiveness of residual stumps as fish attractors in Aliceville Lake, Alabama–Mississippi. North American Journal of Fisheries Management, 5(2B):309–310, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tettey:1986:ITP

[TGPS86] Ernest O. Tettey, Wade L. Griffin, John B. Penson, and John R. Stoll. Implications of tax policy on investment in a commonproperty resource. North American Journal of Fisheries Management, 6(1):100–104, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tomljanovich:1986:PGS

[TH86a] David A. Tomljanovich and John H. Heuer. Passage of gizzard shad and threadfin shad larvae through a larval fish net with 500-μ m openings. North American Journal of Fisheries Management, 6(2): 256–259, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tsumura:1986:TVS

[TH86b] Kanji Tsumura and Jeremy M. B. Hume. Two variations of a salmonid smolt trap for small rivers. North American Journal of Fisheries Management, 6(2):272–276, April 1986. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Timmons:1989:ISF

[THHS89] Tom J. Timmons, Tim Hoffnagle, R. Scott Hale, and John B. Soldo. Incidence of sport fishes in the commercial fish catch from Kentucky Lake, Kentucky and Tennessee. North American Journal of Fisheries Management, 9(2):209–212, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thomas:1985:EDI

[Tho85] Virginia G. Thomas. Experimentally determined impacts of a small, suction gold dredge on a Montana stream. North American Journal of Fisheries Management, 5(3B):480–488, July 1985. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tave:1981:RHA

[TMCS81] Douglas Tave, Andrew S. McGinty, Jesse A. Chappell, and R. O. Smitherman. Relative harvestability by angling of blue catfish, channel catfish, and their reciprocal hybrids. North American Journal of Fisheries Management, 1(1):73–76, January 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:1989:SHS

[TMO89] D. M. Taylor, G. W. Marshall, and P. G. O’Keefe. Shell hardening in snow crabs Chionoecetes opilio tagged in soft-shelled condition. North American Journal of Fisheries Management, 9(4):504–508, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Townsend:1985:EER

[Tow85] Ralph E. Townsend. An economic evaluation of restricted entry in Maine’s soft-shell clam industry. North American Journal of Fisheries Management, 5(1):57–64, January 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:1984:IUC

[TP84] Timothy G. Taylor and Fred J. Prochaska. Incorporating unobserved cyclical stock movements in fishery catch equations: an application to the Florida blue crab fishery. North American Journal of Fisheries Management, 4(1):67–74, January 1984. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Temple:1985:IMS

[TPM85] Alan J. Temple, David J. Price, and Brian R. Murphy. An inexpensive method for sectioning otoliths of large fish. North American Journal of Fisheries Management, 5(4):612, October 1985. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Timmons:1982:MFA

[TS82] Tom J. Timmons and William L. Shelton. Marking fish attractors in lakes with fluctuating water levels. North American Journal of Fisheries Management, 2(1):97, January 1982. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tsumura:1987:STI

[Tsu87] Kanji Tsumura. Simple technique to improve microfiche prints of fish scales. North American Journal of Fisheries Management, 7(3):441–443, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tsumura:1987:SUB

[TW87] Kanji Tsumura and Bruce R. Ward. Sling used for boulder placement in instream enhancement projects. North American Journal of Fisheries Management, 7(3):456–457, July 1987. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tyus:1988:LTR

[Tyu88] Harold M. Tyus. Long-term retention of implanted transmitters in Colorado squawfish and razorback sucker. North American Journal of Fisheries Management, 8(2):264–267, May 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ueber:1986:FYD

[Ueb86] Edward Ueber. Fillet yields of Dover sole versus depth of capture and length. North American Journal of Fisheries Management, 6(2):282–284, April 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanDenAvyle:1983:HPF

[VHJB83] Michael J. Van Den Avyle, Billy J. Higginbotham, Brian T. James, and Frank J. Bulow. Habitat preferences and food habits of youngof-the-year striped bass, white bass, and yellow bass in Watts Bar Reservoir, Tennessee. North American Journal of Fisheries Management, 3(2):163–170, April 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vincent:1987:ESC

[Vin87] E. Richard Vincent. Effects of stocking catchable-size hatchery rainbow trout on two wild trout species in the Madison River and O’Dell Creek, Montana. North American Journal of Fisheries Management, 7(1):91–105, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Villegas:1982:MSS

[VJL82] Luis Villegas, Albert C. Jones, and Ronald F. Labisky. Management strategies for the spiny lobster resources in the Western Central Atlantic: a cooperative approach. North American Journal of Fisheries Management, 2(3):216–223, July 1982. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wagner:1981:RPL

[Wag81] Wilbert C. Wagner. Reproduction of planted lake trout in Lake Michigan. North American Journal of Fisheries Management, 1(2): 159–164, April 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walsh:1988:DVT

[Wal88] S. J. Walsh. Diel variability in trawl catches of juvenile and adult yellowtail flounder on the Grand Banks and the effect on resource assessment. North American Journal of Fisheries Management, 8 (3):373–381, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weisberg:1987:RIE

[WBJR87] Stephen B. Weisberg, William H. Burton, Fred Jacobs, and Eric A. Ross. Reductions in ichthyoplankton entrainment with fine-mesh, wedge-wire screens. North American Journal of Fisheries Management, 7(3):386–393, July 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wooley:1983:BPE

[WC83] Charles M. Wooley and Edouard J. Crateau. Biology, population estimates, and movement of, native and introduced striped bass, Apalachicola River, Florida. North American Journal of Fisheries Management, 3(4):383–394, October 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wooley:1985:MME

[WC85] Charles M. Wooley and Edouard J. Crateau. Movement, microhabitat, exploitation, and management of Gulf of Mexico sturgeon, Apalachicola River, Florida. North American Journal of Fisheries Management, 5(4):590–605, October 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weaver:1986:RTD

[WE86] Oda R. Weaver and Russell H. England. Return of tags with different rewards in Lake Lanier, Georgia. North American Journal of Fisheries Management, 6(1):132–133, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wertheimer:1988:HMC

[Wer88] Alex Wertheimer. Hooking mortality of Chinook salmon released by commercial trollers. North American Journal of Fisheries Management, 8(3):346–355, August 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilkins:1983:CPO

[WG83] Mark E. Wilkins and James T. Golden. Condition of the Pacific Ocean perch resource off Washington and Oregon during 1979: Results of a cooperative trawl survey. North American Journal of Fisheries Management, 3(2):103–122, April 1983. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wesche:1987:CRV

[WGF87] Thomas A. Wesche, Chris M. Goertler, and Carrie B. Frye. Contribution of riparian vegetation to trout cover in small streams. North American Journal of Fisheries Management, 7(1):151–153, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wesche:1987:MHS

[WGH87] Thomas A. Wesche, Chris M. Goertler, and Wayne A. Hubert. Modified habitat suitability index model for brown trout in Southeastern Wyoming. North American Journal of Fisheries Management, 7(2):232–237, April 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weiss-Glanz:1984:PSI

[WGS84] Lori S. Weiss-Glanz and Jon G. Stanley. Population structure indices of largemouth bass and smallmouth bass determined from angler catches. North American Journal of Fisheries Management, 4(1):89–98, January 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Waldman:1988:RSD

[WGW88] John R. Waldman, Joseph Grossfield, and Isaac Wirgin. Review of stock discrimination techniques for striped bass. North American Journal of Fisheries Management, 8(4):410–425, November 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wiley:1984:RBA

[WGWP84] M. J. Wiley, R. W. Gorden, S. W. Waite, and T. Powless. The relationship between aquatic macrophytes and sport fish production in Illinois ponds: a simple model. North American Journal of Fisheries Management, 4(1):111–119, January 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Whitworth:1989:MBE

[Whi89] Walter R. Whitworth. Management brief: Evaluation of largemouth bass growth estimates obtained from angler-collected scale samples. North American Journal of Fisheries Management, 9 (1):116–119, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wiley:1984:RSC

[Wil84] Robert W. Wiley. A review of sodium cyanide for use in sampling stream fishes. North American Journal of Fisheries Management, 4(3):249–256, July 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Willis:1987:RRG

[Wil87a] David W. Willis. Reproduction and recruitment of gizzard shad in Kansas reservoirs. North American Journal of Fisheries Management, 7(1):71–80, January 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Willis:1987:UGN

[Wil87b] David W. Willis. Use of gill-net data to provide a recruitment index for walleyes. North American Journal of Fisheries Management, 7 (4):591–592, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Willis:1989:PSL

[Wil89] David W. Willis. Proposed standard length–weight equation for northern pike. North American Journal of Fisheries Management, 9(2):203–208, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Winans:1989:GVC

[Win89] Gary A. Winans. Genetic variability in Chinook salmon stocks from the Columbia River Basin. North American Journal of Fisheries Management, 9(1):47–52, February 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Willis:1986:FSC

[WJ86] David W. Willis and Lanny D. Jones. Fish standing crops in wooded and nonwooded coves of Kansas reservoirs. North American Journal of Fisheries Management, 6(1):105–108, January 1986. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Welcomme:1983:SEM

[WKC83] R. L. Welcomme, Christopher C. Kohler, and Walter R. Courtenay, Jr. Stock enhancement in the management of freshwater fisheries: a European perspective. North American Journal of Fisheries Management, 3(3):265–275, July 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wentworth:1984:FYS

[WL84] Roderick S. Wentworth and George W. Labar. First-year survival and growth of steelhead stocked as fry in Lewis Creek, Vermont. North American Journal of Fisheries Management, 4(1):103–110, January 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Willis:1985:CSD

[WMG85] David W. Willis, Kenneth D. McCloskey, and Donald W. Gabelhouse, Jr. Calculation of stock density indices based on adjustments for efficiency of gill-net mesh size. North American Journal of Fisheries Management, 5(2A):126–137, April 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Whelan:1989:ITD

[WOH89] W. G. Whelan, M. F. O’Connell, and R. N. Hefford. Improved trap design for counting migrating fish in rivers. North American Journal of Fisheries Management, 9(2):245–248, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wolfert:1981:CFW

[Wol81] D. R. Wolfert. The commercial fishery for walleyes in New York waters of Lake Erie, 1959–1978. North American Journal of Fisheries Management, 1(2):112–126, April 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Workman:1981:RRT

[Wor81] Dennis L. Workman. Recovery of rainbow trout and brown trout populations following chemical poisoning in Sixteenmile Creek, Montana. North American Journal of Fisheries Management, 1(2): 144–150, April 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wesche:1989:NTM

[WRH^{+}89] Thomas A. Wesche, Dudley W. Reiser, Victor R. Hasfurther, Wayne A. Hubert, and Quentin D. Skinner. New technique for measuring fine sediment in streams. North American Journal of Fisheries Management, 9(2):234–238, May 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wright:1981:CPS

[Wri81] Sam Wright. Contemporary Pacific salmon fisheries management. North American Journal of Fisheries Management, 1(1):29–40, January 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Willis:1987:RBS

[WS87] David W. Willis and James L. Stephen. Relationships between storage ratio and population density, natural recruitment, and stocking success of walleye in Kansas reservoirs. North American Journal of Fisheries Management, 7(2):279–282, April 1987. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Willis:1989:RBP

[WS89] David W. Willis and Charles G. Scalet. Relations between proportional stock density and growth and condition of northern pike populations. North American Journal of Fisheries Management, 9 (4):488–492, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wiley:1983:REE

[WT83] Martin L. Wiley and Chu-Fa Tsai. The relative efficiencies of electrofishing vs. seines in Piedmont streams of Maryland. North American Journal of Fisheries Management, 3(3):243–253, July 1983. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wespestad:1984:BEY

[WT84] Vidar G. Wespestad and Joseph M. Terry. Biological and economic yields for Eastern Bering Sea walleye pollock under differing fishing regimes. North American Journal of Fisheries Management, 4(2): 204–215, April 1984. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williams:1989:RMT

[WW89] Cynthia Deacon Williams and Jack E. Williams. Refuge management for the threatened Railroad Valley springfish in Nevada. North American Journal of Fisheries Management, 9(4):465–470, Fall 1989. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Yeager:1985:CHS

[Yea85] David M. Yeager. Creation of a hybrid striped bass fishery in the Escambia River, Florida. North American Journal of Fisheries Management, 5(3A):389–392, July 1985. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zale:1987:PHS

[Zal87] Alexander V. Zale. Periodicity of habitation of a stenothermal spring run in North-Central Florida by blue tilapia. North American Journal of Fisheries Management, 7(4):575–579, October 1987. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zubik:1988:CSM

[ZF88] Raymond J. Zubik and John J. Fraley. Comparison of snorkel and mark–recapture estimates for trout populations in large streams. North American Journal of Fisheries Management, 8(1):58–62, February 1988. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zuboy:1981:NTF

[Zub81] James R. Zuboy. A new tool for fishery managers: The Delphi technique. North American Journal of Fisheries Management, 1(1): 55–59, January 1981. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).